### v2

# Stationary solution to Stochastically Forced Euler-Poisson Equations in Bounded Domain: Part 1. 3-D Insulating Boundary

Abstract.
This paper is concerned with $3$-D stochastic Euler-Poisson equations with insulating boundary conditions forced by the Wiener process. We first establish the global existence and uniqueness of the solution to the system, then we prove that the solution converges to its steady-state time-asymptotically. Compared with the deterministic case, due to the stochastic forces, the momentum does not have the temporal derivative, which causes some difficulty to the energy estimates. To overcome the obstacle caused by the stochastic force, we develop some new techniques to treat the stochastic problems. Furthermore, we observe that the invariant measure is just the Dirac measure generated by the steady-state, in which the time-exponential convergence rate to the steady-state plays an essential role.
Keywords: Stability, stationary solution, stochastic Euler-Poisson equations, cylindrical Brownian motion, insulating boundary conditions
2020 AMS Subject Classification: 34D05, 34D20,35B35, 35K51, 60H15, 82D37.
1. Introduction
Euler-Poisson equations is important in the analysis and design of semiconductor devices, offering a more precise description of physical phenomena [12] compared to the conventional drift-diffusion model. Furthermore, in the extreme ultraviolet (EUV) lithography, stochastic effects sometimes cause unwanted defects and pattern roughness in chips [3], that may impact the performance of a chip, or cause a device to fail. Hence, there is a pressing need to investigate the dynamic model of semiconductors perturbed by stochastic forces within mathematical frameworks. The stochastically forced Euler-Poisson equations (SEP for short) in a bounded smooth domain $U\subset\mathds{R}^{3}$ reads as
$\left\{\begin{array}[]{l}\rho_{t}+\nabla\cdot\left(\rho{\bf u}\right)=0,\\ \operatorname{d}_{t}\left(\rho{\bf u}\right)+\left(\nabla\cdot\left(\rho{\bf u}\otimes{\bf u}\right)+\nabla P\left(\rho\right)-\rho\nabla\Phi\right)\operatorname{d}t=-\frac{\rho{\bf u}}{\tau}\operatorname{d}t+\mathbb{F}\left(\rho,\mathbf{u}\right)\operatorname{d}_{t}W,\\ \triangle\Phi=\rho-b,\end{array}\right.$ (1.1)
where $\operatorname{d}_{t}$ is the differential notation with respect to time $t$, $\rho$ is the electron density of semiconductors, ${\bf u}$ denotes the particle velocity, $P\left(\rho\right)$ is the pressure, $\Phi$ is the electrostatic potential, $\tau$ is the velocity relaxation time and $b(x)$ is called the doping profile, which is positive and immobile. The above mentioned unknowns $\rho=\rho(\omega,t,x)$, ${\bf u}={\bf u}(\omega,t,x)$, $\Phi=\Phi$, and $P\left(\rho\right)=P\left(\rho(\omega,t,x)\right)$ are stochastic processes as functions with respect to $\omega$, $t$, and $x$, where $\omega$ is a sample in the probability space $\left(\Omega,\mathbb{P}\right)$. For convenience, we use the simplified notions $\rho$, ${\bf u}$, $\Phi$, and $P\left(\rho\right)$ here and hereafter. $W$ is an $\mathcal{H}$-valued cylindrical Brownian motion defined on the filtrated probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$, where $\mathcal{H}$ is an auxiliary separable Hilbert space, $\mathcal{F}$ is the filtration, see the definitions of filtration and Wiener process in Appendix 5.
[…]
$\langle\mathbb{F}\left(\rho,\mathbf{u}\right),e_{k}\rangle=\mathbf{F}_{k}\left(\rho,{\bf u}\right)$ (1.2)
is an $H$-valued vector function, which shows the strength of the external stochastic forces by
$\mathbb{F}\left(\rho,\mathbf{u}\right)\operatorname{d}_{t}W=\sum\limits_{k=1}^{+\infty}\mathbf{F}_{k}\left(\rho,{\bf u}\right)\operatorname{d}_{t}\beta_{k}.$ (1.3)
Throughout the paper, we assume that
$\mathbf{F}_{k}\left(\rho,{\bf u}\right)=a_{k}\rho{\bf u}Y\left(\rho,{\bf u}\right),$ (1.4)
where ${a_{k}}$ are positive constants, $Y\left(\rho,{\bf u}\right)$ is a smooth function of $\rho$ and ${\bf u}$, satisfying
$\sum a_{k}^{2}=1,\quad\left|Y\left(\rho,{\bf u}\right)\right|\leqslant C\left|\rho{\bf u}\right|,\quad\left\|\nabla_{\rho,{\bf u}}Y\left(\rho,{\bf u}\right)\right\|_{L^{\infty}}\leqslant C,\quad\left\|\nabla_{\rho,{\bf u}}^{2}Y\left(\rho,{\bf u}\right)\right\|_{L^{\infty}}\leqslant C.$ (1.5)
Here, $\nabla_{\rho,{\bf u}}$ denotes the differential operator with respect to $\rho$ and ${\bf u}$.
Subjected to the stochastic Euler-Poisson equations (1.1), the proposed boundary is the insulating boundary:
${\bf u}\cdot\nu=0,\quad\nabla\Phi\cdot\nu=0,$ (1.6)
where $\nu$ is the outer normal vector of $U$; and the initial data is:
$(\rho,{\bf u}_{0},\Phi)|_{t=0}=\left(\rho_{0}\left(\omega,x\right),~{}{\bf u}_{0}\left(\omega,x\right),~{}\Phi_{0}\left(\omega,x\right)\right),$ (1.7)
which is given in the probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$. Here and hereafter, we simply denote the initial data by $\left(\rho_{0},{\bf u}_{0},\Phi_{0}\right)$ without confusion.
The hydrodynamic model of semiconductors was first introduced by Blotekjaer [4], which is the deterministically dynamical model presented by Euler-Poisson equations mathematically. For 1-D case, the initial-boundary value problems to Euler-Poisson equations with the insulating boundary and the Ohmic contact boundary were studied by Hsiao-Yang [13], Li-Markowich-Mei [23], respectively, where the solutions are showed to converge to the corresponding subsonic steady-states time-asymptotically, where the doping profile is needed to be flat: $|b^{\prime}(x)|\ll 1$. Such a restriction was then released by Nishibata-Suzuki [27] and Guo-Strass [12] independently. For $N$-D case, Guo-Strass [12] first considered the deterministic 3-D Euler-Poisson equations in bounded domain with insulating boundary, and showed the convergence of solutions to the 3-D subsonic steady-states. Subsequently, Mei-Wu-Zhang [26] investigated the convergence to the steady-states for the n-D radial Euler-Poisson equations with the Ohmic contact boundary. For the whole space without boundary effects, the Cauchy problems to deterministic Euler-Poisson equations were extensively studied in [9, 14, 15, 16, 17, 20]. For the case of free boundary with vacuum, we refer to [24, 25, 29] and the references therein.
When the hydrodynamic model of semiconductors is counted into the stochastic affections, it then becomes the stochastic Euler-Poisson equations with uncertain extra disturbances, see (1.1) with the Wiener process $\mathbb{F}\left(\rho,\mathbf{u}\right)\operatorname{d}_{t}W$. This is a new model for semiconductor devices and never touched yet. The main issue of the paper is to investigate this 3-D stochastic Euler-Poisson equations in bounded domain with insulating boundary, and are going to prove the convergence of solutions to the stochastic steady states. The coefficient function of Wiener process $\mathbf{F}_{k}\left(\rho,{\bf u}\right)$, depending on the solutions $\rho$ and ${\bf u}$, is called the multiplicative noise. In most cases, the multiplicative noise magnifies the perturbation and thereby complicating the well-posedness of solutions for evolution systems. The stochastic forces are at most Hölder-$\frac{1}{2}-$continuous in time $t$, resulting in reduced regularity of velocity with respect to time. So from a mathematical standpoint, the study of the stochastic problem helps us to study how the solutions to stochastic Euler-Poisson equations behave in the absence of strong regularity in time. Further, this encourages exploring whether the desirable property remains under the influence of particular types of noise. This is the first attempt to study the asymptotic behavior of solutions to stochastic 3-D Euler-Poisson equations.
For stochastic evolution systems, the solution is called the stationary solution provided that the increment of solutions during evolution is time-independent. Da Prato-Zabczyk [8] listed the basic theory of stationary solutions of general stochastic PDEs in view of invariant measure. As an application of invariant measure for concrete models, Breit-Feireisl-Hofmanová-Maslowski [6] proved the existence of stationary solutions for stochastic compressible Navier-Stokes equations. Gess-Souganidis [11] obtained the large time behavior and existence of invariant measure for the stochastic scalar conservation laws, and the algebraic decay in time holds. In [11], they gave a kind of noise which has better regularization to the problem. Then Dong-Zhang-Zhang [10] proved the existence of stationary solutions to stochastic scalar conservation laws with the multiplicative noise. Bedrossian-Liss [1] gave the existence of stationary measures for stochastic differential equations with a nonlinear term.
It is worth noting that the stationary solution we consider is in view of invariant measure. In this paper, the concepts of stationary solution for stochastically forced system (1.1) and steady state $\left(\bar{\rho}(\omega,x),\bar{{\bf u}}(\omega,x),\bar{\Phi}(\omega,x)\right)$ for the following deterministic system (1.11) are distinguished. Firstly, we establish the global existence and uniqueness of perturbed solutions around the steady state for the Euler-Poisson equations. Subsequently, we demonstrate the existence of stationary solutions and invariant measure based on the a priori energy estimates and weighted energy estimates.
We recall the steady state and recount the basic conclusion on the existence and uniqueness of $\left(\bar{\rho}(\omega,x),\bar{{\bf u}}(\omega,x),\bar{\Phi}(\omega,x)\right)$. Within the probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$, the steady state $\left(\bar{\rho}(\omega,x),\bar{{\bf u}}(\omega,x),\bar{\Phi}(\omega,x)\right)$ are assumed to adhere to the following equations
$\left\{\begin{array}[]{l}\nabla\cdot\left(\bar{\rho}\bar{{\bf u}}\right)=0,\\ \nabla\cdot\left(\bar{\rho}\bar{{\bf u}}\otimes\bar{{\bf u}}\right)+\nabla P\left(\bar{\rho}\right)-\bar{\rho}\nabla\bar{\Phi}=-\frac{\bar{\rho}\bar{{\bf u}}}{\tau},\\ \triangle\bar{\Phi}=\bar{\rho}-b(x).\end{array}\right.$ (1.11)
[…]
$\nabla\cdot\left(\bar{\rho}\bar{{\bf u}}\cdot\nabla\bar{{\bf u}}\right)+\triangle P\left(\bar{\rho}\right)-\nabla\cdot\left(\bar{\rho}\nabla\bar{\Phi}\right)=0.$ (1.12)
If $\bar{{\bf u}}=0$, it deduces to
$P^{\prime}\left(\bar{\rho}\right)\triangle\bar{\rho}+P^{\prime\prime}\left(\bar{\rho}\right)\left|\nabla\bar{\rho}\right|^{2}-\nabla\bar{\rho}\nabla\bar{\Phi}-\bar{\rho}\left(\bar{\rho}-b\right)=0,$ (1.13)
where $P^{\prime}\left(\bar{\rho}\right)>0=\left|\bar{{\bf u}}\right|^{2}$ so that the equation of $\bar{\rho}$ given in (1.13), is uniformly elliptic. In this paper, we consider the subsonic case, i.e., the condition $P^{\prime}\left(\rho\right)>\left|{\bf u}\right|^{2}$ holds under consideration. For every $\omega\in\Omega$, $\left(\bar{\rho}(\omega,x),0,\bar{\Phi}(\omega,x)\right)=\left(\bar{\rho}(x),0,\bar{\Phi}(x)\right)$ is the unique solution of (1.11), which is called steady state in this paper. We will denote $\left(\bar{\rho}(\omega,x),0,\bar{\Phi}(\omega,x)\right)$ by $\left(\bar{\rho},0,\bar{\Phi}\right)$ for convenience in the following. The law of steady state is Dirac measure $\delta_{\bar{\rho}}\times\delta_{0}\times\delta_{\bar{\Phi}}$, see Appendix 5. We conclude the following lemma for steady state. Here $\bar{U}$ denotes the closed set of $U$.
Proposition 1.1.
Let $b(x)>0$ in $\bar{U}$ and $P:\left(0,\infty\right)\rightarrow\left(0,\infty\right)$ be smooth with $P\left(0\right)=0$. Then there exists $\left(\bar{\rho},\bar{{\bf u}},\bar{\Phi}\right)$, $\forall\omega\in\Omega$, a unique smooth steady-state solution of the insulating problem with the Neumann boundary condition
$\frac{\partial\bar{\Phi}}{\partial\nu}|_{\partial U}\equiv 0,$ (1.14)
such that there hold
$\bar{\rho}>\underline{\rho}>0,\quad\left|\nabla\bar{\rho}\right|>0,\quad\bar{\Phi}>0,\quad\forall x\in\bar{U},\quad\mathbb{P}~{}{\rm a.s.},$ (1.15)
where $\underline{\rho}$ is a constant, and
$\int_{U}\bar{\rho}\operatorname{d}x=\int_{U}b(x)\operatorname{d}x,\quad\mathbb{P}~{}{\rm a.s.}$ (1.16)
Let $Q\left(\rho\right)$ be such that $\nabla Q\left(\rho\right)=\nabla\Phi$ (cf. [12]). Then, the steady state satisfies
$\nabla\bar{Q}(\bar{\rho})=\nabla\bar{\Phi},\quad\triangle\bar{\Phi}=\bar{\rho}-b(x).$ (1.17)
We consider the solutions $\left(\rho,{\bf u},\Phi\right)$ of hydrodynamic system around the steady state $\left(\bar{\rho},0,\bar{\Phi}\right)$ and we denote
$\sigma=\rho-\bar{\rho},\quad\phi=\Phi-\bar{\Phi}.$ (1.18)
Our main result is on the existence of strong solutions near the steady state, and asymptotic stability for insulating boundary condition.
[…]
Theorem 1.1.
Let $U$ be a smooth bounded domain in $\mathds{R}^{3}$ and the pressure $P:(0,\infty)\rightarrow(0,\infty)$ be a smooth function, with $P(\cdot)>0$ and $P^{\prime}(\cdot)>0$. Let $\left(\bar{\rho},0,\bar{\Phi}\right)$ be the smooth steady state in Proposition 1.1. There exists a constant $\varepsilon>0$ such that if the initial condition $\left(\rho_{0},{\bf u}_{0},\Phi_{0}\right)$ satisfies
$\mathbb{E}\left[\left(\left\|\rho_{0}-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}_{0}\right\|_{3}^{2}+\left\|\nabla\Phi_{0}-\nabla\bar{\Phi}\right\|^{2}\right)^{m}\right]\leqslant\varepsilon^{2m},\quad\forall m\geqslant 2,$ (1.19)
and
$\triangle\Phi_{0}=\rho_{0}-b(x),$ (1.20)
then in $\left(\Omega,\mathcal{F},\mathbb{P}\right)$, there exists a unique global-in-time solution $(\rho,{\bf u},\Phi)$ of (1.1):
$\rho,\ {\bf u}\in L^{2m}\left(\Omega;C\left([0,T];H^{3}\left(U\right)\right)\right),\ \Phi\in L^{2m}\left(\Omega;C\left([0,T];H^{5}\left(U\right)\right)\right),$ (1.21)
up to a modification. Moreover, there hold the decay rate and the existence of invariant measure:
  1. (1)
    there are positive constants $C$ and $\alpha$ such that the expectation
    $\mathbb{E}\left[\left(\sup\limits_{s\in[0,t]}\left(\left\|\rho-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}\right\|_{3}^{2}+\left\|\nabla\Phi-\nabla\bar{\Phi}\right\|^{2}\right)\right)^{m}\right]$
    $\leqslant$ $Ce^{-\alpha mt}\mathbb{E}\left[\left(\left\|\rho_{0}-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}_{0}\right\|_{3}^{2}+\left\|\nabla\Phi_{0}-\nabla\bar{\Phi}\right\|^{2}\right)^{m}\right],$ (1.22)
    holds, where $C$ is independent on $t$ and $C$ is the $m$-th power of some constant;
  2. (2)
    the invariant measure generated by $\frac{1}{T}\int_{0}^{T}\mathcal{L}\left(\rho\right)\times\mathcal{L}\left({\bf u}\right)\times\mathcal{L}\left(\Phi\right)\operatorname{d}t$ is exactly the Dirac measure of steady state $\left(\bar{\rho},0,\bar{\Phi}\right)$.
Remark 1.1.
After passing to the limit $t\rightarrow\infty$ in (1), the stationary solution coincides with the steady state $\mathbb{P}$ a.s., since the $m$-th moment of their difference tends to zero.
Remark 1.2.
If for any $\omega\in\Omega$,
$\left(\left\|\rho_{0}-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}_{0}\right\|_{3}^{2}+\left\|\nabla\Phi_{0}-\nabla\bar{\Phi}\right\|^{2}\right)\leqslant\varepsilon^{2},$ (1.23)
then there exists some constant $\tilde{C}$, such that the asymptotic stability holds $\mathbb{P}$ a.s.:
$\sup\limits_{s\in[0,t]}\left(\left\|\rho-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}\right\|_{3}^{2}+\left\|\nabla\Phi-\nabla\bar{\Phi}\right\|^{2}\right)\leqslant 2\tilde{C}e^{-\alpha t}\varepsilon^{2}.$ (1.24)
Actually, by Chebyshev’s inequality (see Appendix 5), it holds
[$\mathbb{P}\left[\left\{\omega\in\Omega|\left\|\rho-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}\right\|_{3}^{2}+\left\|\nabla\Phi-\nabla\bar{\Phi}\right\|^{2}>2\tilde{C}e^{-\alpha t}\varepsilon^{2}\right\}\right]$
$\leqslant$ $\frac{\mathbb{E}\left[\left|\left\|\rho-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}\right\|_{3}^{2}+\left\|\nabla\Phi-\nabla\bar{\Phi}\right\|^{2}\right|^{m}\right]}{\left(2\tilde{C}e^{-\alpha t}\varepsilon^{2}\right)^{m}}$ (1.25)
$\leqslant$ $\frac{\mathbb{E}\left[\tilde{C}e^{-\alpha t}\left(\left\|\rho_{0}-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}_{0}\right\|_{3}^{2}+\left\|\nabla\Phi_{0}-\nabla\bar{\Phi}\right\|^{2}\right)^{m}\right]}{\left(2\tilde{C}e^{-\alpha t}\varepsilon^{2}\right)^{m}}=\frac{1}{2^{m}}.$
Let $m\rightarrow\infty$, then it holds
[$\mathbb{P}\left[\left\{\omega\in\Omega|\left\|\rho-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}\right\|_{3}^{2}+\left\|\nabla\Phi-\nabla\bar{\Phi}\right\|^{2}>2\tilde{C}e^{-\alpha t}\varepsilon^{2}\right\}\right]\rightarrow 0,$
i.e.,
$\left\|\rho-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}\right\|_{3}^{2}+\left\|\nabla\Phi-\nabla\bar{\Phi}\right\|^{2}\leqslant 2\tilde{C}e^{-\alpha t}\varepsilon^{2}$ holds $\mathbb{P}$
a.s. for every $s\in[0,t]$.
Remark 1.3.
The argument in this paper implies the same existence and asymptotic stability of solutions around the steady state for the 2-D system with insulating boundary conditions. Repeating the argument, by Sobolev’s embedding, the existence of perturbed solutions and asymptotic stability of steady state for 1-D system with insulating boundary conditions holds: $\rho$ and ${\bf u}$ are in $L^{2m}\left(\Omega;C\left([0,T];H^{2}\left(U\right)\right)\right)$, $\Phi\in L^{2m}\left(\Omega;C\left([0,T];H^{4}\left(U\right)\right)\right)$ in $\left(\Omega,\mathcal{F},\mathbb{P}\right)$.
As mentioned before, the study of stochastic Euler-Poisson equations, totally from the existing studies for the deterministic case, is new and challenging. The idea of the proof is as follows. We first prove the local existence by Banach’s fixed point theorem, then we establish the uniform energy estimates in time $t$ to show the global existence of $\left(\sigma,{\bf u},\phi\right)$. Furthermore, we prove the weighted energy estimates so that we can obtain the asymptotic stability for steady states with the insulating boundary conditions. The a priori estimates imply the tightness of approximates measures, which will converge to an invariant measure by Krylov-Bogoliubov’s theorem in a complete probability space. From the weighted energy estimates, we then prove that the invariant measure for (1.1) is exactly the law of steady state, c.f. Section 4. This intricate relationship has not been uncovered in the asymptotic behavior analysis of stochastic Navier-Stokes equations yet [6].
Here we explain in detail the main difficulties we have to face to and the strategies we are going to propose.
  1. (1)
    No temporal solutions due to the stochastic term. Since Brownian motion is at most Höder-$\frac{1}{2}-$ continuous with respect to $t$ and it is nowhere differentiable, we do not have $\frac{\operatorname{d}W}{\operatorname{d}t}$ or $\frac{\operatorname{d}\left(\rho{\bf u}\right)}{\operatorname{d}t}$ either. No temporal derivative is involved in the norm of solutions. Thus, in deterministic cases [12, 26], the spatial estimates bounded by the temporal derivatives estimates like
    $\qquad\|\left(\rho-\bar{\rho}\right)\|^{2}+\|\nabla\left(\rho-\bar{\rho}\right)\|^{2}+\|\nabla\cdot{\bf u}\|^{2}\leqslant C\left(\left\|u_{t}\right\|^{2}+\left\|\left(\rho-\bar{\rho}\right)_{t}\right\|^{2}+\|u\|^{2}+\interleave w\interleave^{3}\right),$ (1.26)
[…]
    It is interesting that the noise in form of (1.4) is in the higher order of ${\bf u}$ than the Lipschitz continuous on ${\bf u}$. This is reasonable when we consider the small perturbation around the steady state, which is different with most cases in which Lipshcitz continuous coefficients give birth to wellposedness. In this case the influence of stochastic force does not been exaggerated so much.
  2. (2)
    Weighted energy estimates on account of the estimates of the stochastic integral. Recalling the 3-D deterministic case, for instance, one can multiply the ordinary differential equation of solutions with the exponential function of $t$ directly to facilitate the stability analysis. However, in this paper, in order to estimates the stochastic term, we apply the Burkholder-Davis-Gundy’s inequality to the stochastic integral of the Wiener process. Then the a priori estimates (2.218) is already in the form of time integrals rather than a differential equation. Integration with respect to time twice could not imply the asymptotic stability. Consequently, direct acquisition of asymptotic stability becomes challenging. To overcome this obstacle, we employ the weighted energy estimates. Moreover, the weight is determined by the a priori estimates which should be obtained first, cf. Section 3.1.
This paper is organized as follows. Section 2 is dedicated to establishing the global existence of solutions around the steady state. In Section 3, we investigate the asymptotic stability of semiconductor equations. Finally, in Section 4, we demonstrate the existence and property of invariant measures. Section 5 is the Appendix, in which we provide an overview of stochastic analysis theories that are employed in this study.
2. Global existence of solutions
In this section, we first establish the local existence of solutions by Banach’s fixed point theorem. Then under the assumption that $\left\|\rho-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}\right\|_{3}^{2}+\left\|\nabla\Phi-\nabla\bar{\Phi}\right\|^{2}$ is small, we get the uniform estimates so as to obtain the global existence of $\rho,{\bf u}\in L^{2m}\left(\Omega;C\left([0,T];H^{3}\left(U\right)\right)\right)$.
In form of $\left(\sigma,{\bf u},\phi\right)=\left(\rho-\bar{\rho},{\bf u}-\bar{{\bf u}},\Phi-\bar{\Phi}\right)$, the hydrodynamic system deforms into
$\left\{\begin{array}[]{l}\sigma_{t}+\nabla\cdot\left(\left(\bar{\rho}+\sigma\right){\bf u}\right)=0,\\ \operatorname{d}_{t}\left({\bf u}\right)+\left(\left({\bf u}\cdot\nabla\right){\bf u}+{\bf u}+\nabla Q\left(\bar{\rho}+\sigma\right)-\nabla Q\left(\bar{\rho}\right)\right)\operatorname{d}t=\nabla\phi\operatorname{d}t+\frac{\mathbb{F}}{\bar{\rho}+\sigma}\operatorname{d}_{t}W,\\ \triangle\phi=\sigma.\end{array}\right.$ (2.1)
Here we view $\tau$ as a constant without loss of generality for the stability analysis. In terms of component, by Taylor’s expansion, it holds
$\nabla Q\left(\bar{\rho}+\sigma\right)-\nabla Q\left(\bar{\rho}\right)=\left(Q\left(\bar{\rho}+\sigma\right)-Q\left(\bar{\rho}\right)\right)_{,i}$
$=$ $Q^{\prime}\left(\bar{\rho}+\sigma\right)\left(\bar{\rho}+\sigma\right)_{,i}-Q^{\prime}\left(\bar{\rho}\right)\bar{\rho}_{,i}$
$=$ $Q^{\prime}\left(\bar{\rho}+\sigma\right)\sigma_{,i}+\left(Q^{\prime}\left(\bar{\rho}+\sigma\right)-Q^{\prime}\left(\bar{\rho}\right)\right)\bar{\rho}_{,i}$ (2.2)
$=$ $Q^{\prime}\left(\bar{\rho}+\sigma\right)\sigma_{,i}+Q^{\prime\prime}\left(\bar{\rho}\right)\sigma\bar{\rho}_{i}+\left(Q^{\prime}\left(\bar{\rho}+\sigma\right)-Q^{\prime}\left(\bar{\rho}\right)-Q^{\prime\prime}\left(\bar{\rho}\right)\sigma\right)\bar{\rho}_{,i}$
::=:= $Q^{\prime}\left(\bar{\rho}+\sigma\right)\sigma_{,i}+Q^{\prime\prime}\left(\bar{\rho}\right)\sigma\bar{\rho}_{,i}+h_{i},$
where $\left(\cdot\right)_{,i}$ means the derivative with respect to $x_{i}$, and
$h_{i}=O\left(\sigma^{2}\right).$ (2.3)
In term of component, there holds
$\phi_{,i}=\triangle^{-1}\sigma_{,i},$ (2.4)
where $\triangle^{-1}$ is well-defined under the condition (1.6). In matrix notation, denoting $\mathbf{w}=\left[\begin{array}[]{l}\sigma\\ {\bf u}\end{array}\right]$, we write the system as
$\operatorname{d}_{t}\mathbf{w}+\left(\mathcal{A}^{1}\mathbf{w}_{,1}+\mathcal{A}^{2}\mathbf{w}_{,2}+\mathcal{A}^{3}\mathbf{w}_{,3}+\mathcal{B}\mathbf{w}+\mathcal{L}_{{\bf u}}\right)\operatorname{d}t=\mathcal{L}_{\mathcal{\phi}}\operatorname{d}t+f,$ (2.5)
where
$\mathcal{A}^{1}=\left[\begin{array}[]{cccc}u^{1}&\bar{\rho}+\sigma&0&0\\ Q^{\prime}\left(\bar{\rho}+\sigma\right)&u^{1}&0&0\\ 0&0&u^{1}&0\\ 0&0&0&u^{1}\end{array}\right],$ (2.10)
$\mathcal{A}^{2}=\left[\begin{array}[]{cccc}u^{2}&0&\bar{\rho}+\sigma&0\\ 0&u^{2}&0&0\\ Q^{\prime}\left(\bar{\rho}+\sigma\right)&0&u^{2}&0\\ 0&0&0&u^{2}\end{array}\right],$ (2.15)
$\mathcal{A}^{3}=\left[\begin{array}[]{cccc}u^{3}&0&0&\bar{\rho}+\sigma\\ 0&u^{3}&0&0\\ 0&0&u^{3}&0\\ Q^{\prime}\left(\bar{\rho}+\sigma\right)&0&0&u^{3}\end{array}\right],$ (2.20)
$\mathcal{B}=\left[\begin{array}[]{cccc}0&\bar{\rho}_{,1}&\bar{\rho}_{,2}&\bar{\rho}_{,3}\\ Q^{\prime\prime}\left(\bar{\rho}\right)\bar{\rho}_{,1}&0&0&0\\ Q^{\prime\prime}\left(\bar{\rho}\right)\bar{\rho}_{,2}&0&0&0\\ Q^{\prime\prime}\left(\bar{\rho}\right)\bar{\rho}_{,3}&0&0&0\end{array}\right],$ (2.25)
$\mathcal{L}_{{\bf u}}=\left[\begin{array}[]{c}0\\ u^{1}\\ u^{2}\\ u^{3}\end{array}\right],\quad\mathcal{L}_{\mathcal{\phi}}=\left[\begin{array}[]{c}0\\ \triangle^{-1}\sigma_{,1}\\ \triangle^{-1}\sigma_{,2}\\ \triangle^{-1}\sigma_{,3}\end{array}\right],\quad f=-\left[\begin{array}[]{c}0\\ h\left(\sigma\right)_{,1}-\mathbb{F}^{1}\operatorname{d}_{t}W\\ h\left(\sigma\right)_{,2}-\mathbb{F}^{2}\operatorname{d}_{t}W\\ h\left(\sigma\right)_{,3}-\mathbb{F}^{3}\operatorname{d}_{t}W\end{array}\right].$ (2.38)
We define the symmetrizer $\mathcal{D}=\operatorname{diag}\left[Q^{\prime}\left(\bar{\rho}+\sigma\right),\bar{\rho}+\sigma,\bar{\rho}+\sigma,\bar{\rho}+\sigma\right]:=\operatorname{diag}\left[d_{1},d_{2},d_{3},d_{4}\right]$, then the system deforms into
$\mathcal{D}\operatorname{d}_{t}\mathbf{w}+\left(\tilde{\mathcal{A}}^{1}\mathbf{w}_{,1}+\tilde{\mathcal{A}}^{2}\mathbf{w}_{,2}+\tilde{\mathcal{A}}^{3}\mathbf{w}_{,3}+\tilde{\mathcal{B}}\mathbf{w}+\tilde{\mathcal{L}}_{{\bf u}}\right)\operatorname{d}t=\tilde{\mathcal{L}}_{\phi}\operatorname{d}t+\tilde{f},$ (2.39)
where
$\tilde{\mathcal{A}}^{1}=\left[\begin{array}[]{cccc}u^{1}Q^{\prime}\left(\bar{\rho}+\sigma\right)&\left(\bar{\rho}+\sigma\right)Q^{\prime}\left(\bar{\rho}+\sigma\right)&0&0\\ \left(\bar{\rho}+\sigma\right)Q^{\prime}\left(\bar{\rho}+\sigma\right)&\left(\bar{\rho}+\sigma\right)u^{1}&0&0\\ 0&0&\left(\bar{\rho}+\sigma\right)u^{1}&0\\ 0&0&0&\left(\bar{\rho}+\sigma\right)u^{1}\end{array}\right],$ (2.44)
$\tilde{\mathcal{A}}^{2}=\left[\begin{array}[]{cccc}u^{2}Q^{\prime}\left(\bar{\rho}+\sigma\right)&0&\left(\bar{\rho}+\sigma\right)Q^{\prime}\left(\bar{\rho}+\sigma\right)&0\\ 0&\left(\bar{\rho}+\sigma\right)u^{2}&0&0\\ \left(\bar{\rho}+\sigma\right)Q^{\prime}\left(\bar{\rho}+\sigma\right)&0&\left(\bar{\rho}+\sigma\right)u^{2}&0\\ 0&0&0&\left(\bar{\rho}+\sigma\right)u^{2}\end{array}\right],$ (2.49)
$\tilde{\mathcal{A}}^{3}=\left[\begin{array}[]{cccc}u^{3}Q^{\prime}\left(\bar{\rho}+\sigma\right)&0&0&\left(\bar{\rho}+\sigma\right)Q^{\prime}\left(\bar{\rho}+\sigma\right)\\ 0&\left(\bar{\rho}+\sigma\right)u^{3}&0&0\\ 0&0&\left(\bar{\rho}+\sigma\right)u^{3}&0\\ \left(\bar{\rho}+\sigma\right)Q^{\prime}\left(\bar{\rho}+\sigma\right)&0&0&\left(\bar{\rho}+\sigma\right)u^{3}\end{array}\right],$ (2.54)
~$\tilde{\mathcal{B}}=\left[\begin{array}[]{cccc}0&\bar{\rho}_{,1}&\bar{\rho}_{,2}&\bar{\rho}_{,3}\\ Q^{\prime\prime}\left(\bar{\rho}\right)\bar{\rho}_{,1}&0&0&0\\ Q^{\prime\prime}\left(\bar{\rho}\right)\bar{\rho}_{,2}&0&0&0\\ Q^{\prime\prime}\left(\bar{\rho}\right)\bar{\rho}_{,3}&0&0&0\end{array}\right],$ (2.59)
~$\tilde{\mathcal{L}}_{{\bf u}}=\left(\bar{\rho}+\sigma\right)\left[\begin{array}[]{c}0\\ u^{1}\\ u^{2}\\ u^{3}\end{array}\right],\quad\tilde{\mathcal{L}}_{\phi}=\left(\bar{\rho}+\sigma\right)\left[\begin{array}[]{c}0\\ \triangle^{-1}\sigma_{,1}\\ \triangle^{-1}\sigma_{,2}\\ \triangle^{-1}\sigma_{,3}\end{array}\right],$ (2.68)
$\tilde{f}=-\left(\bar{\rho}+\sigma\right)\left[\begin{array}[]{c}0\\ h\left(\sigma\right)_{,1}\operatorname{d}t-\mathbb{F}^{1}\operatorname{d}_{t}W\\ h\left(\sigma\right)_{,2}\operatorname{d}t-\mathbb{F}^{2}\operatorname{d}_{t}W\\ h\left(\sigma\right)_{,3}\operatorname{d}t-\mathbb{F}^{3}\operatorname{d}_{t}W\end{array}\right].$ (2.73)
2.1. Local existence
Similarly as in [21, 28], we first linearize the system, then we use Banach’s fixed point theorem to get the local existence by the a priori energy estimates.
The linearized system is
$\mathcal{D}(\hat{\sigma})\operatorname{d}_{t}\mathbf{w}+\left(\tilde{\mathcal{A}}^{1}\left(\hat{\mathbf{w}}\right)\mathbf{w}_{,1}+\tilde{\mathcal{A}}^{2}(\hat{\mathbf{w}})\mathbf{w}_{,2}+\tilde{\mathcal{A}}^{3}(\hat{\mathbf{w}})\mathbf{w}_{,3}+\tilde{\mathcal{B}}\mathbf{w}\right)\operatorname{d}t$ (2.74)
$=$ ~$-\tilde{\mathcal{L}}_{\hat{{\bf u}}}\left(\hat{\sigma},{\bf u}\right)\operatorname{d}t+\tilde{\mathcal{L}}_{\hat{\phi}}\left(\hat{\sigma},\hat{\phi}\right)\operatorname{d}t+\tilde{f}(\hat{\mathbf{w}}).$
where $\hat{\mathbf{w}}=\left[\begin{array}[]{l}\hat{\sigma}\\ \hat{{\bf u}}\end{array}\right]$ is given, $\hat{\sigma}\in C\left([0,T];H^{3}\left(U\right)\right)$, $\hat{{\bf u}}\in C\left([0,T];H^{3}\left(U\right)\right)$, and we denote $M=\sup\limits_{t\in[0,T]}\left\|\hat{\sigma},\hat{{\bf u}}\right\|_{3}$.
Step 1: Uniform bound.
By Itô’s formula (see Appendix 5), it holds
$\int_{U}\operatorname{d}_{t}\left(\frac{1}{2}\mathcal{D}\mathbf{w}\cdot\mathbf{w}\right)\operatorname{d}x=\int_{U}\frac{1}{2}\operatorname{d}_{t}\mathcal{D}\mathbf{w}\cdot\mathbf{w}\operatorname{d}x+\int_{U}\mathcal{D}\mathbf{w}\cdot\operatorname{d}_{t}\mathbf{w}\operatorname{d}x+\int_{U}\mathcal{D}\mathbb{F}\cdot\mathbb{F}\operatorname{d}x\operatorname{d}t.$ (2.75)
We integrate
$\mathcal{D}\operatorname{d}_{t}\mathbf{w}\cdot\mathbf{w}+\left(\tilde{\mathcal{A}}^{1}\left(\hat{\mathbf{w}}\right)\mathbf{w}_{,1}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{2}\left(\hat{\mathbf{w}}\right)\mathbf{w}_{,2}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{3}\left(\hat{\mathbf{w}}\right)\mathbf{w}_{,3}\cdot\mathbf{w}+\tilde{\mathcal{B}}\mathbf{w}\cdot\mathbf{w}\right)\operatorname{d}t$ (2.76)
$=$ ~$-\tilde{\mathcal{L}}_{\hat{{\bf u}}}\cdot\mathbf{w}\operatorname{d}t+\tilde{\mathcal{L}}_{\hat{\phi}}\cdot\mathbf{w}\operatorname{d}t+\tilde{f}\cdot\mathbf{w}$
over the domain $U$, we gain
$\int_{U}\mathcal{D}\mathbf{w}\cdot\operatorname{d}_{t}\mathbf{w}\operatorname{d}x$
$=$ $\int_{U}\left(-\left(\tilde{\mathcal{A}}^{1}\mathbf{w}_{,1}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{2}\mathbf{w}_{,2}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{3}\mathbf{w}_{,3}\cdot\mathbf{w}+\tilde{\mathcal{B}}\mathbf{w}\cdot\mathbf{w}\right)-\tilde{\mathcal{L}}_{\hat{{\bf u}}}\cdot\mathbf{w}+\tilde{\mathcal{L}}_{\hat{\phi}}\cdot\mathbf{w}\right)\operatorname{d}x\operatorname{d}t$ (2.77)
$+\int_{U}\nabla h\left(\hat{\sigma}\right)\cdot\mathbf{w}\operatorname{d}x\operatorname{d}t+\int_{U}\mathcal{D}\mathbb{F}\operatorname{d}_{t}W\cdot\mathbf{w}\operatorname{d}x.$
By the integration by parts, we have
$\int_{U}\left(\tilde{\mathcal{A}}^{1}\mathbf{w}_{,1}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{2}\mathbf{w}_{,2}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{3}\mathbf{w}_{,3}\cdot\mathbf{w}+\tilde{\mathcal{B}}\mathbf{w}\cdot\mathbf{w}\right)\operatorname{d}x\operatorname{d}t$
$=$ $\int_{U}\left(-\frac{1}{2}\left(\mathbf{w}\tilde{\mathcal{A}}^{1}_{1}\mathbf{w}+\mathbf{w}\tilde{\mathcal{A}}^{2}_{2}\mathbf{w}+\mathbf{w}\tilde{\mathcal{A}}^{3}_{3}\mathbf{w}\right)+\tilde{\mathcal{B}}\left|\mathbf{w}\right|^{2}\right)\operatorname{d}x\operatorname{d}t+\int_{U}\left(\mathbf{w}\tilde{\mathcal{A}}^{j}\mathbf{w}\right)_{,j}\operatorname{d}x\operatorname{d}t.$ (2.78)
On account of the insulated boundary condition $\hat{{\bf u}}\cdot\nu|_{\partial U}=0$, it holds
$\int_{U}\left(\mathbf{w}\tilde{\mathcal{A}}^{j}\mathbf{w}\right)_{,j}\operatorname{d}x$ (2.79)
$=$ $\int_{\partial U}\left(\left(\hat{{\bf u}}\cdot\nu\right)\left(Q^{\prime}\left(\bar{\rho}+\hat{\sigma}\right)\sigma^{2}+\left(\bar{\rho}+\hat{\sigma}\right)\left|{\bf u}\right|^{2}+2\hat{\rho}{\bf u}Q^{\prime}\left(\bar{\rho}+\hat{\sigma}\right)\left(\bar{\rho}+\hat{\sigma}\right)\right)\right)\operatorname{d}S\equiv 0.$
In summary, there holds
$\int_{U}\operatorname{d}_{t}\left(\frac{1}{2}\mathcal{D}\mathbf{w}\cdot\mathbf{w}\right)\operatorname{d}x$
$=$ $\int_{U}\frac{1}{2}\operatorname{d}_{t}\mathcal{D}\mathbf{w}\cdot\mathbf{w}\operatorname{d}x-\int_{U}\left(-\frac{1}{2}\left(\mathbf{w}\tilde{\mathcal{A}}^{1}_{1}\mathbf{w}+\mathbf{w}\tilde{\mathcal{A}}^{2}_{2}\mathbf{w}+\mathbf{w}\tilde{\mathcal{A}}^{3}_{3}\mathbf{w}\right)+\tilde{\mathcal{B}}\left|\mathbf{w}\right|^{2}\right)\operatorname{d}x\operatorname{d}t$ (2.80)
$+\int_{U}\left(-\tilde{\mathcal{L}}_{\hat{{\bf u}}}\cdot\mathbf{w}+\tilde{\mathcal{L}}_{\hat{\phi}}\cdot\mathbf{w}\right)\operatorname{d}x\operatorname{d}t+\int_{U}\nabla h\left(\hat{\sigma}\right)\cdot\mathbf{w}\operatorname{d}x\operatorname{d}t+\int_{U}\mathcal{D}\mathbb{F}\operatorname{d}_{t}W\cdot\mathbf{w}\operatorname{d}x$
$+\int_{U}\mathcal{D}\mathbb{F}\cdot\mathbb{F}\operatorname{d}x\operatorname{d}t.$
Direct calculation shows that
$-\frac{1}{2}\tilde{\mathcal{A}}^{i}_{,i}+\tilde{\mathcal{B}}-\operatorname{diag}\left[-\frac{1}{2}\left(u^{i}Q^{\prime}\left(\rho\right)\right)_{,i},-\frac{1}{2}\left(u^{i}\rho\right)_{,i},-\frac{1}{2}\left(u^{i}\rho\right)_{,i},-\frac{1}{2}\left(u^{i}\rho\right)_{,i}\right]$
[…]