### v3

# Stationary solution to Stochastically Forced Euler-Poisson Equations in Bounded Domain: Part 1. 3-D Insulating Boundary

Abstract.
This paper is concerned with $3$-D stochastic Euler-Poisson equations with insulating boundary conditions forced by the Wiener process. We first establish the global existence and uniqueness of the solution to the system, then we prove that the solution converges to its steady-state time-asymptotically. To obtain the converging rate, we need to develop weighted energy estimates, which are not required for the deterministic counterpart of the problem. Moreover, we observe that the invariant measure is just the Dirac measure generated by the steady-state, in which the time-exponential convergence rate to the steady-state plays an essential role.
Keywords: stability, stationary solution, stochastic Euler-Poisson equations, cylindrical Brownian motion, insulating boundary conditions
2020 AMS Subject Classification: 34D05, 34D20,35B35, 35K51, 60H15, 82D37.
1. Introduction
Euler-Poisson equations is important in the analysis and design of semiconductor devices, offering a more precise description of physical phenomena [21] compared to the conventional drift-diffusion model. Furthermore, in the extreme ultraviolet (EUV) lithography, stochastic effects sometimes cause unwanted defects and pattern roughness in chips [3], that may impact the performance of a chip, or cause a device to fail. Hence, there is a pressing need to investigate the dynamic model of semiconductors perturbed by stochastic forces within mathematical frameworks. The stochastically forced Euler-Poisson equations (SEP for short) in a bounded smooth domain $U\subset\mathds{R}^{3}$ reads as
$\left\{\begin{array}[]{l}\rho_{t}+\nabla\cdot\left(\rho{\bf u}\right)=0,\\ \operatorname{d}\left(\rho{\bf u}\right)+\left(\nabla\cdot\left(\rho{\bf u}\otimes{\bf u}\right)+\nabla P\left(\rho\right)-\rho\nabla\Phi\right)\operatorname{d}t=-\frac{\rho{\bf u}}{\tau}\operatorname{d}t+\mathbb{F}\left(\rho,\mathbf{u}\right)\operatorname{d}W,\\ \triangle\Phi=\rho-b,\end{array}\right.$ (1.1)
where ``$\operatorname{d}"$ in (1.1) is the differential notation with respect to time $t$, in comparison to gradient $\nabla$ and Laplacian $\triangle$ for spatial derivatives, $\rho$ is the electron density of semiconductors, ${\bf u}$ denotes the particle velocity. $P\left(\rho\right)$ is the pressure, $\Phi$ is the electrostatic potential, $\tau$ is the velocity relaxation time and $b(x)$ is called the doping profile, which is positive and immobile. The above mentioned unknowns $\rho=\rho(\omega,t,x)$, ${\bf u}={\bf u}(\omega,t,x)$, $\Phi=\Phi(\omega,t,x)$, and $P\left(\rho\right)=P\left(\rho(\omega,t,x)\right)$ are stochastic processes as functions with respect to $\omega$, $t$, and $x$, where $\omega$ is a sample in the complete probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$. For convenience, we use the simplified notions $\rho$, ${\bf u}$, $\Phi$, and $P\left(\rho\right)$ here and hereafter. $W$ is an $\mathcal{H}$-valued cylindrical Brownian motion defined on the filtrated probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$, where $\mathcal{H}$ is an auxiliary separable Hilbert space, $\mathcal{F}$ is the filtration, see the definitions of filtration and Wiener process in Appendix 5.
[…]
$\langle\mathbb{F}\left(\rho,\mathbf{u}\right),e_{k}\rangle=\mathbf{F}_{k}\left(\rho,{\bf u}\right)$ (1.2)
is an $H$-valued vector function, which shows the strength of the external stochastic forces by
$\mathbb{F}\left(\rho,\mathbf{u}\right)\operatorname{d}W=\sum\limits_{k=1}^{+\infty}\mathbf{F}_{k}\left(\rho,{\bf u}\right)\operatorname{d}\beta_{k}e_{k}.$ (1.3)
Throughout the paper, we assume that
$\mathbf{F}_{k}\left(\rho,{\bf u}\right)=a_{k}\rho{\bf u}Y\left(\rho,{\bf u}\right),$ (1.4)
where ${a_{k}}$ are positive constants, $Y\left(\rho,{\bf u}\right)$ is a smooth function of $\rho$ and ${\bf u}$, and can be bounded by the homogeneous polynomials.
Subjected to the stochastic Euler-Poisson equations (1.1), the proposed boundary is the insulating boundary:
${\bf u}\cdot\nu=0,\quad\nabla\Phi\cdot\nu=0,$ (1.5)
where $\nu$ is the outer normal vector of $U$; and the initial data is:
$(\rho,{\bf u}_{0},\Phi)|_{t=0}=\left(\rho_{0}\left(\omega,x\right),~{}{\bf u}_{0}\left(\omega,x\right),~{}\Phi_{0}\left(\omega,x\right)\right),$ (1.6)
which is given in the probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$, $\rho_{0}\left(\omega,x\right)>0$. Here and hereafter, we simply denote the initial data by $\left(\rho_{0},{\bf u}_{0},\Phi_{0}\right)$ without confusion.
The hydrodynamic model of semiconductors was first introduced by Blotekjaer [4], which is the deterministically dynamical model presented by Euler-Poisson equations mathematically. For 1-D case, the initial-boundary value problems to Euler-Poisson equations with the insulating boundary and the Ohmic contact boundary were studied by Hsiao-Yang [29], Li-Markowich-Mei [39], respectively, where the solutions are showed to converge to the corresponding subsonic steady-states time-asymptotically, where the doping profile is needed to be flat: $|b^{\prime}(x)|\ll 1$. Such a restriction was then released by Nishibata-Suzuki [44] and Guo-Strauss [21] independently. For $N$-D case, Guo-Strauss [21] first considered the deterministic 3-D Euler-Poisson equations in bounded domain with insulating boundary, and showed the convergence of solutions to the 3-D subsonic steady-states. Subsequently, Mei-Wu-Zhang [42] investigated the convergence to the steady-states for the $N$-D radial Euler-Poisson equations with the Ohmic contact boundary. For the whole space without boundary effects, the Cauchy problems to deterministic Euler-Poisson equations were extensively studied in [10, 30, 31, 32, 33, 36]. For the case of free boundary with vacuum, we refer to [40, 43, 51] and the references therein. For the formulation of singularities in compressible Euler-Poisson equations and the large time behavior of Euler equations with damping, one can refer to [49] and [50], respectively.
When the hydrodynamic model of semiconductors is counted into the stochastic affections, it then becomes the stochastic Euler-Poisson equations with uncertain extra disturbances, see (1.1) with the Wiener process $\mathbb{F}\left(\rho,\mathbf{u}\right)\operatorname{d}W$. This is a new model for semiconductor devices and never touched yet. The main issue of the paper is to investigate this 3-D SEP in bounded domain with insulating boundary, and are going to prove the convergence of solutions to the stochastic steady states. The coefficient function of Wiener process $\mathbf{F}_{k}\left(\rho,{\bf u}\right)$, depending on the solutions $\rho$ and ${\bf u}$, is called the multiplicative noise. In most cases, the multiplicative noise magnifies the perturbation and thereby complicating the well-posedness of solutions for evolution systems. The stochastic forces are at most Hölder-$\frac{1}{2}-$continuous in time $t$, resulting in reduced regularity of velocity with respect to time. So from a mathematical standpoint, the study of the stochastic problem helps us to study how the solutions to stochastic Euler-Poisson equations behave in the absence of strong regularity in time. Further, this encourages exploring whether the desirable property remains under the influence of particular types of noise. This is the first attempt to study the asymptotic behavior of solutions to stochastic 3-D Euler-Poisson equations.
For stochastic evolution systems, the solution is called the stationary solution provided that the increment of solutions during evolution is time-independent. Originally, the study of stationary measures dates back to the works of Hopf [28], Doeblin [9], Doob [12], Halmos [22, 23], Feller [13], and Harris and Robbins [25, 24], who contributed to the theory of discrete Markov processes from 1930s to 1950s. The study of invariant measure of fluid models dates back to Cruzeiro [8] for stochastic incompressible Navier-Stokes equations in 1989, by Galerkin approximation with dimensions $D\geqslant 2$. Flandoli [14] proved existence of an invariant measure by the “remote start” method for 2-D incompressible Navier-Stokes equations in 1994. One year later Flandoli-Gatarek [15] showed existence of stationary solution for 3-D incompressible Navier-Stokes equations by a different method with [8]. In 2002, Mattingly [41] proved the existence of exponentially attracting invariant measure with respect to initial data, for incompressible N-S equations. Later, Goldys-Maslowski [20] showed that transition measures of the 2-D stochastic Navier-Stokes equations converge exponentially fast to the corresponding invariant measures in the distance of total variation. Then for 3-D case, Da Prato and Debussche [46] constructed a transition semigroup for 3-D stochastic Navier-Stokes equations without the uniqueness, which allows for rather irregular solutions. Flandoli-Romito [17] used the classical Stroock-Varadhan type argument to find the almost sure Markov selection. The above works are for the incompressible case. For stochastic compressible Navier-Stokes equations, Breit-Feireisl-Hofmanová-Maslowski [6] proved the existence of stationary solutions. Compared to Navier–Stokes equations, the regularity effect of viscosity is lost for Euler system. Hofmanová-Zhu-Zhu [27] selected the dissipative global martingale solutions to the stochastic incompressible Euler system, and obtained the non-uniqueness of strong Markov solutions. Very recently, they [26] showed that stationary solution to the Euler equations is a vanishing viscosities limit in law of stationary analytically weak solutions to Navier-Stokes equations. In terms of the non-uniqueness studies, some scholars believe that a certain stochastic perturbation can provide a regularizing effect of the underlying PDE dynamics. For instance, Flandoli-Luo [16] showed that a noise of transport type prevents a vorticity blow-up in the incompressible Navier-Stokes equations. A linear multiplicative noise prevents the blow up of the velocity with high probability for the 3-D Euler system, which was shown by Glatt-Holtz-Vicol [19]. Gess-Souganidis [18] investigated the large-time behavior and established the existence of an invariant measure for stochastic scalar conservation laws, demonstrating that an algebraic decay rate in time holds. In their work, they introduced a particular type of noise that provided stronger regularization properties for the problem. Then Dong-Zhang-Zhang [11] proved the existence of stationary solutions with the multiplicative noise. For stochastic conservation laws, Da Prato-Gatarek studied the existence and uniqueness of invariant measure for stochastic Burgers equation [47]. Da Prato-Zabczyk listed the basic theory of stationary solutions of general stochastic PDEs in view of invariant measure in book [48]. Bedrossian-Liss [1] gave the existence of stationary measures for stochastic ordinary differential equations with a nonlinear term. To the best of our knowledge, the stationary solutions of SEP have not been explored previously. For our SEP, the electrostatic potential term $\rho\nabla\Phi\operatorname{d}t$ and the relaxation term $\frac{\rho{\bf u}}{\tau}\operatorname{d}t$ are actually damping terms providing better regularity than Euler equations. In this paper, we could show the existence and uniqueness of invariant measure in more regular space.
It is worth noting that the stationary solution we consider is in view of invariant measure. In this paper, the concepts of stationary solution for stochastically forced system (1.1) and steady state $\left(\bar{\rho}(\omega,x),\bar{{\bf u}}(\omega,x),\bar{\Phi}(\omega,x)\right)$ for the following deterministic system (1.10) are distinguished. Firstly, we establish the global existence and uniqueness of perturbed solutions around the steady state for the Euler-Poisson equations. Subsequently, we demonstrate the existence of stationary solutions and invariant measure based on the a priori energy estimates and weighted energy estimates.
We recall the steady state and recount the basic conclusion on the existence and uniqueness of $\left(\bar{\rho}(\omega,x),\bar{{\bf u}}(\omega,x),\bar{\Phi}(\omega,x)\right)$. Within the probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$, the steady state $\left(\bar{\rho}(\omega,x),\bar{{\bf u}}(\omega,x),\bar{\Phi}(\omega,x)\right)$ are assumed to adhere to the following equations
$\left\{\begin{array}[]{l}\nabla\cdot\left(\bar{\rho}\bar{{\bf u}}\right)=0,\\ \nabla\cdot\left(\bar{\rho}\bar{{\bf u}}\otimes\bar{{\bf u}}\right)+\nabla P\left(\bar{\rho}\right)-\bar{\rho}\nabla\bar{\Phi}=-\frac{\bar{\rho}\bar{{\bf u}}}{\tau},\\ \triangle\bar{\Phi}=\bar{\rho}-b(x).\end{array}\right.$ (1.10)
[…]
$\nabla\cdot\left(\bar{\rho}\bar{{\bf u}}\cdot\nabla\bar{{\bf u}}\right)+\triangle P\left(\bar{\rho}\right)-\nabla\cdot\left(\bar{\rho}\nabla\bar{\Phi}\right)=0.$ (1.11)
If $\bar{{\bf u}}=0$, it deduces to
$P^{\prime}\left(\bar{\rho}\right)\triangle\bar{\rho}+P^{\prime\prime}\left(\bar{\rho}\right)\left|\nabla\bar{\rho}\right|^{2}-\nabla\bar{\rho}\nabla\bar{\Phi}-\bar{\rho}\left(\bar{\rho}-b\right)=0,$ (1.12)
where $P^{\prime}\left(\bar{\rho}\right)>0=\left|\bar{{\bf u}}\right|^{2}$ so that the equation of $\bar{\rho}$ given in (1.12), is uniformly elliptic. In this paper, we consider the subsonic case, i.e., the condition $P^{\prime}\left(\rho\right)>\left|{\bf u}\right|^{2}$ holds under consideration. For every $\omega\in\Omega$, $\left(\bar{\rho}(\omega,x),0,\bar{\Phi}(\omega,x)\right)=\left(\bar{\rho}(x),0,\bar{\Phi}(x)\right)$ is the unique solution of (1.10), which is called steady state in this paper. We will denote $\left(\bar{\rho}(\omega,x),0,\bar{\Phi}(\omega,x)\right)$ by $\left(\bar{\rho},0,\bar{\Phi}\right)$ for convenience in the following. The law of steady state is Dirac measure $\delta_{\bar{\rho}}\times\delta_{0}\times\delta_{\bar{\Phi}}$, see Appendix 5. We conclude the following lemma for steady state. Here $\bar{U}$ denotes the closed set of $U$.
Proposition 1.1.
Let $b(x)>0$ in $\bar{U}$ and $P:\left(0,\infty\right)\rightarrow\left(0,\infty\right)$ be smooth with $P\left(0\right)=0$. Then there exists $\left(\bar{\rho},\bar{{\bf u}},\bar{\Phi}\right)$, $\forall\omega\in\Omega$, a unique smooth steady-state solution of the insulating problem with the Neumann boundary condition
$\frac{\partial\bar{\Phi}}{\partial\nu}|_{\partial U}\equiv 0,$ (1.13)
such that there holds
$\bar{\rho}>\underline{\rho}>0,\quad\left|\nabla\bar{\rho}\right|>0,\quad\bar{\Phi}>0,\quad\forall x\in\bar{U},\quad\mathbb{P}~{}{\rm a.s.},$ (1.14)
where $\underline{\rho}$ is a constant, and
$\int_{U}\bar{\rho}\operatorname{d}x=\int_{U}b(x)\operatorname{d}x,\quad\mathbb{P}~{}{\rm a.s.}$ (1.15)
Let $Q\left(\rho\right)$ be such that $\nabla Q\left(\rho\right)=\nabla\Phi$ (cf. [21]). Then, the steady state satisfies
$\nabla\bar{Q}(\bar{\rho})=\nabla\bar{\Phi},\quad\triangle\bar{\Phi}=\bar{\rho}-b(x).$ (1.16)
We consider the solutions $\left(\rho,{\bf u},\Phi\right)$ of hydrodynamic system around the steady state $\left(\bar{\rho},0,\bar{\Phi}\right)$ and we denote
$\sigma=\rho-\bar{\rho},\quad\phi=\Phi-\bar{\Phi}.$ (1.17)
Our main result is on the existence of solutions near the steady state, and asymptotic stability for insulating boundary condition.
[…]
Theorem 1.1 (Global existence).
Let $U$ be a smooth bounded domain in $\mathds{R}^{3}$ and the pressure $P:(0,\infty)\rightarrow(0,\infty)$ be a smooth function, with $P(\cdot)>0$ and $P^{\prime}(\cdot)>0$. Let $\left(\bar{\rho},0,\bar{\Phi}\right)$ be the smooth steady state in Proposition 1.1. and
$\triangle\Phi_{0}=\rho_{0}-b(x),$ (1.18)
then in $\left(\Omega,\mathcal{F},\mathbb{P}\right)$, there exists a unique global-in-time strong solution $(\rho,{\bf u},\Phi)$ to the initial and boundary problem (1.1)-(1.5)-(1.6):
$\rho,~{}{\bf u}\in L^{2m}\left(\Omega;C\left([0,T];H^{3}\left(U\right)\right)\right),\quad\Phi\in L^{2m}\left(\Omega;C\left([0,T];H^{5}\left(U\right)\right)\right),\forall~{}T>0,$ (1.19)
up to a modification, for any fixed integer $m\geqslant 2$.
Moreover, for the small perturbation problem, there hold the the existence of invariant measure and decay rate.
Theorem 1.2 (Convergence to steady state).
Assume that the stochastic forces satisfies
$\sum a_{k}^{2}=1,\quad\left|Y\left(\rho,{\bf u}\right)\right|\leqslant C\left|\rho{\bf u}\right|,\quad\left\|\nabla_{\rho,{\bf u}}Y\left(\rho,{\bf u}\right)\right\|_{L^{\infty}}\leqslant C,\quad\left\|\nabla_{\rho,{\bf u}}^{2}Y\left(\rho,{\bf u}\right)\right\|_{L^{\infty}}\leqslant C.$ (1.20)
Here, $\nabla_{\rho,{\bf u}}$ denotes the differential operator with respect to $\rho$ and ${\bf u}$. If there exists a constant $\varepsilon>0$ such that the initial condition $\left(\rho_{0},{\bf u}_{0},\Phi_{0}\right)$ satisfies (1.18) and
$\mathbb{E}\left[\left(\left\|\rho_{0}-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}_{0}\right\|_{3}^{2}+\left\|\nabla\Phi_{0}-\nabla\bar{\Phi}\right\|^{2}\right)^{m}\right]\leqslant\varepsilon^{2m},\quad\forall~{}m\geqslant 2,$ (1.21)
then there hold the decay rate and the existence of invariant measure:
  1. (1)
    there are positive constants $C$ and $\alpha$ such that the expectation
    $\mathbb{E}\left[\left(\sup\limits_{s\in[0,t]}\left(\left\|\rho-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}\right\|_{3}^{2}+\left\|\nabla\Phi-\nabla\bar{\Phi}\right\|^{2}\right)\right)^{m}\right]$
    $\leqslant$ $Ce^{-\alpha mt}\mathbb{E}\left[\left(\left\|\rho_{0}-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}_{0}\right\|_{3}^{2}+\left\|\nabla\Phi_{0}-\nabla\bar{\Phi}\right\|^{2}\right)^{m}\right],$ (1.22)
    holds, where $C$ is independent on $t$ and $C$ is the $m$-th power of some constant;
  2. (2)
    the invariant measure generated by $\frac{1}{T}\int_{0}^{T}\mathcal{L}\left(\rho\right)\times\mathcal{L}\left({\bf u}\right)\times\mathcal{L}\left(\Phi\right)\operatorname{d}t$ is exactly the Dirac measure of steady state $\left(\bar{\rho},0,\bar{\Phi}\right)$.
Remark 1.1.
After passing to the limit $t\rightarrow\infty$ in (1), the stationary solution coincides with the steady state $\mathbb{P}$ a.s., since the $m$-th moment of their difference tends to zero.
Remark 1.2.
If for any $\omega\in\Omega$,
$\left(\left\|\rho_{0}-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}_{0}\right\|_{3}^{2}+\left\|\nabla\Phi_{0}-\nabla\bar{\Phi}\right\|^{2}\right)\leqslant\varepsilon^{2},$ (1.23)
then there exists some constant $\tilde{C}$, such that the asymptotic stability holds $\mathbb{P}$ a.s.:
$\sup\limits_{s\in[0,t]}\left(\left\|\rho-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}\right\|_{3}^{2}+\left\|\nabla\Phi-\nabla\bar{\Phi}\right\|^{2}\right)\leqslant 2\tilde{C}e^{-\alpha t}\varepsilon^{2}.$ (1.24)
Actually, by Chebyshev’s inequality (see Appendix 5), it holds
[$\mathbb{P}\left[\left\{\omega\in\Omega|\left\|\rho-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}\right\|_{3}^{2}+\left\|\nabla\Phi-\nabla\bar{\Phi}\right\|^{2}>2\tilde{C}e^{-\alpha t}\varepsilon^{2}\right\}\right]$
$\leqslant$ $\frac{\mathbb{E}\left[\left|\left\|\rho-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}\right\|_{3}^{2}+\left\|\nabla\Phi-\nabla\bar{\Phi}\right\|^{2}\right|^{m}\right]}{\left(2\tilde{C}e^{-\alpha t}\varepsilon^{2}\right)^{m}}$ (1.25)
$\leqslant$ $\frac{\mathbb{E}\left[\tilde{C}e^{-\alpha t}\left(\left\|\rho_{0}-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}_{0}\right\|_{3}^{2}+\left\|\nabla\Phi_{0}-\nabla\bar{\Phi}\right\|^{2}\right)^{m}\right]}{\left(2\tilde{C}e^{-\alpha t}\varepsilon^{2}\right)^{m}}=\frac{1}{2^{m}}.$
Let $m\rightarrow\infty$, then it holds
[$\mathbb{P}\left[\left\{\omega\in\Omega|\left\|\rho-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}\right\|_{3}^{2}+\left\|\nabla\Phi-\nabla\bar{\Phi}\right\|^{2}>2\tilde{C}e^{-\alpha t}\varepsilon^{2}\right\}\right]\rightarrow 0,$
i.e.,
$\left\|\rho-\bar{\rho}\right\|_{3}^{2}+\left\|{\bf u}\right\|_{3}^{2}+\left\|\nabla\Phi-\nabla\bar{\Phi}\right\|^{2}\leqslant 2\tilde{C}e^{-\alpha t}\varepsilon^{2}$ holds $\mathbb{P}$
a.s. for every $s\in[0,t]$.
Remark 1.3.
The argument in this paper implies the same existence and asymptotic stability of solutions around the steady state for the 2-D system with insulating boundary conditions. Repeating the argument, by Sobolev’s embedding, the existence of perturbed solutions and asymptotic stability of steady state for 1-D system with insulating boundary conditions holds: $\rho$ and ${\bf u}$ are in $L^{2m}\left(\Omega;C\left([0,T];H^{2}\left(U\right)\right)\right)$, $\Phi\in L^{2m}\left(\Omega;C\left([0,T];H^{4}\left(U\right)\right)\right)$ in $\left(\Omega,\mathcal{F},\mathbb{P}\right)$.
As mentioned before, the study of stochastic Euler-Poisson equations, totally from the existing studies for the deterministic case, is new and challenging. The idea of the proof is as follows. We first prove the local existence by Banach’s fixed point theorem, then we establish the uniform energy estimates in time $t$ to show the global existence of $\left(\sigma,{\bf u},\phi\right)$. Furthermore, we prove the weighted energy estimates so that we can obtain the asymptotic stability for steady states with the insulating boundary conditions. The a priori estimates imply the tightness of approximates measures, which will converge to an invariant measure by Krylov-Bogoliubov’s theorem in a complete probability space. The global existence does not require the small perturbation condition (1.20) and (1.21). However, the existence of invariant measure in Theorem 4.2, requires (1.20) and (1.21). From the weighted energy estimates, we then prove that the invariant measure for (1.1) is exactly the law of steady state, c.f. Section 4. This intricate relationship has not been uncovered in the asymptotic behavior analysis of stochastic Navier-Stokes equations yet [41, 20].
Here we explain in detail the main difficulties we face to and the strategies we are going to propose.
  1. (1)
    No temporal solutions due to the stochastic term. Since Brownian motion is at most Höder-$\frac{1}{2}-$ continuous with respect to $t$ and it is nowhere differentiable, we do not have $\frac{\operatorname{d}W}{\operatorname{d}t}$ or $\frac{\operatorname{d}\left(\rho{\bf u}\right)}{\operatorname{d}t}$ either. No temporal derivative is involved in the norm of solutions. Thus, in deterministic cases [21, 42], the spatial estimates bounded by the temporal derivatives estimates like
    $\qquad\|\left(\rho-\bar{\rho}\right)\|^{2}+\|\nabla\left(\rho-\bar{\rho}\right)\|^{2}+\|\nabla\cdot{\bf u}\|^{2}\leqslant C\left(\left\|u_{t}\right\|^{2}+\left\|\left(\rho-\bar{\rho}\right)_{t}\right\|^{2}+\|u\|^{2}+\interleave w\interleave^{3}\right),$ (1.26)
[…]
    It is interesting that the noise in form of (1.4) is in the higher order of ${\bf u}$ than the Lipschitz continuous on ${\bf u}$. This is reasonable when we consider the small perturbation around the steady state, which is different with most cases in which Lipshcitz continuous coefficients give birth to wellposedness. In this case the influence of stochastic force does not been exaggerated so much.
  2. (2)
    Weighted energy estimates on account of the estimates of the stochastic integral. Recalling the 3-D deterministic case [21], for instance, based on the energy estimates, one can obtain ordinary differential inequality (ODI). Then they multiply the ODI with the exponential function of $t$ directly to facilitate the stability analysis. However, in this paper, in order to estimates the stochastic term, we apply the Burkholder-Davis-Gundy’s inequality to the stochastic integral of the Wiener process. Then the a priori estimates (2.221) is already in the form of time integrals rather than an ODI. Integration with respect to time twice could not imply the asymptotic stability. Consequently, direct acquisition of asymptotic stability becomes challenging. To overcome this obstacle, we employ the weighted energy estimates. Moreover, the weight is determined by the a priori estimates which should be obtained first, cf. Section 3.1.
This paper is organized as follows. Section 2 is dedicated to establishing the global existence of solutions around the steady state. In Section 3, we investigate the asymptotic stability of semiconductor equations. Finally, in Section 4, we demonstrate the existence and property of invariant measures. Section 5 is the Appendix, in which we provide an overview of stochastic analysis theories that are employed in this study.
2. Global existence of solutions
In this section, we first establish the local existence of strong solutions by Banach’s fixed point theorem. Specifically, we derive the system of perturbed solutions in matrix notation by (1.1). In Step 1, We symmetrize it to simplify the energy estimates and proceed to linearize the system. In Step 2, following a standard procedure in view of Picard interation, we establish the uniform estimates onto mapping. We utilize Itô’s formula and the Burkholder-Davis-Gundy inequality to estimate the stochastic force. In Step 3, we demonstrate contraction. In Step 4, we get the a priori estimates in §2.2 so as to obtain the global existence of $(\rho-\bar{\rho},{\bf u})$, or equivalently, $\rho,{\bf u}\in L^{2m}\left(\Omega;C\left([0,T];H^{3}\left(U\right)\right)\right)$ in §2.3. Step 5 is about the proof of global existence.
In form of $\left(\sigma,{\bf u},\phi\right)=\left(\rho-\bar{\rho},{\bf u},\Phi-\bar{\Phi}\right)$, the hydrodynamic system deforms into
$\left\{\begin{array}[]{l}\sigma_{t}+\nabla\cdot\left(\left(\bar{\rho}+\sigma\right){\bf u}\right)=0,\\ \operatorname{d}\left({\bf u}\right)+\left(\left({\bf u}\cdot\nabla\right){\bf u}+{\bf u}+\nabla Q\left(\bar{\rho}+\sigma\right)-\nabla Q\left(\bar{\rho}\right)\right)\operatorname{d}t=\nabla\phi\operatorname{d}t+\frac{\mathbb{F}}{\bar{\rho}+\sigma}\operatorname{d}W,\\ \triangle\phi=\sigma.\end{array}\right.$ (2.1)
Here we view $\tau$ as a constant $1$ without loss of generality for the stability analysis. In terms of component, by Taylor’s expansion, it holds
$\nabla Q\left(\bar{\rho}+\sigma\right)-\nabla Q\left(\bar{\rho}\right)=\left(Q\left(\bar{\rho}+\sigma\right)-Q\left(\bar{\rho}\right)\right)_{,i}$
$=$ $Q^{\prime}\left(\bar{\rho}+\sigma\right)\left(\bar{\rho}+\sigma\right)_{,i}-Q^{\prime}\left(\bar{\rho}\right)\bar{\rho}_{,i}$
$=$ $Q^{\prime}\left(\bar{\rho}+\sigma\right)\sigma_{,i}+\left(Q^{\prime}\left(\bar{\rho}+\sigma\right)-Q^{\prime}\left(\bar{\rho}\right)\right)\bar{\rho}_{,i}$ (2.2)
$=$ $Q^{\prime}\left(\bar{\rho}+\sigma\right)\sigma_{,i}+Q^{\prime\prime}\left(\bar{\rho}\right)\sigma\bar{\rho}_{i}+\left(Q^{\prime}\left(\bar{\rho}+\sigma\right)-Q^{\prime}\left(\bar{\rho}\right)-Q^{\prime\prime}\left(\bar{\rho}\right)\sigma\right)\bar{\rho}_{,i}$
::=:= $Q^{\prime}\left(\bar{\rho}+\sigma\right)\sigma_{,i}+Q^{\prime\prime}\left(\bar{\rho}\right)\sigma\bar{\rho}_{,i}+h_{i},$
where $\left(\cdot\right)_{,i}$ means the derivative with respect to $x_{i}$, and
$h_{i}=O\left(\sigma^{2}\right).$ (2.3)
In term of component, there holds
$\phi_{,i}=\triangle^{-1}\sigma_{,i},$ (2.4)
where $\triangle^{-1}$ is well-defined under the condition (1.5). In matrix notation, denoting $\mathbf{w}=\left[\begin{array}[]{l}\sigma\\ {\bf u}\end{array}\right]$, we write the system as
$\operatorname{d}\mathbf{w}+\left(\mathcal{A}^{1}\mathbf{w}_{,1}+\mathcal{A}^{2}\mathbf{w}_{,2}+\mathcal{A}^{3}\mathbf{w}_{,3}+\mathcal{B}\mathbf{w}+\mathcal{L}_{{\bf u}}\right)\operatorname{d}t=\mathcal{L}_{\mathcal{\phi}}\operatorname{d}t+f,$ (2.5)
where
$\mathcal{A}^{1}=\left[\begin{array}[]{cccc}u^{1}&\bar{\rho}+\sigma&0&0\\ Q^{\prime}\left(\bar{\rho}+\sigma\right)&u^{1}&0&0\\ 0&0&u^{1}&0\\ 0&0&0&u^{1}\end{array}\right],$ (2.10)
$\mathcal{A}^{2}=\left[\begin{array}[]{cccc}u^{2}&0&\bar{\rho}+\sigma&0\\ 0&u^{2}&0&0\\ Q^{\prime}\left(\bar{\rho}+\sigma\right)&0&u^{2}&0\\ 0&0&0&u^{2}\end{array}\right],$ (2.15)
$\mathcal{A}^{3}=\left[\begin{array}[]{cccc}u^{3}&0&0&\bar{\rho}+\sigma\\ 0&u^{3}&0&0\\ 0&0&u^{3}&0\\ Q^{\prime}\left(\bar{\rho}+\sigma\right)&0&0&u^{3}\end{array}\right],$ (2.20)
$\mathcal{B}=\left[\begin{array}[]{cccc}0&\bar{\rho}_{,1}&\bar{\rho}_{,2}&\bar{\rho}_{,3}\\ Q^{\prime\prime}\left(\bar{\rho}\right)\bar{\rho}_{,1}&0&0&0\\ Q^{\prime\prime}\left(\bar{\rho}\right)\bar{\rho}_{,2}&0&0&0\\ Q^{\prime\prime}\left(\bar{\rho}\right)\bar{\rho}_{,3}&0&0&0\end{array}\right],$ (2.25)
$\mathcal{L}_{{\bf u}}=\left[\begin{array}[]{c}0\\ u^{1}\\ u^{2}\\ u^{3}\end{array}\right],\quad\mathcal{L}_{\mathcal{\phi}}=\left[\begin{array}[]{c}0\\ \triangle^{-1}\sigma_{,1}\\ \triangle^{-1}\sigma_{,2}\\ \triangle^{-1}\sigma_{,3}\end{array}\right],\quad f=-\left[\begin{array}[]{c}0\\ h\left(\sigma\right)_{,1}-\mathbb{F}^{1}\operatorname{d}W\\ h\left(\sigma\right)_{,2}-\mathbb{F}^{2}\operatorname{d}W\\ h\left(\sigma\right)_{,3}-\mathbb{F}^{3}\operatorname{d}W\end{array}\right].$ (2.38)
Step 1: Symmetrizing and Linearizing.
We define the symmetrizer $\mathcal{D}=\operatorname{diag}\left[Q^{\prime}\left(\bar{\rho}+\sigma\right),\bar{\rho}+\sigma,\bar{\rho}+\sigma,\bar{\rho}+\sigma\right]:=\operatorname{diag}\left[d_{1},d_{2},d_{3},d_{4}\right]$. Then the system deforms into
$\mathcal{D}\operatorname{d}\mathbf{w}+\left(\tilde{\mathcal{A}}^{1}\mathbf{w}_{,1}+\tilde{\mathcal{A}}^{2}\mathbf{w}_{,2}+\tilde{\mathcal{A}}^{3}\mathbf{w}_{,3}+\tilde{\mathcal{B}}\mathbf{w}+\tilde{\mathcal{L}}_{{\bf u}}\right)\operatorname{d}t=\tilde{\mathcal{L}}_{\phi}\operatorname{d}t+\tilde{f},$ (2.39)
where
$\tilde{\mathcal{A}}^{1}=\left[\begin{array}[]{cccc}u^{1}Q^{\prime}\left(\bar{\rho}+\sigma\right)&\left(\bar{\rho}+\sigma\right)Q^{\prime}\left(\bar{\rho}+\sigma\right)&0&0\\ \left(\bar{\rho}+\sigma\right)Q^{\prime}\left(\bar{\rho}+\sigma\right)&\left(\bar{\rho}+\sigma\right)u^{1}&0&0\\ 0&0&\left(\bar{\rho}+\sigma\right)u^{1}&0\\ 0&0&0&\left(\bar{\rho}+\sigma\right)u^{1}\end{array}\right],$ (2.44)
$\tilde{\mathcal{A}}^{2}=\left[\begin{array}[]{cccc}u^{2}Q^{\prime}\left(\bar{\rho}+\sigma\right)&0&\left(\bar{\rho}+\sigma\right)Q^{\prime}\left(\bar{\rho}+\sigma\right)&0\\ 0&\left(\bar{\rho}+\sigma\right)u^{2}&0&0\\ \left(\bar{\rho}+\sigma\right)Q^{\prime}\left(\bar{\rho}+\sigma\right)&0&\left(\bar{\rho}+\sigma\right)u^{2}&0\\ 0&0&0&\left(\bar{\rho}+\sigma\right)u^{2}\end{array}\right],$ (2.49)
$\tilde{\mathcal{A}}^{3}=\left[\begin{array}[]{cccc}u^{3}Q^{\prime}\left(\bar{\rho}+\sigma\right)&0&0&\left(\bar{\rho}+\sigma\right)Q^{\prime}\left(\bar{\rho}+\sigma\right)\\ 0&\left(\bar{\rho}+\sigma\right)u^{3}&0&0\\ 0&0&\left(\bar{\rho}+\sigma\right)u^{3}&0\\ \left(\bar{\rho}+\sigma\right)Q^{\prime}\left(\bar{\rho}+\sigma\right)&0&0&\left(\bar{\rho}+\sigma\right)u^{3}\end{array}\right],$ (2.54)
~$\tilde{\mathcal{B}}=\left[\begin{array}[]{cccc}0&\bar{\rho}_{,1}&\bar{\rho}_{,2}&\bar{\rho}_{,3}\\ Q^{\prime\prime}\left(\bar{\rho}\right)\bar{\rho}_{,1}&0&0&0\\ Q^{\prime\prime}\left(\bar{\rho}\right)\bar{\rho}_{,2}&0&0&0\\ Q^{\prime\prime}\left(\bar{\rho}\right)\bar{\rho}_{,3}&0&0&0\end{array}\right],$ (2.59)
~$\tilde{\mathcal{L}}_{{\bf u}}=\left(\bar{\rho}+\sigma\right)\left[\begin{array}[]{c}0\\ u^{1}\\ u^{2}\\ u^{3}\end{array}\right],\quad\tilde{\mathcal{L}}_{\phi}=\left(\bar{\rho}+\sigma\right)\left[\begin{array}[]{c}0\\ \triangle^{-1}\sigma_{,1}\\ \triangle^{-1}\sigma_{,2}\\ \triangle^{-1}\sigma_{,3}\end{array}\right],$ (2.68)
$\tilde{f}=-\left(\bar{\rho}+\sigma\right)\left[\begin{array}[]{c}0\\ h\left(\sigma\right)_{,1}\operatorname{d}t-\mathbb{F}^{1}\operatorname{d}W\\ h\left(\sigma\right)_{,2}\operatorname{d}t-\mathbb{F}^{2}\operatorname{d}W\\ h\left(\sigma\right)_{,3}\operatorname{d}t-\mathbb{F}^{3}\operatorname{d}W\end{array}\right].$ (2.73)
2.1. Local existence
In this subsection, the main estimates for the stochastic forces are taking the assumption of (1.20) for instance. Similar to the approach in [37, 45], we first linearize the system and then we use Banach’s fixed point theorem to get the local existence by the a priori energy estimates.
The linearized system is
$\mathcal{D}(\hat{\sigma})\operatorname{d}\mathbf{w}+\left(\tilde{\mathcal{A}}^{1}\left(\hat{\mathbf{w}}\right)\mathbf{w}_{,1}+\tilde{\mathcal{A}}^{2}(\hat{\mathbf{w}})\mathbf{w}_{,2}+\tilde{\mathcal{A}}^{3}(\hat{\mathbf{w}})\mathbf{w}_{,3}+\tilde{\mathcal{B}}\mathbf{w}\right)\operatorname{d}t$ (2.74)
$=$ ~$-\tilde{\mathcal{L}}_{\hat{{\bf u}}}\left(\hat{\sigma},{\bf u}\right)\operatorname{d}t+\tilde{\mathcal{L}}_{\hat{\phi}}\left(\hat{\sigma},\hat{\phi}\right)\operatorname{d}t+\tilde{f}(\hat{\mathbf{w}}),$
where $\hat{\mathbf{w}}=\left[\begin{array}[]{l}\hat{\sigma}\\ \hat{{\bf u}}\end{array}\right]$ is given, $\hat{\sigma}\in C\left([0,T];H^{3}\left(U\right)\right)$, $\hat{{\bf u}}\in C\left([0,T];H^{3}\left(U\right)\right)$. We denote $M=\sup\limits_{t\in[0,T]}\left\|\hat{\sigma},\hat{{\bf u}}\right\|_{3}$.
Step 2: Estimates for the uniform upper bound.
By Itô’s formula (see Appendix 5), it holds
$\int_{U}\operatorname{d}\left(\frac{1}{2}\mathcal{D}\mathbf{w}\cdot\mathbf{w}\right)\operatorname{d}x=\int_{U}\frac{1}{2}\operatorname{d}\mathcal{D}\mathbf{w}\cdot\mathbf{w}\operatorname{d}x+\int_{U}\mathcal{D}\mathbf{w}\cdot\operatorname{d}\mathbf{w}\operatorname{d}x+\int_{U}\mathcal{D}\mathbb{F}\cdot\mathbb{F}\operatorname{d}x\operatorname{d}t.$ (2.75)
We integrate
$\mathcal{D}\operatorname{d}\mathbf{w}\cdot\mathbf{w}+\left(\tilde{\mathcal{A}}^{1}\left(\hat{\mathbf{w}}\right)\mathbf{w}_{,1}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{2}\left(\hat{\mathbf{w}}\right)\mathbf{w}_{,2}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{3}\left(\hat{\mathbf{w}}\right)\mathbf{w}_{,3}\cdot\mathbf{w}+\tilde{\mathcal{B}}\mathbf{w}\cdot\mathbf{w}\right)\operatorname{d}t$ (2.76)
$=$ ~$-\tilde{\mathcal{L}}_{\hat{{\bf u}}}\cdot\mathbf{w}\operatorname{d}t+\tilde{\mathcal{L}}_{\hat{\phi}}\cdot\mathbf{w}\operatorname{d}t+\tilde{f}\cdot\mathbf{w}$
over the domain $U$, we gain
$\int_{U}\mathcal{D}\mathbf{w}\cdot\operatorname{d}\mathbf{w}\operatorname{d}x$
$=$ $\int_{U}\left(-\left(\tilde{\mathcal{A}}^{1}\mathbf{w}_{,1}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{2}\mathbf{w}_{,2}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{3}\mathbf{w}_{,3}\cdot\mathbf{w}+\tilde{\mathcal{B}}\mathbf{w}\cdot\mathbf{w}\right)-\tilde{\mathcal{L}}_{\hat{{\bf u}}}\cdot\mathbf{w}+\tilde{\mathcal{L}}_{\hat{\phi}}\cdot\mathbf{w}\right)\operatorname{d}x\operatorname{d}t$ (2.77)
$+\int_{U}\nabla h\left(\hat{\sigma}\right)\cdot\mathbf{w}\operatorname{d}x\operatorname{d}t+\int_{U}\mathcal{D}\mathbb{F}\operatorname{d}W\cdot\mathbf{w}\operatorname{d}x.$
By the integration by parts, we have
$\int_{U}\left(\tilde{\mathcal{A}}^{1}\mathbf{w}_{,1}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{2}\mathbf{w}_{,2}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{3}\mathbf{w}_{,3}\cdot\mathbf{w}+\tilde{\mathcal{B}}\mathbf{w}\cdot\mathbf{w}\right)\operatorname{d}x\operatorname{d}t$
$=$ $\int_{U}\left(-\frac{1}{2}\left(\mathbf{w}\tilde{\mathcal{A}}^{1}_{1}\mathbf{w}+\mathbf{w}\tilde{\mathcal{A}}^{2}_{2}\mathbf{w}+\mathbf{w}\tilde{\mathcal{A}}^{3}_{3}\mathbf{w}\right)+\tilde{\mathcal{B}}\left|\mathbf{w}\right|^{2}\right)\operatorname{d}x\operatorname{d}t+\int_{U}\left(\mathbf{w}\tilde{\mathcal{A}}^{j}\mathbf{w}\right)_{,j}\operatorname{d}x\operatorname{d}t.$ (2.78)
On account of the insulated boundary condition $\hat{{\bf u}}\cdot\nu|_{\partial U}=0$, it holds
$\int_{U}\left(\mathbf{w}\tilde{\mathcal{A}}^{j}\mathbf{w}\right)_{,j}\operatorname{d}x$ (2.79)
$=$ $\int_{\partial U}\left(\left(\hat{{\bf u}}\cdot\nu\right)\left(Q^{\prime}\left(\bar{\rho}+\hat{\sigma}\right)\sigma^{2}+\left(\bar{\rho}+\hat{\sigma}\right)\left|{\bf u}\right|^{2}+2\hat{\rho}{\bf u}Q^{\prime}\left(\bar{\rho}+\hat{\sigma}\right)\left(\bar{\rho}+\hat{\sigma}\right)\right)\right)\operatorname{d}S\equiv 0.$
In summary, there holds
$\int_{U}\operatorname{d}\left(\frac{1}{2}\mathcal{D}\mathbf{w}\cdot\mathbf{w}\right)\operatorname{d}x$
$=$ $\int_{U}\frac{1}{2}\operatorname{d}\mathcal{D}\mathbf{w}\cdot\mathbf{w}\operatorname{d}x-\int_{U}\left(-\frac{1}{2}\left(\mathbf{w}\tilde{\mathcal{A}}^{1}_{1}\mathbf{w}+\mathbf{w}\tilde{\mathcal{A}}^{2}_{2}\mathbf{w}+\mathbf{w}\tilde{\mathcal{A}}^{3}_{3}\mathbf{w}\right)+\tilde{\mathcal{B}}\left|\mathbf{w}\right|^{2}\right)\operatorname{d}x\operatorname{d}t$ (2.80)
$+\int_{U}\left(-\tilde{\mathcal{L}}_{\hat{{\bf u}}}\cdot\mathbf{w}+\tilde{\mathcal{L}}_{\hat{\phi}}\cdot\mathbf{w}\right)\operatorname{d}x\operatorname{d}t+\int_{U}\nabla h\left(\hat{\sigma}\right)\cdot\mathbf{w}\operatorname{d}x\operatorname{d}t+\int_{U}\mathcal{D}\mathbb{F}\operatorname{d}W\cdot\mathbf{w}\operatorname{d}x$
$+\int_{U}\mathcal{D}\mathbb{F}\cdot\mathbb{F}\operatorname{d}x\operatorname{d}t.$
Direct calculation shows that
$-\frac{1}{2}\tilde{\mathcal{A}}^{i}_{,i}+\tilde{\mathcal{B}}-\operatorname{diag}\left[-\frac{1}{2}\left(u^{i}Q^{\prime}\left(\rho\right)\right)_{,i},-\frac{1}{2}\left(u^{i}\rho\right)_{,i},-\frac{1}{2}\left(u^{i}\rho\right)_{,i},-\frac{1}{2}\left(u^{i}\rho\right)_{,i}\right]$
$=$ $\left[\begin{array}[]{cccc}0&\footnotesize{-\frac{1}{2}\left\{\rho q\right\}_{,1}+\bar{\rho}_{,1}q}&-\frac{1}{2}\left\{\rho q\right\}_{,2}+\bar{\rho}_{,2}q&-\frac{1}{2}\left\{\rho q\right\}_{,3}+\bar{\rho}_{,3}q\\ -\frac{1}{2}\left\{\rho q\right\}_{,1}+\rho Q^{\prime\prime}\left(\bar{\rho}\right)\bar{\rho}&0&0&0\\ -\frac{1}{2}\left\{\rho q\right\}_{,2}+\rho Q^{\prime\prime}\left(\bar{\rho}\right)\bar{\rho}&0&0&0\\ -\frac{1}{2}\left\{\rho q\right\}_{,3}+\rho Q^{\prime\prime}\left(\bar{\rho}\right)\bar{\rho}&0&0&0\end{array}\right]$ (2.85)
is anti-symmetric [21], where $q=Q^{\prime}\left(\rho\right)$. Then we estimate
$\int_{U}\left(-\frac{1}{2}\left(\mathbf{w}\tilde{\mathcal{A}}^{1}_{1}\mathbf{w}+\mathbf{w}\tilde{\mathcal{A}}^{2}_{2}\mathbf{w}+\mathbf{w}\tilde{\mathcal{A}}^{3}_{3}\mathbf{w}\right)+\tilde{\mathcal{B}}\left|\mathbf{w}\right|^{2}\right)\operatorname{d}x\operatorname{d}t\leqslant C\left\|\mathbf{w}\right\|^{2}\left(\left\|\hat{\sigma}\right\|_{3}+\left\|\hat{{\bf u}}\right\|_{3}\right)\operatorname{d}t.$ (2.86)
Recalling (2.68), we have
$\int_{U}\tilde{\mathcal{L}}_{\hat{{\bf u}}}\left(\hat{\sigma},{\bf u}\right)\cdot\mathbf{w}\operatorname{d}x\operatorname{d}t=\int_{U}\left(\bar{\rho}+\hat{\sigma}\right)\left|{\bf u}\right|^{2}\operatorname{d}x\operatorname{d}t\geqslant C\int_{U}\bar{\rho}\left|{\bf u}\right|^{2}\operatorname{d}x\operatorname{d}t,$ (2.87)
and
$\int_{U}\tilde{\mathcal{L}}_{\phi}\left(\hat{\sigma},\phi\right)\cdot\mathbf{w}\operatorname{d}x\operatorname{d}t=$ $\int_{U}\left(\bar{\rho}+\hat{\sigma}\right)\nabla\phi\cdot{\bf u}\operatorname{d}x\operatorname{d}t=-\int_{U}\nabla\cdot\left(\left(\bar{\rho}+\hat{\sigma}\right){\bf u}\right)\phi\operatorname{d}x\operatorname{d}t$ (2.88)
$=$ $\int_{U}\sigma_{t}\phi\operatorname{d}x\operatorname{d}t=\int_{U}\left(\triangle\phi\right)_{t}\phi\operatorname{d}x\operatorname{d}t=-\operatorname{d}\int_{U}\left|\nabla\phi\right|^{2}\operatorname{d}x.$
For $\tilde{f}$, there holds
$\int_{0}^{t}\int_{U}\tilde{f}\cdot\mathbf{w}\operatorname{d}x=$ $C\int_{0}^{t}\int_{U}\hat{\sigma}^{2}\cdot{\bf u}\operatorname{d}x\operatorname{d}t+\left|\int_{0}^{t}\int_{U}\left(\bar{\rho}+\hat{\sigma}\right)\mathbb{F}\cdot{\bf u}\operatorname{d}x\operatorname{d}W\right|,$ (2.89)
where
$\int_{0}^{t}\int_{U}\hat{\sigma}^{2}\cdot{\bf u}\operatorname{d}x\operatorname{d}s\leqslant\int_{0}^{t}\left\|\hat{\sigma}\right\|_{2}\left\|\hat{\sigma}\right\|\left\|\mathbf{w}\right\|\operatorname{d}s.$ (2.90)
One can see the definition of stochastic integral $\int_{0}^{t}\int_{U}\left(\bar{\rho}+\hat{\sigma}\right)\mathbb{F}\cdot{\bf u}\operatorname{d}x\operatorname{d}W$ in Appendix 5. Since $\left|\mathbb{F}\left(\hat{\sigma},\hat{{\bf u}}\right)\right|^{2}\leqslant C\left|\left(\bar{\rho}+\hat{\sigma}\right)\hat{{\bf u}}\right|^{4}$, there hold
$\int_{U}\mathcal{D}\mathbb{F}\cdot\mathbb{F}\operatorname{d}x\operatorname{d}t\leqslant C\left\|\hat{{\bf u}}\right\|^{2}\left(\left\|\hat{{\bf u}}\right\|_{3}^{2}\left\|\hat{\sigma}\right\|_{3}^{4}\right)\operatorname{d}t\leqslant CM^{8}\operatorname{d}t,$ (2.91)
and
$\mathbb{E}\left[\left|\int_{0}^{t}\int_{U}\mathbb{F}\cdot{\bf u}\operatorname{d}x\operatorname{d}W\right|^{m}\right]\leqslant\mathbb{E}\left[\left(\int_{0}^{t}\left|\int_{U}\mathbb{F}\cdot{\bf u}\operatorname{d}x\right|^{2}\operatorname{d}s\right)^{\frac{m}{2}}\right]$
$\leqslant$ $\mathbb{E}\left[\left(C\int_{0}^{t}\left|\int_{U}\left|\left(\bar{\rho}+\hat{\sigma}\right)\hat{{\bf u}}\right|^{2}{\bf u}\operatorname{d}x\right|^{2}\operatorname{d}s\right)^{\frac{m}{2}}\right]$ (2.92)
$\leqslant$ $\mathbb{E}\left[\left(C\sup\limits_{s\in[0,t]}\left\|{\bf u}\right\|^{2}\int_{0}^{t}\left\|\left(\bar{\rho}+\hat{\sigma}\right)\right\|_{3}^{4}\left\|\hat{{\bf u}}\right\|_{3}^{4}\operatorname{d}s\right)^{\frac{m}{2}}\right]$
$\leqslant$ $\delta_{1}^{m}\mathbb{E}\left[\left(\sup\limits_{s\in[0,t]}\left\|{\bf u}\right\|^{2}\right)^{m}\right]+C_{\delta_{1}}^{m}\mathbb{E}\left[\left(\int_{0}^{t}\left\|\hat{{\bf u}}\right\|_{3}^{4}\left\|\left(\bar{\rho}+\hat{\sigma}\right)\right\|_{3}^{4}\operatorname{d}s\right)^{m}\right],$
by Burkholder-Davis-Gundy’s inequality (see Appendix 5), where $\delta_{1}$ is taken such that $\delta_{1}\sup\limits_{s\in[0,t]}\left\|{\bf u}\right\|^{2}$ can be balanced by the left side. We estimate
$\int_{U}\mathbf{w}\left(\operatorname{d}\mathcal{D}\right)\mathbf{w}\operatorname{d}x$
$=$ $\int_{U}\mathbf{w}\left(\operatorname{diag}\left\{Q^{\prime}\left(\bar{\rho}+\hat{\sigma}\right)_{t},\left(\bar{\rho}+\hat{\sigma}\right)_{t},\left(\bar{\rho}+\hat{\sigma}\right)_{t},\left(\bar{\rho}+\hat{\sigma}\right)_{t}\right\}\right)\mathbf{w}\operatorname{d}x\operatorname{d}t$
$=$ $\int_{U}\left(Q^{\prime\prime}\left(\bar{\rho}+\hat{\sigma}\right)\hat{\sigma}_{t}\sigma^{2}+\hat{\sigma}_{t}\left|{\bf u}\right|^{2}\right)\operatorname{d}x\operatorname{d}t$ (2.93)
$=$ $\int_{U}\left(Q^{\prime\prime}\left(\bar{\rho}\right)+O\left(\hat{\sigma}\right)\right)\left(-\nabla\cdot\left(\left(\bar{\rho}+\hat{\sigma}\right)\hat{{\bf u}}\right)\right)\sigma^{2}\operatorname{d}x+\int_{U}\left(-\nabla\cdot\left(\left(\bar{\rho}+\hat{\sigma}\right)\hat{{\bf u}}\right)\right)\left|{\bf u}\right|^{2}\operatorname{d}x\operatorname{d}t$
$\leqslant$ $C\left\|\mathbf{w}\right\|^{2}\left(\left\|\hat{{\bf u}}\right\|_{2}+\left\|\hat{\sigma}\right\|_{2}\left\|\hat{{\bf u}}\right\|_{2}+\left\|\hat{\sigma}\right\|_{2}\left\|\hat{\sigma}\right\|_{3}\left\|\hat{{\bf u}}\right\|_{2}\right)\operatorname{d}t,$
where $O$ means the same order. In summary, there holds
$\mathbb{E}\left[\left(\sup\limits_{s\in[0,t]}\int_{0}^{s}\operatorname{d}\left(\int_{U}\bar{\rho}\left|\mathbf{w}\right|^{2}\operatorname{d}x+\int_{U}\left|\nabla\phi\right|^{2}\operatorname{d}x\right)+c_{1}\int_{0}^{t}\int_{U}\bar{\rho}\left|{\bf u}\right|^{2}\operatorname{d}x\operatorname{d}s\right)^{m}\right]$
$\leqslant$ $\mathbb{E}\left[\left(C\int_{0}^{t}\left(\left\|\hat{{\bf u}}\right\|^{2}\left(1+\left\|\hat{\sigma}\right\|_{H^{1}}^{2}\right)+\left\|\mathbf{w}\right\|^{2}\left\|\hat{\mathbf{w}}\right\|+\left\|\hat{\mathbf{w}}\right\|^{2}\left\|\mathbf{w}\right\|\right)\operatorname{d}s\right)^{m}\right]$
$+\mathbb{E}\left[\left(C\int_{0}^{t}\interleave\hat{{\bf u}}\interleave^{4}\interleave\left(\bar{\rho}+\hat{\sigma}\right)\interleave^{4}\operatorname{d}s\right)^{m}\right]$ (2.94)
$+\mathbb{E}\left[\left(C\int_{0}^{t}\left\|\mathbf{w}\right\|^{2}\left(\left\|\hat{{\bf u}}\right\|_{2}+\left\|\hat{\sigma}\right\|_{2}\left\|\hat{{\bf u}}\right\|_{2}+\left\|\hat{\sigma}\right\|_{2}\left\|\hat{\sigma}\right\|_{3}\left\|\hat{{\bf u}}\right\|_{2}\right)\operatorname{d}s\right)^{m}\right]$
$\leqslant$ $\mathbb{E}\left[\left(C\int_{0}^{t}\left(M+M^{2}+M^{4}\right)\operatorname{d}s+\int_{0}^{t}M\left\|\mathbf{w}\right\|^{2}\operatorname{d}s+\int_{0}^{t}M^{2}\left\|\mathbf{w}\right\|\operatorname{d}s\right)^{m}\right]$
$+\mathbb{E}\left[\left(C\int_{0}^{t}(M^{4}+M^{8})\operatorname{d}s\right)^{m}\right]+C\mathbb{E}\left[\left(\int_{0}^{t}\left\|\mathbf{w}\right\|^{2}\left(M+M^{2}+M^{3}\right)\operatorname{d}s\right)^{m}\right].$
Furthermore, for $\bar{\rho}$ with a positive lower bound, we have
$\mathbb{E}\left[\left(\sup\limits_{s\in[0,t]}\int_{0}^{s}\operatorname{d}\left(\int_{U}\left|\mathbf{w}\right|^{2}\operatorname{d}x+\int_{U}\left|\nabla\phi\right|^{2}\operatorname{d}x\right)\right)^{m}\right]\leqslant C_{M,m}\left(t^{m}+\mathbb{E}\left[\left(\int_{0}^{t}\left\|\mathbf{w}\right\|^{2}\operatorname{d}s\right)^{m}\right]\right),$ (2.95)
where $C_{M,m}$ is a constant depending on $m,M$. Similarly, we take higher-order derivatives to the system (2.74) up to third order, and we do the a priori estimates. There holds
$\mathbb{E}\left[\left(\sup\limits_{s\in[0,t]}\int_{0}^{s}\operatorname{d}\left(\left\|\mathbf{w}\right\|_{3}^{2}+\left\|\nabla\phi\right\|^{2}\right)\right)^{m}\right]$ (2.96)
$\leqslant$ $C_{M,m}\left(t^{m}+\mathbb{E}\left[\left(\int_{0}^{t}\left(\left\|\mathbf{w}\right\|_{3}^{2}+\left\|\nabla\phi\right\|^{2}\right)\operatorname{d}s\right)^{m}\right]\right).$
By Grönwall’s inequality, we have $\mathbf{w}\in L^{2m}\left(\Omega;C\left([0,T];H^{3}\left(U\right)\right)\right)$. More precisely,
$\mathbb{E}\left[\left(\sup\limits_{s\in[0,t]}\left(\left\|\mathbf{w}\right\|_{3}^{2}+\left\|\nabla\phi\right\|^{2}\right)(s)\right)^{m}\right]$
$\leqslant$ $\mathbb{E}\left[\left(\left(\left\|\mathbf{w}\right\|_{3}^{2}+\left\|\nabla\phi\right\|^{2}\right)(0)\right)^{m}\right]+C_{M,m}t^{m}$ (2.97)
$+\int_{0}^{t}\left(\mathbb{E}\left[\left(\left(\left\|\mathbf{w}\right\|_{3}^{2}+\left\|\nabla\phi\right\|^{2}\right)(0)\right)^{m}\right]+C_{M,m}t^{m}\right)C_{M,m}e^{\int_{0}^{s}C_{M,m}\operatorname{d}\tau}\operatorname{d}s$
$\leqslant$ $\left(\mathbb{E}\left[\left(\left(\left\|\mathbf{w}_{0}\right\|_{3}^{2}+\left\|\nabla\phi_{0}\right\|^{2}\right)\right)^{m}\right]+C_{M,m}t^{m}\right)e^{C_{M,m}t}.$
From the estimates of time shift of solutions similar as (2.1), by applying Kolmogorov-Centov’s theorem (see Appendix 5), following the standard argument in stochastic analysis [5], we deduce the time continuity of $\mathbf{w}$ up to a modification in probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$, and we omit the details.
The iteration scheme is
$\mathcal{D}\left(\sigma_{n-1}\right)\operatorname{d}\mathbf{w}_{n}+\left(\tilde{\mathcal{A}}^{1}\left(\mathbf{w}_{n-1}\right)\mathbf{w}_{n,1}+\tilde{\mathcal{A}}^{2}\left(\mathbf{w}_{n-1}\right)\mathbf{w}_{n,2}+\tilde{\mathcal{A}}^{3}\left(\mathbf{w}_{n-1}\right)\mathbf{w}_{n,3}+\tilde{\mathcal{B}}\mathbf{w}_{n}\right)\operatorname{d}t$ (2.98)
$=$ $-L_{{\bf u}_{n-1}}\left(\sigma_{n-1},{\bf u}_{n}\right)\operatorname{d}t+\tilde{\mathcal{L}}_{\phi}\left(\sigma_{n-1},\phi_{n}\right)\operatorname{d}t+\tilde{f}\left(\mathbf{w}_{n-1}\right).$
By energy estimates (2.1), we take $T_{0}$ such that
$e^{C_{M,m}T_{0}}\leqslant 2,\quad C_{M,m}T_{0}\leqslant\mathbb{E}\left[\left(\left(\left\|\mathbf{w}_{0}\right\|_{3}^{2}+\left\|\nabla\phi_{0}\right\|^{2}\right)\right)^{m}\right],$ (2.99)
if
$\mathbb{E}\left[\left(\sup\limits_{s\in[0,t]}\left\|\mathbf{w}_{n-1}(s)\right\|_{3}^{2}\right)^{m}\right]\leqslant 4\mathbb{E}\left[\left(\left(\left\|\mathbf{w}_{0}\right\|_{3}^{2}+\left\|\nabla\phi_{0}\right\|^{2}\right)\right)^{m}\right],$ (2.100)
then
$\mathbb{E}\left[\left(\sup\limits_{s\in[0,t]}\left\|\mathbf{w}_{n}(s)\right\|_{3}^{2}\right)^{m}\right]\leqslant 4\mathbb{E}\left[\left(\left(\left\|\mathbf{w}_{0}\right\|_{3}^{2}+\left\|\nabla\phi_{0}\right\|^{2}\right)\right)^{m}\right].$ (2.101)
Remark 2.1.
For general stochastic forces without the condition (1.20), there also holds
$\mathbb{E}\left[\left(\sup\limits_{s\in[0,t]}\int_{0}^{s}\operatorname{d}\left(\left\|\mathbf{w}\right\|_{3}^{2}+\left\|\nabla\phi\right\|^{2}\right)\right)^{m}\right]$ (2.102)
$\leqslant$ $C_{M,m}\left(t^{m}+\mathbb{E}\left[\left(\int_{0}^{t}\left(\left\|\mathbf{w}\right\|_{3}^{2}+\left\|\nabla\phi\right\|^{2}\right)\operatorname{d}s\right)^{m}\right]\right).$
with another expression of the constant $C_{M,m}$. Thus, we get the uniform bound by Grönwall’s inequality similarly to the above statement.
Step 3: Contraction.
For $\left\|\mathbf{w}_{n}-\mathbf{w}_{n-1}\right\|_{3}$, we show that it is a Cauchy sequence. $\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)$ satisfies
$\mathcal{D}\left(\sigma_{n-1}\right)\left(\operatorname{d}\mathbf{w}_{n}-\operatorname{d}\mathbf{w}_{n-1}\right)+\left(\mathcal{D}\left(\sigma_{n-1}\right)-\mathcal{D}\left(\sigma_{n-2}\right)\right)\operatorname{d}\mathbf{w}_{n-1}$
$+\tilde{\mathcal{A}}^{1}\left(\mathbf{w}_{n-1}\right)\left(\mathbf{w}_{n,1}-\mathbf{w}_{n-1,1}\right)\operatorname{d}t+\left(\tilde{\mathcal{A}}^{1}\left(\mathbf{w}_{n-1}\right)-\tilde{\mathcal{A}}^{1}\left(\mathbf{w}_{n-2}\right)\right)\mathbf{w}_{n-1,1}\operatorname{d}t$
$+\tilde{\mathcal{A}}^{2}\left(\mathbf{w}_{n-1}\right)\left(\mathbf{w}_{n,2}-\mathbf{w}_{n-1,2}\right)\operatorname{d}t+\left(\tilde{\mathcal{A}}^{2}\left(\mathbf{w}_{n-1}\right)-\tilde{\mathcal{A}}^{2}\left(\mathbf{w}_{n-2}\right)\right)\mathbf{w}_{n-1,2}\operatorname{d}t$
$+\tilde{\mathcal{A}}^{3}\left(\mathbf{w}_{n-1}\right)\left(\mathbf{w}_{n,3}-\mathbf{w}_{n-1,3}\right)\operatorname{d}t+\left(\tilde{\mathcal{A}}^{3}\left(\mathbf{w}_{n-1}\right)-\tilde{\mathcal{A}}^{3}\left(\mathbf{w}_{n-2}\right)\right)\mathbf{w}_{n-1,3}\operatorname{d}t$ (2.103)
+~$+\tilde{\mathcal{B}}\left(\bar{\mathbf{w}}\right)\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\operatorname{d}t$
$=$ ~$-\tilde{\mathcal{L}}_{{\bf u}}\left(\sigma_{n-1},{\bf u}_{n}\right)\operatorname{d}t+\tilde{\mathcal{L}}_{{\bf u}}\left(\sigma_{n-2},{\bf u}_{n-1}\right)\operatorname{d}t+\left(\tilde{\mathcal{L}}_{\phi}\left(\sigma_{n-1},\phi_{n}\right)-\tilde{\mathcal{L}}_{\phi}\left(\sigma_{n-2},\phi_{n-1}\right)\right)\operatorname{d}t$
$+\left(\tilde{f}\left(\mathbf{w}_{n-1}\right)-\tilde{f}\left(\mathbf{w}_{n-2}\right)\right).$
Then we multiply the above formula with $\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)$, the estimates of some terms
$\tilde{\mathcal{A}}^{1}\left(\mathbf{w}_{n-1}\right)\left(\mathbf{w}_{n,1}-\mathbf{w}_{n-1,1}\right)\operatorname{d}t+\tilde{\mathcal{A}}^{2}\left(\mathbf{w}_{n-1}\right)\left(\mathbf{w}_{n,2}-\mathbf{w}_{n-1,2}\right)\operatorname{d}t$ (2.104)
$+\tilde{\mathcal{A}}^{3}\left(\mathbf{w}_{n-1}\right)\left(\mathbf{w}_{n,3}-\mathbf{w}_{n-1,3}\right)\operatorname{d}t+\tilde{\mathcal{B}}\left(\bar{\mathbf{w}}\right)\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\operatorname{d}t$
are similar to (2.1), (2.79) and (2.86), we omit it here. We focus on the estimates of
$\sum\left(\tilde{\mathcal{A}}^{i}\left(\mathbf{w}_{n-1}\right)-\tilde{\mathcal{A}}^{i}\left(\mathbf{w}_{n-2}\right)\right)\mathbf{w}_{n-1,i}\operatorname{d}t,$ (2.105)
and the right-hand side terms in (2.1). By the expression formula of $\tilde{\mathcal{A}}^{i}$, it holds
$\int_{0}^{1}\sum\left(\tilde{\mathcal{A}}^{i}\left(\mathbf{w}_{n-1}\right)-\tilde{\mathcal{A}}^{i}\left(\mathbf{w}_{n-2}\right)\right)\mathbf{w}_{n-1,i}\cdot\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\operatorname{d}x\operatorname{d}t$ (2.106)
$\leqslant$ $C\left\|\mathbf{w}_{n}-\mathbf{w}_{n-1}\right\|\left\|\mathbf{w}_{n-1}-\mathbf{w}_{n-2}\right\|\operatorname{d}t.$
Since
~$-\tilde{\mathcal{L}}_{{\bf u}}\left(\sigma_{n-1},{\bf u}_{n}\right)\operatorname{d}t+\tilde{\mathcal{L}}_{{\bf u}}\left(\sigma_{n-2},{\bf u}_{n-1}\right)\operatorname{d}t$ (2.107)
$=$ (~$\left(\tilde{\mathcal{L}}_{{\bf u}}\left(\sigma_{n-1},{\bf u}_{n-1}\right)-\tilde{\mathcal{L}}_{{\bf u}}\left(\sigma_{n-1},{\bf u}_{n}\right)\right)\operatorname{d}t+\left(\tilde{\mathcal{L}}_{{\bf u}}\left(\sigma_{n-2},{\bf u}_{n-1}\right)-\tilde{\mathcal{L}}_{{\bf u}}\left(\sigma_{n-1},{\bf u}_{n-1}\right)\right)\operatorname{d}t,$
we estimate
$\int_{0}^{1}\left(-\tilde{\mathcal{L}}_{{\bf u}}\left(\sigma_{n-1},{\bf u}_{n}\right)\operatorname{d}t+\tilde{\mathcal{L}}_{{\bf u}}\left(\sigma_{n-2},{\bf u}_{n-1}\right)\right)\cdot\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\operatorname{d}x\operatorname{d}t$
$=$ $-\int_{0}^{1}\left(\bar{\rho}+\sigma_{n-1}\right)\left|{\bf u}_{n}-{\bf u}_{n-1}\right|^{2}\operatorname{d}x\operatorname{d}t-\int_{0}^{1}\left(\sigma_{n-1}-\sigma_{n-2}\right){\bf u}_{n-1}\cdot\left({\bf u}_{n}-{\bf u}_{n-1}\right)\operatorname{d}x\operatorname{d}t$ (2.108)
$\leqslant$ $-\int_{0}^{1}\frac{\bar{\rho}}{2}\left|{\bf u}_{n}-{\bf u}_{n-1}\right|^{2}\operatorname{d}x\operatorname{d}t-\int_{0}^{1}\left(\sigma_{n-1}-\sigma_{n-2}\right){\bf u}_{n-1}\cdot\left({\bf u}_{n}-{\bf u}_{n-1}\right)\operatorname{d}x\operatorname{d}t,$
where
$\int_{0}^{1}\left(\sigma_{n-1}-\sigma_{n-2}\right){\bf u}_{n-1}\left({\bf u}_{n}-{\bf u}_{n-1}\right)\operatorname{d}x\operatorname{d}t\leqslant C\left\|\mathbf{w}_{n}-\mathbf{w}_{n-1}\right\|\left\|\mathbf{w}_{n-1}-\mathbf{w}_{n-2}\right\|\operatorname{d}t.$ (2.109)
Since
(~$\left(\tilde{\mathcal{L}}_{\phi}\left(\sigma_{n-1},\phi_{n}\right)-\tilde{\mathcal{L}}_{\phi}\left(\sigma_{n-2},\phi_{n-1}\right)\right)\operatorname{d}t$
$=$ (~$\left(\tilde{\mathcal{L}}_{\phi}\left(\sigma_{n-1},\phi_{n}\right)-\tilde{\mathcal{L}}_{\phi}\left(\sigma_{n-1},\phi_{n-1}\right)\right)\operatorname{d}t+\left(\tilde{\mathcal{L}}_{\phi}\left(\sigma_{n-1},\phi_{n-1}\right)-\tilde{\mathcal{L}}_{\phi}\left(\sigma_{n-2},\phi_{n-1}\right)\right)\operatorname{d}t,$ (2.110)
then we have
$\int_{0}^{1}\left(\tilde{\mathcal{L}}_{\phi}\left(\sigma_{n-1},\phi_{n}\right)-\tilde{\mathcal{L}}_{\phi}\left(\sigma_{n-2},\phi_{n-1}\right)\right)\cdot\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\operatorname{d}x\operatorname{d}t$
$=$ $\int_{0}^{1}\left(\tilde{\mathcal{L}}_{\phi}\left(\sigma_{n-1},\phi_{n}\right)-\tilde{\mathcal{L}}_{\phi}\left(\sigma_{n-1},\phi_{n-1}\right)\right)\cdot\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\operatorname{d}x\operatorname{d}t$ (2.111)
$+\int_{0}^{1}\left(\tilde{\mathcal{L}}_{\phi}\left(\sigma_{n-1},\phi_{n-1}\right)-\tilde{\mathcal{L}}_{\phi}\left(\sigma_{n-2},\phi_{n-1}\right)\right)\cdot\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\operatorname{d}x\operatorname{d}t.$
By the continuity equation, there holds
$\int_{U}\tilde{\mathcal{L}}_{\phi}\left(\sigma_{n-1},\phi_{n}-\phi_{n-1}\right)\cdot\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\operatorname{d}x\operatorname{d}t$
$=$ $\int_{U}\left(\bar{\rho}+\sigma_{n-1}\right)\nabla\left(\phi_{n}-\phi_{n-1}\right)\cdot\left({\bf u}_{n}-{\bf u}_{n-1}\right)\operatorname{d}x\operatorname{d}t$
$=$ $-\int_{U}\nabla\cdot\left(\left(\bar{\rho}+\sigma_{n-1}\right)\left({\bf u}_{n}-{\bf u}_{n-1}\right)\right)\left(\phi_{n}-\phi_{n-1}\right)\operatorname{d}x\operatorname{d}t$
$=$ $\int_{U}\left(\sigma_{n}-\sigma_{n-1}\right)_{t}\left(\phi_{n}-\phi_{n-1}\right)\operatorname{d}x\operatorname{d}t+\int_{U}\nabla\cdot\left(\left(\sigma_{n-1}-\sigma_{n-2}\right){\bf u}_{n-1}\right)\left(\phi_{n}-\phi_{n-1}\right)\operatorname{d}x\operatorname{d}t$ (2.112)
$=$ $\int_{U}\left(\triangle\left(\phi_{n}-\phi_{n-1}\right)\right)_{t}\left(\phi_{n}-\phi_{n-1}\right)\operatorname{d}x\operatorname{d}t+\int_{U}\nabla\cdot\left(\left(\sigma_{n-1}-\sigma_{n-2}\right){\bf u}_{n-1}\right)\left(\phi_{n}-\phi_{n-1}\right)\operatorname{d}x\operatorname{d}t$
$=$ $-\operatorname{d}\int_{U}\left|\nabla\left(\phi_{n}-\phi_{n-1}\right)\right|^{2}\operatorname{d}x-\int_{U}\left(\sigma_{n-1}-\sigma_{n-2}\right){\bf u}_{n-1}\cdot\nabla\left(\phi_{n}-\phi_{n-1}\right)\operatorname{d}x\operatorname{d}t$
$\leqslant$ $-\operatorname{d}\int_{U}\left|\nabla\left(\phi_{n}-\phi_{n-1}\right)\right|^{2}\operatorname{d}x+C\left\|\mathbf{w}_{n}-\mathbf{w}_{n-1}\right\|\left\|\mathbf{w}_{n-1}-\mathbf{w}_{n-2}\right\|\operatorname{d}t,$
and
$\int_{0}^{1}\left(\tilde{\mathcal{L}}_{\phi}\left(\sigma_{n-1},\phi_{n-1}\right)-\tilde{\mathcal{L}}_{\phi}\left(\sigma_{n-2},\phi_{n-1}\right)\right)\cdot\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\operatorname{d}x\operatorname{d}t$
$=$ $\int_{U}\left(\sigma_{n-1}-\sigma_{n-2}\right)\nabla\phi_{n-1}\cdot\left({\bf u}_{n}-{\bf u}_{n-1}\right)\operatorname{d}x\operatorname{d}t$ (2.113)
$\leqslant$ $C\left\|\mathbf{w}_{n}-\mathbf{w}_{n-1}\right\|\left\|\mathbf{w}_{n-1}-\mathbf{w}_{n-2}\right\|\operatorname{d}t.$
For the terms in $\tilde{f}$, similarly, we have
$\int_{0}^{1}\left(\left(\bar{\rho}+\sigma_{n-1}\right)\nabla h\left(\sigma_{n-1}\right)-\left(\bar{\rho}+\sigma_{n-2}\right)\nabla h\left(\sigma_{n-2}\right)\right)\cdot\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\operatorname{d}x\operatorname{d}t$ (2.114)
$\leqslant$ $C\left\|\mathbf{w}_{n}-\mathbf{w}_{n-1}\right\|\left\|\mathbf{w}_{n-1}-\mathbf{w}_{n-2}\right\|\operatorname{d}t;$
and
$\mathbb{E}\left[\left|\int_{0}^{t}\int_{0}^{1}\left(\mathbb{F}_{n-1}-\mathbb{F}_{n-2}\right)\operatorname{d}W\cdot\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\operatorname{d}x\right|^{m}\right]$
$=$ $\mathbb{E}\left[\left|\int_{0}^{t}\int_{0}^{1}\left(\mathbb{F}_{n-1}-\mathbb{F}_{n-2}\right)\cdot\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\operatorname{d}x\operatorname{d}W\right|^{m}\right]$
$\leqslant$ $\mathbb{E}\left[\left|C\int_{0}^{t}\left|\int_{0}^{1}\left(\mathbb{F}_{n-1}-\mathbb{F}_{n-2}\right)\cdot\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\operatorname{d}x\right|^{2}\operatorname{d}s\right|^{\frac{m}{2}}\right]$ (2.115)
$\leqslant$ $\mathbb{E}\left[\left|C\int_{0}^{t}\left\|\mathbf{w}_{n}-\mathbf{w}_{n-1}\right\|^{2}\left\|\mathbf{w}_{n-1}-\mathbf{w}_{n-2}\right\|^{2}\operatorname{d}s\right|^{\frac{m}{2}}\right]$
$\leqslant$ $\mathbb{E}\left[\left(C\sup_{s\in[0,t]}\left\|\mathbf{w}_{n-1}-\mathbf{w}_{n-2}\right\|^{2}\int_{0}^{t}\left\|\mathbf{w}_{n}-\mathbf{w}_{n-1}\right\|^{2}\operatorname{d}s\right)^{\frac{m}{2}}\right].$
$\leqslant$ $\mathbb{E}\left[\left(\delta_{2}\sup_{s\in[0,t]}\left\|\mathbf{w}_{n-1}-\mathbf{w}_{n-2}\right\|^{2}\right)^{m}\right]+\mathbb{E}\left[\left(\int_{0}^{t}C_{\delta_{2}}\left\|\mathbf{w}_{n}-\mathbf{w}_{n-1}\right\|^{2}\operatorname{d}s\right)^{m}\right].$
By Itô’s formula, we have
$\operatorname{d}\left(\mathcal{D}\left(\sigma_{n-1}\right)\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\cdot\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\right)$
$=$ $\operatorname{d}\mathcal{D}\left(\sigma_{n-1}\right)\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\cdot\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)+2\mathcal{D}\left(\sigma_{n-1}\right)\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\cdot\operatorname{d}\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)$ (2.116)
$+\mathcal{D}\left(\sigma_{n-1}\right)\left\langle\operatorname{d}\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right),\operatorname{d}\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\right\rangle,$
where $\left\langle\operatorname{d}\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right),\operatorname{d}\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\right\rangle$ is a shorthand for the more detailed expression for quadratic variation
$\left\langle~{}\left\langle\operatorname{d}\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right),\operatorname{d}\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\right\rangle,\left\langle\operatorname{d}\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right),\operatorname{d}\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\right\rangle~{}\right\rangle_{\mathcal{H}}^{\frac{1}{2}},$
with a slight abuse of notation. Moreover, we have
$\mathcal{D}\left(\sigma_{n-1}\right)\left\langle\operatorname{d}\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right),\operatorname{d}\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\right\rangle$ (2.117)
$=$ $\left(\bar{\rho}+\sigma_{n-1}\right)\left|\mathbb{F}_{n-1}-\mathbb{F}_{n-2}\right|^{2}\operatorname{d}t,$
and
$\int_{0}^{1}\mathcal{D}\left(\sigma_{n-1}\right)\left\langle\operatorname{d}\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right),\operatorname{d}\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\right\rangle\operatorname{d}x\leqslant C\left\|\mathbf{w}_{n-1}-\mathbf{w}_{n-2}\right\|^{2}\operatorname{d}t.$ (2.118)
Hence it holds
$\int_{0}^{1}\operatorname{d}\left(\mathcal{D}\left(\sigma_{n-1}\right)\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\cdot\left(\mathbf{w}_{n}-\mathbf{w}_{n-1}\right)\right)\operatorname{d}x\leqslant C\left\|\mathbf{w}_{n}-\mathbf{w}_{n-1}\right\|^{2}\operatorname{d}t.$ (2.119)
Combining the above estimates, for some $m\geqslant 2$, we have
$\mathbb{E}\left[\left(\sup\limits_{s\in[0,t]}\int_{0}^{s}\operatorname{d}\left\|\mathbf{w}_{n}-\mathbf{w}_{n-1}\right\|^{2}\right)^{m}\right]$ (2.120)
$\leqslant$ $\mathbb{E}\left[\left|\int_{0}^{t}C\left(\left\|\mathbf{w}_{n}-\mathbf{w}_{n-1}\right\|^{2}+\left\|\mathbf{w}_{n-1}-\mathbf{w}_{n-2}\right\|^{2}+\left\|\mathbf{w}_{n}-\mathbf{w}_{n-1}\right\|\left\|\mathbf{w}_{n-1}-\mathbf{w}_{n-2}\right\|\right)\operatorname{d}s\right|^{m}\right]$
$+\mathbb{E}\left[\left(\delta_{2}\sup_{s\in[0,t]}\left\|\mathbf{w}_{n-1}-\mathbf{w}_{n-2}\right\|^{2}\right)^{m}\right]+\mathbb{E}\left[\left(\int_{0}^{t}C_{\delta_{2}}\left\|\mathbf{w}_{n}-\mathbf{w}_{n-1}\right\|^{2}\operatorname{d}s\right)^{m}\right],$
where $C$ depends on $M$. By Cauchy’s inequality and Jensen’s inequality, we have
$\mathbb{E}\left[\left(\sup_{s\in[0,t]}\left\|\mathbf{w}_{n}-\mathbf{w}_{n-1}\right\|^{2}\right)^{m}\right]$
$\leqslant$ $\mathbb{E}\left[\left(\int_{0}^{t}C\left(\left\|\mathbf{w}_{n}-\mathbf{w}_{n-1}\right\|^{2}+\left\|\mathbf{w}_{n-1}-\mathbf{w}_{n-2}\right\|^{2}+\left\|\mathbf{w}_{n}-\mathbf{w}_{n-1}\right\|\left\|\mathbf{w}_{n-1}-\mathbf{w}_{n-2}\right\|\right)\operatorname{d}s\right)^{m}\right]$
$\leqslant$ $\mathbb{E}\left[\left(\int_{0}^{t}C\left(\left\|\mathbf{w}_{n}-\mathbf{w}_{n-1}\right\|^{2}+\left\|\mathbf{w}_{n-1}-\mathbf{w}_{n-2}\right\|^{2}\right)\operatorname{d}s\right)^{m}\right]$ (2.121)
$\leqslant$ $\int_{0}^{t}\left(\mathbb{E}\left[\left(C_{0}\left\|\mathbf{w}_{n}-\mathbf{w}_{n-1}\right\|^{2}\right)^{m}\right]+\mathbb{E}\left[\left(C_{0}\left\|\mathbf{w}_{n-1}-\mathbf{w}_{n-2}\right\|^{2}\right)^{m}\right]\right)\operatorname{d}s.$
The higher order contraction estimates are proved similarly to zeroth-order, with the same symmetrizing matrix and the important insulating boundary condition, and the detailed proof is omitted here. In summary, we have
$\mathbb{E}\left[\left(\sup_{s\in[0,t]}\left\|\mathbf{w}_{n}-\mathbf{w}_{n-1}\right\|_{3}^{2}\right)^{m}\right]$ (2.122)
$\leqslant$ $\int_{0}^{t}\left(\mathbb{E}\left[\left(C_{0}\left\|\mathbf{w}_{n}-\mathbf{w}_{n-1}\right\|_{3}^{2}\right)^{m}\right]+\mathbb{E}\left[\left(C_{0}\left\|\mathbf{w}_{n-1}-\mathbf{w}_{n-2}\right\|_{3}^{2}\right)^{m}\right]\right)\operatorname{d}s.$
By Grönwall’s inequality, we have
$\mathbb{E}\left[\left(\sup_{s\in[0,t]}\left\|\mathbf{w}_{n}-\mathbf{w}_{n-1}\right\|_{3}^{2}\right)^{m}\right]$ (2.123)
$\leqslant$ $\mathbb{E}\left[\left(\sup_{s\in[0,t]}\left\|\mathbf{w}_{n-1}-\mathbf{w}_{n-2}\right\|_{3}^{2}\right)^{m}\right]C_{0}^{m}t+\int_{0}^{t}\mathbb{E}\left[\left(\sup_{s\in[0,\tau]}\left\|\mathbf{w}_{n-1}-\mathbf{w}_{n-2}\right\|_{3}^{2}\right)^{m}\right]\tau C_{0}^{2r}e^{C_{0}^{m}\tau}\operatorname{d}\tau$
$\leqslant$ $3C_{0}^{m}\mathbb{E}\left[\left(\sup_{s\in[0,\tau]}\left\|\mathbf{w}_{n-1}-\mathbf{w}_{n-2}\right\|_{3}^{2}\right)^{m}\right]t.$
Let $T_{1}\leqslant T_{0}$ and $13C_{0}^{m}T_{1}<$ 1, $e^{C_{0}^{m}T_{1}}\leqslant 2$, then
$\mathbb{E}\left[\left(\sup\limits_{s\in[0,t]}\left\|\mathbf{w}_{n}-\mathbf{w}_{n-1}\right\|_{3}\right)^{m}\right]\leqslant a~{}\mathbb{E}\left[\left(\sup\limits_{s\in[0,t]}\left\|\mathbf{w}_{n-1}-\mathbf{w}_{n-2}\right\|_{3}\right)^{m}\right],\quad a<1,$ (2.124)
where $a=3C_{0}^{m}T_{1}$ with $C_{0}$ depending on the initial data by the onto mapping estimates. Hence, $\mathbf{w}_{n}$ is a Cauchy sequence. By Banach’s fixed point theorem, there exists a unique solution $\mathbf{w}$ in $L^{2m}\left(\Omega;C\left([0,T_{1}];H^{3}\left(U\right)\right)\right)$. Since $\triangle\phi=\sigma$ holds, $\phi$ is also a unique solution in $L^{2m}\left(\Omega;C\left([0,T_{1}];H^{5}\left(U\right)\right)\right)$ up to a constant, with the boundary condition $\nabla\phi\cdot\nu=0$.
By the proof of theorem 5.2.9 in [35], $(\rho,{\bf u},\Phi)$ is the unique strong solutions to SEP, where $\rho,{\bf u}\in C\left([0,T_{1}];H^{3}\left(U\right)\right)$ and $\Phi\in C\left([0,T_{1}];H^{5}\left(U\right)\right)$ hold $\mathbb{P}$ a.s. We give the definition of the local strong solution as follows.
Definition 2.1.
Let $\left(\Omega,\mathcal{F},\mathbb{P}\right)$ be a fixed stochastic basis with a complete right-continuous filtration $\mathcal{F}=\left(\mathcal{F}_{s}\right)_{s\geqslant 0}$ and $W$ be the fixed Wiener process. $\left(\rho,{\bf u},\Phi\right)$ is called a strong solution to initial and boundary problem (1.1)-(1.5)-(1.6)-(1.18)-(1.4), if:
  1. (1)
    $\left(\rho,{\bf u},\Phi\right)$ is adapted to the filtration $\left(\mathcal{F}_{s}\right)_{s\geqslant 0}$;
  2. (2)
    [$\mathbb{P}[\{\left(\rho(0),{\bf u}(0),\Phi(0)\right)=\left(\rho_{0},{\bf u}_{0},\Phi_{0}\right)\}]=1$;
  3. (3)
    the equation of continuity
    $\rho(t)=\rho_{0}-\int_{0}^{t}\nabla\cdot(\rho{\bf u})\operatorname{d}s,$
    holds $\mathbb{P}$ a.s., for any $t\in[0,T_{1}]$;
  4. (4)
    the momentum equation
    ${\bf u}(t)=$ ${\bf u}_{0}-\int_{0}^{t}({\bf u}\cdot\nabla){\bf u}\operatorname{d}s-\int_{0}^{t}\frac{\nabla P(\rho)}{\rho}\operatorname{d}s+\int_{0}^{t}\nabla\Phi\operatorname{d}s-\int_{0}^{t}{\bf u}\operatorname{d}s$ (2.125)
    $+\int_{0}^{t}\frac{\mathbb{F}(\rho,{\bf u})}{\rho}\operatorname{d}W(s),$
    holds $\mathbb{P}$ a.s., for any $t\in[0,T_{1}]$;
  5. (5)
    the electrostatic potential equation
    $\triangle\Phi=\rho-b,$ (2.126)
    holds $\mathbb{P}$ a.s. for any $t\in[0,T_{1}]$.
Remark 2.2.
Reviewing the above proof, (2.124) holds for general stochastic forces without (1.20). Thus, the local existence also holds.
Step 4: Energy estimates.
2.2. Estimates up to third-order
In this subsection, we begin by symmetrizing the system. Then, we proceed with energy estimates up to third order, taking stochastic forces under the condition (1.20) for instance.
2.2.1. Zero-order estimates
For the system (2.5), we define the energy
$\mathcal{E}\left(t\right)=\int_{U}\frac{1}{2}\left(\bar{\rho}\left|{\bf u}\right|^{2}+Q^{\prime}\left(\bar{\rho}\right)\sigma^{2}+\left|\nabla\phi\right|^{2}\right)\operatorname{d}x.$ (2.127)
By Itô’s formula, we have
$\operatorname{d}\int_{U}\frac{1}{2}\left(\bar{\rho}+\sigma\right)\left|{\bf u}\right|^{2}\operatorname{d}x$ (2.128)
$=$ $\int_{U}\frac{1}{2}\operatorname{d}\left(\bar{\rho}+\sigma\right)\left|{\bf u}\right|^{2}\operatorname{d}x+\int_{U}\left(\bar{\rho}+\sigma\right){\bf u}\cdot\operatorname{d}{\bf u}\operatorname{d}x+\int_{U}\frac{1}{2}\left(\bar{\rho}+\sigma\right)\left|\mathbb{F}\right|^{2}\operatorname{d}t\operatorname{d}x.$
But here we will deal with $\bar{\rho}+\sigma$ and ${\bf u}$ together by considering the symmetrized system of $\mathbf{w}$. By Itô’s formula, it holds
$\int_{U}\operatorname{d}\left(\frac{1}{2}\mathcal{D}\mathbf{w}\cdot\mathbf{w}\right)\operatorname{d}x=\int_{U}\frac{1}{2}\mathbf{w}\left(\operatorname{d}\mathcal{D}\right)\mathbf{w}\operatorname{d}x+\int_{U}\mathcal{D}\operatorname{d}\mathbf{w}\cdot\mathbf{w}\operatorname{d}x+\int_{U}\frac{1}{2}\mathcal{D}\mathbb{F}\cdot\mathbb{F}\operatorname{d}x\operatorname{d}t,$ (2.129)
which is $\operatorname{d}\int_{U}\frac{1}{2}\left(\bar{\rho}\left|{\bf u}\right|^{2}+Q^{\prime}\left(\bar{\rho}\right)\sigma^{2}\right)\operatorname{d}x$. Over the domain $U$, we integrate
$\mathcal{D}\operatorname{d}\mathbf{w}\cdot\mathbf{w}+\left(\tilde{\mathcal{A}}^{1}\mathbf{w}_{,1}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{2}\mathbf{w}_{,2}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{3}\mathbf{w}_{,3}\cdot\mathbf{w}+\tilde{\mathcal{B}}\mathbf{w}\cdot\mathbf{w}+\tilde{\mathcal{L}}_{{\bf u}}\cdot\mathbf{w}\right)\operatorname{d}t$ (2.130)
$=$ ~$\tilde{\mathcal{L}}_{\phi}\cdot\mathbf{w}\operatorname{d}t+\tilde{f}\cdot\mathbf{w},$
then we have
$\int_{U}\operatorname{d}\left(\frac{1}{2}\mathcal{D}\mathbf{w}\cdot\mathbf{w}\right)\operatorname{d}x=\int_{U}\frac{1}{2}\mathbf{w}\left(\operatorname{d}\mathcal{D}\right)\mathbf{w}\operatorname{d}x+\int_{U}\mathcal{D}\operatorname{d}\mathbf{w}\cdot\mathbf{w}\operatorname{d}x+\int_{U}\frac{1}{2}\mathcal{D}\mathbb{F}\cdot\mathbb{F}\operatorname{d}x\operatorname{d}t$
$=$ $\int_{U}\frac{1}{2}\mathbf{w}\left(\operatorname{d}\mathcal{D}\right)\mathbf{w}\operatorname{d}x-\int_{U}\left(\tilde{\mathcal{A}}^{1}\mathbf{w}_{,1}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{2}\mathbf{w}_{,2}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{3}\mathbf{w}_{,3}\cdot\mathbf{w}+\tilde{\mathcal{B}}\mathbf{w}\cdot\mathbf{w}\right)\operatorname{d}x\operatorname{d}t$ (2.131)
$-\int_{U}\tilde{\mathcal{L}}_{{\bf u}}\cdot\mathbf{w}\operatorname{d}x\operatorname{d}t+\int_{U}\tilde{\mathcal{L}}_{\phi}\cdot\mathbf{w}\operatorname{d}x\operatorname{d}t+\int_{U}\nabla h\left(\sigma\right)\cdot\mathbf{w}\operatorname{d}x\operatorname{d}t$
$+\int_{U}\mathcal{D}\mathbb{F}\operatorname{d}W\cdot\mathbf{w}\operatorname{d}x+\int_{U}\mathcal{D}\mathbb{F}\cdot\mathbb{F}\operatorname{d}x\operatorname{d}t.$
Direct calculation shows that
$-\frac{1}{2}\tilde{\mathcal{A}}^{i}_{,i}+\tilde{\mathcal{B}}-\operatorname{diag}\left[-\frac{1}{2}\left(u^{i}Q^{\prime}\left(\rho\right)\right)_{,i},-\frac{1}{2}\left(u^{i}\rho\right)_{,i},-\frac{1}{2}\left(u^{i}\rho\right)_{,i},-\frac{1}{2}\left(u^{i}\rho\right)_{,i}\right]$
is anti-symmetric [21]. Hence, we have
$\int_{U}\left(\tilde{\mathcal{A}}^{1}\mathbf{w}_{,1}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{2}\mathbf{w}_{,2}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{3}\mathbf{w}_{,3}\cdot\mathbf{w}+\tilde{\mathcal{B}}\mathbf{w}\cdot\mathbf{w}\right)\operatorname{d}x\operatorname{d}t$
$=$ $\int_{U}\left(-\frac{1}{2}\left(\mathbf{w}\tilde{\mathcal{A}}^{1}_{,1}\mathbf{w}+\mathbf{w}\tilde{\mathcal{A}}^{2}_{,2}\mathbf{w}+\mathbf{w}\tilde{\mathcal{A}}^{3}_{,3}\mathbf{w}\right)+\tilde{\mathcal{B}}\left|\mathbf{w}\right|^{2}\right)\operatorname{d}x\operatorname{d}t+\int_{U}\left(\mathbf{w}\tilde{\mathcal{A}}^{j}\mathbf{w}\right)_{,j}\operatorname{d}x\operatorname{d}t.$
On account of the insulated boundary condition ${\bf u}\cdot\nu|_{\partial U}=0$, it holds
$\int_{U}\left(\mathbf{w}\tilde{\mathcal{A}}^{j}\mathbf{w}\right)_{,j}\operatorname{d}x=\int_{\partial U}\left(\left({\bf u}\cdot\nu\right)\left(Q^{\prime}\left(\bar{\rho}\right)\sigma^{2}+\rho\left|{\bf u}\right|^{2}+2\rho Q^{\prime}\left(\bar{\rho}\right)\right)\right)\operatorname{d}S\equiv 0.$ (2.132)
Hence, it holds
$\int_{U}\left(-\frac{1}{2}\left(\mathbf{w}\tilde{\mathcal{A}}^{1}_{1}\mathbf{w}+\mathbf{w}\tilde{\mathcal{A}}^{2}_{2}\mathbf{w}+\mathbf{w}\tilde{\mathcal{A}}^{3}_{3}\mathbf{w}\right)+\tilde{\mathcal{B}}\left|\mathbf{w}\right|^{2}\right)\operatorname{d}x\operatorname{d}t\leqslant C\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}t.$ (2.133)
Recalling (2.68), there hold
$\int_{U}\tilde{\mathcal{L}}_{{\bf u}}\cdot\mathbf{w}\operatorname{d}x\operatorname{d}t=\int_{U}\left(\bar{\rho}+\sigma\right)\left|{\bf u}\right|^{2}\operatorname{d}x\operatorname{d}t\geqslant\int_{U}C\bar{\rho}\left|{\bf u}\right|^{2}\operatorname{d}x\operatorname{d}t;$ (2.134)
$\int_{U}\tilde{\mathcal{L}}_{\phi}\cdot\mathbf{w}\operatorname{d}x\operatorname{d}t=$ $\int_{U}\left(\bar{\rho}+\sigma\right)\nabla\phi\cdot{\bf u}\operatorname{d}x\operatorname{d}t=-\int_{U}\nabla\cdot\left(\left(\bar{\rho}+\sigma\right){\bf u}\right)\phi\operatorname{d}x\operatorname{d}t$ (2.135)
$=$ $\int_{U}\sigma_{t}\phi\operatorname{d}x\operatorname{d}t=\int_{U}\left(\triangle\phi\right)_{t}\phi\operatorname{d}x\operatorname{d}t=-\operatorname{d}\int_{U}\left|\nabla\phi\right|^{2}\operatorname{d}x.$
For the stochastic term, it holds
$\int_{U}\tilde{f}\cdot\mathbf{w}\operatorname{d}x=\int_{U}\left(O\left(\sigma^{2}\right)\operatorname{d}t-\mathbb{F}\operatorname{d}W\right)\cdot{\bf u}\operatorname{d}x\leqslant\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}t+\left|\int_{U}\mathbb{F}\operatorname{d}W\cdot{\bf u}\operatorname{d}x\right|.$ (2.136)
For $\left|\mathbb{F}\right|\leqslant C\left|\rho{\bf u}\right|^{2}$, we estimate
$\mathbb{E}\left[\left|\int_{0}^{t}\int_{U}\mathbb{F}\cdot{\bf u}\operatorname{d}x\operatorname{d}W\right|^{m}\right]\leqslant\mathbb{E}\left[\left(\int_{0}^{t}\left|C\int_{U}\mathbb{F}\cdot{\bf u}\operatorname{d}x\right|^{2}\operatorname{d}s\right)^{\frac{m}{2}}\right]$
$\leqslant$ $\mathbb{E}\left[\left(C\int_{0}^{t}\left|\int_{U}\left|\bar{\rho}{\bf u}\right|^{2}\left|{\bf u}\right|^{2}\cdot{\bf u}\operatorname{d}x\right|^{2}\operatorname{d}s\right)^{\frac{m}{2}}\right]\leqslant\mathbb{E}\left[\left(C\sup\limits_{s\in[0,t]}\left\|{\bf u}\right\|^{2}\int_{0}^{t}\left\|{\bf u}\right\|_{3}^{4}\operatorname{d}s\right)^{\frac{m}{2}}\right]$ (2.137)
$\leqslant$ $\delta_{3}^{m}\mathbb{E}\left[\left(\sup\limits_{s\in[0,t]}\left\|{\bf u}\right\|^{2}\right)^{m}\right]+C_{\delta_{3}}^{m}\mathbb{E}\left[\left(\int_{0}^{t}\left\|{\bf u}\right\|_{3}^{3}\operatorname{d}s\right)^{m}\right],$
where $\delta_{3}$ is taken such that $\delta_{3}^{m}\mathbb{E}\left[\left(\sup\limits_{s\in[0,t]}\left\|{\bf u}\right\|^{2}\right)^{m}\right]$ can be balanced by the left side by the time continuity of solutions. Similarly, it holds
$\int_{U}\frac{1}{2}\mathcal{D}\mathbb{F}\cdot\mathbb{F}\operatorname{d}x\operatorname{d}t\leqslant C\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}t.$ (2.138)
Besides, there holds
$\int_{U}\mathbf{w}\left(\operatorname{d}\mathcal{D}\right)\mathbf{w}\operatorname{d}x$
$=$ $\int_{U}\mathbf{w}\left(\operatorname{diag}\left\{Q^{\prime}\left(\bar{\rho}+\sigma\right)_{t},\left(\bar{\rho}+\sigma\right)_{t},\left(\bar{\rho}+\sigma\right)_{t},\left(\bar{\rho}+\sigma\right)_{t}\right\}\right)\mathbf{w}\operatorname{d}x\operatorname{d}t$
$=$ $\int_{U}\left(Q^{\prime\prime}\left(\bar{\rho}+\sigma\right)\sigma_{t}\sigma^{2}+\sigma_{t}\left|{\bf u}\right|^{2}\right)\operatorname{d}x\operatorname{d}t$ (2.139)
$=$ $\int_{U}\left(Q^{\prime\prime}\left(\bar{\rho}\right)+O\left(\sigma\right)\right)\left(-\nabla\cdot\left(\left(\bar{\rho}+\sigma\right){\bf u}\right)\right)\sigma^{2}\operatorname{d}x\operatorname{d}t+\int_{U}\left(-\nabla\cdot\left(\left(\bar{\rho}+\sigma\right){\bf u}\right)\right)\left|{\bf u}\right|^{2}\operatorname{d}x\operatorname{d}t$
$\leqslant$ $C\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}t.$
In conclusion, as $\bar{\rho}$ have a positive lower bound, we have
$\mathbb{E}\left[\left(\sup\limits_{s\in[0,t]}\int_{0}^{s}\operatorname{d}\left(\int_{U}\left|\mathbf{w}\right|^{2}\operatorname{d}x+\int_{U}\left|\nabla\phi\right|^{2}\operatorname{d}x\right)\operatorname{d}s\right)^{m}\right]+\mathbb{E}\left[\left(c_{2}\int_{0}^{t}\int_{U}\left|{\bf u}\right|^{2}\operatorname{d}x\operatorname{d}s\right)^{m}\right]$
$\leqslant$ $\mathbb{E}\left[\left(C\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right)^{m}\right].$ (2.140)
Next, we give the estimates of $\int_{0}^{t}\int_{U}\left\|\sigma\right\|^{2}\operatorname{d}x\operatorname{d}s$. From the velocity equation (2.1), we have
$\left(\nabla Q\left(\bar{\rho}+\sigma\right)-\nabla Q\left(\bar{\rho}\right)\right)\operatorname{d}t=-\operatorname{d}{\bf u}-\left(\left({\bf u}\cdot\nabla\right){\bf u}-{\bf u}\right)\operatorname{d}t+\nabla\phi\operatorname{d}t+\frac{\mathbb{F}}{\bar{\rho}+\sigma}\operatorname{d}W,$ (2.141)
with
$\nabla\left(Q\left(\bar{\rho}+\sigma\right)-Q\left(\bar{\rho}\right)\right)=Q^{\prime}\left(\bar{\rho}+\sigma\right)\nabla\sigma+Q^{\prime\prime}\left(\bar{\rho}\right)\sigma\nabla\bar{\rho}+\mathbf{h},$ (2.142)
where
$h_{i}=O\left(\sigma^{2}\right).$ (2.143)
We multiply the equation (2.141) with $\left(\sigma,\sigma,\sigma\right)^{T}$. By the integration by parts and the insulating boundary condition, due to the condition that $\left|\nabla\bar{\rho}\right|>0$, the left side is
$\int_{U}\left|Q^{\prime\prime}\left(\bar{\rho}\right)\nabla\bar{\rho}\right|\left|\sigma\right|^{2}\operatorname{d}x+\int_{U}O\left(\sigma^{3}\right)\operatorname{d}x.$ (2.144)
By Itô’s formula, there holds
$\left(\operatorname{d}u^{i}\right)\sigma=\operatorname{d}\left(u^{i}\sigma\right)-u^{i}\operatorname{d}\sigma,$ (2.145)
where
$-\int_{0}^{t}\operatorname{d}\int_{U}\left(u^{i}\sigma\right)\operatorname{d}x\leqslant\int_{0}^{t}\operatorname{d}\left(\frac{1}{2}\left\|\sigma\right\|^{2}+\frac{1}{2}\left\|u^{i}\right\|^{2}\right).$ (2.146)
By the continuity equation, it holds
$\int_{0}^{t}\int_{U}\left|u^{i}\operatorname{d}\sigma\right|\operatorname{d}x\leqslant C\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s.$ (2.147)
For $-{\bf u}\operatorname{d}t$, we directly estimate
$\int_{0}^{t}\int_{U}\left|-u^{i}\sigma\right|\operatorname{d}x\operatorname{d}s\leqslant\frac{\delta_{4}}{2}\int_{0}^{t}\left\|\sigma\right\|^{2}\operatorname{d}s+C_{\delta_{4}}\int_{0}^{t}\left\|u^{i}\right\|^{2}\operatorname{d}s,$ (2.148)
where $\delta_{4}$ is small such that $\delta_{4}\int_{0}^{t}\left\|\sigma\right\|^{2}\operatorname{d}s$ can be balanced by the left side. For the term $\nabla\phi\operatorname{d}t$ in (2.141), we estimate
$\int_{0}^{t}\int_{U}\left|-\phi_{,i}\sigma\right|\operatorname{d}x\operatorname{d}s\leqslant\frac{\delta_{4}}{2}\int_{0}^{t}\left\|\sigma\right\|^{2}\operatorname{d}s+C_{\delta_{4}}\sup_{s\in[0,t]}\left\|\phi_{,i}\right\|^{2}.$ (2.149)
For the stochastic term, since $\left|\mathbb{F}\right|\leqslant C\left|\rho{\bf u}\right|^{2}$, we estimate
$\mathbb{E}\left[\left|\int_{0}^{t}\int_{U}\frac{\mathbb{F}^{i}}{\bar{\rho}+\sigma}\operatorname{d}W\sigma\operatorname{d}x\right|^{m}\right]\leqslant\mathbb{E}\left[\left|C\int_{0}^{t}\left|\int_{U}\frac{\mathbb{F}^{i}}{\bar{\rho}+\sigma}\sigma\operatorname{d}x\right|^{2}\operatorname{d}s\right|^{\frac{m}{2}}\right]$
$\leqslant$ $\mathbb{E}\left[\left|C\int_{0}^{t}\left|\int_{U}\left|\bar{\rho}+\sigma\right|\left|{\bf u}\right|^{2}\sigma\operatorname{d}x\right|^{2}\operatorname{d}s\right|^{\frac{m}{2}}\right]\leqslant\mathbb{E}\left[\left|C\int_{0}^{t}\left\|{\bf u}\right\|^{2}\left\|\bar{\rho}\sigma+\sigma^{2}\right\|_{\infty}^{2}\operatorname{d}s\right|^{\frac{m}{2}}\right]$ (2.150)
$\leqslant$ $\mathbb{E}\left[\left(\frac{1}{4}\sup_{s\in[0,t]}\left\|{\bf u}\right\|^{2}\right)^{m}\right]+\mathbb{E}\left[\left(C\int_{0}^{t}\left\|{\bf u}\right\|^{2}\left\|\sigma\right\|_{\infty}^{2}\operatorname{d}s\right)^{m}\right]$
$+\mathbb{E}\left[\left(\frac{1}{4}\sup_{s\in[0,t]}\left\|{\bf u}\right\|^{2}\right)^{m}\right]+\mathbb{E}\left[\left(C\int_{0}^{t}\left\|{\bf u}\right\|^{2}\left\|\sigma\right\|_{\infty}^{4}\operatorname{d}s\right)^{m}\right]$
$\leqslant$ $\mathbb{E}\left[\left(\frac{1}{2}\sup_{s\in[0,t]}\left\|{\bf u}\right\|^{2}\right)^{m}\right]+\mathbb{E}\left[\left(C\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right)^{m}\right].$
Therefore, we have
$\mathbb{E}\left[\left(\int_{0}^{t}\left\|\sigma\right\|^{2}\operatorname{d}s\right)^{m}\right]$
$\leqslant$ $\mathbb{E}\left[\left(\int_{0}^{t}\operatorname{d}\left(\frac{1}{2}\left\|\sigma\right\|^{2}+\frac{1}{2}\left\|{\bf u}\right\|^{2}\right)\right)^{m}\right]+\mathbb{E}\left[\left(\frac{1}{2}\sup_{s\in[0,t]}\left\|{\bf u}\right\|^{2}\right)^{m}\right]$ (2.151)
$+\mathbb{E}\left[\left(C\sup_{s\in[0,t]}\left\|\nabla\phi\right\|^{2}\right)^{m}\right]+\mathbb{E}\left[\left(C\int_{0}^{t}\left\|{\bf u}\right\|^{2}\operatorname{d}s\right)^{m}\right]+\mathbb{E}\left[\left(C\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right)^{m}\right].$
Furthermore, we can give the estimate of $\mathbb{E}\left[\left(\int_{0}^{t}\left\|\nabla\phi\right\|^{2}\operatorname{d}s\right)^{m}\right]$. We multiply (2.141) with $\nabla\phi$ and integrate it over $U$, then we have
$\int_{U}\left|\nabla\phi\right|^{2}\operatorname{d}x\operatorname{d}t$
$=$ $-\int_{U}\left(\nabla Q\left(\bar{\rho}+\sigma\right)-\nabla Q\left(\bar{\rho}\right)\right)\cdot\nabla\phi\operatorname{d}x\operatorname{d}t+\int_{U}\operatorname{d}{\bf u}\cdot\nabla\phi\operatorname{d}x$ (2.152)
$+\int_{U}\left(\left({\bf u}\cdot\nabla\right){\bf u}-{\bf u}\right)\cdot\nabla\phi\operatorname{d}x\operatorname{d}t-\int_{U}\frac{\mathbb{F}}{\bar{\rho}+\sigma}\operatorname{d}W\cdot\nabla\phi\operatorname{d}x.$
From (2.142), by integration by parts and $\triangle\phi=\sigma$, we estimate
$\int_{0}^{t}\int_{U}\left(\nabla Q\left(\bar{\rho}+\sigma\right)-\nabla Q\left(\bar{\rho}\right)\right)\cdot\nabla\phi\operatorname{d}x\operatorname{d}s$ (2.153)
$\leqslant$ $C\left(\sup\limits_{s\in[0,t]}\left\|\nabla\phi\right\|^{2}+\int_{0}^{t}\left\|\sigma\right\|^{2}\operatorname{d}s+\int_{0}^{t}\left\|\sigma\right\|_{3}^{3}\operatorname{d}s\right).$
By Itô’s formula, there holds
$\left(\operatorname{d}{\bf u}\right)\nabla\phi=\operatorname{d}\left({\bf u}\nabla\phi\right)-{\bf u}\operatorname{d}\nabla\phi,$ (2.154)
where
$-\int_{0}^{t}\operatorname{d}\int_{U}\left({\bf u}\nabla\phi\right)\operatorname{d}x\leqslant\int_{0}^{t}\operatorname{d}\left(\frac{1}{2}\left\|\nabla\phi\right\|^{2}+\frac{1}{2}\left\|{\bf u}\right\|^{2}\right).$ (2.155)
By the continuity equation, it holds
$\int_{0}^{t}\int_{U}\left|{\bf u}\operatorname{d}\nabla\phi\right|\operatorname{d}x=\int_{0}^{t}\int_{U}\left|{\bf u}\operatorname{d}\nabla\triangle^{-1}\sigma\right|\operatorname{d}x$ (2.156)
$=$ $\int_{0}^{t}\int_{U}\left|{\bf u}\nabla\triangle^{-1}\operatorname{d}\sigma\right|\operatorname{d}x\leqslant C\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s.$
It is clear that
$\int_{0}^{t}\int_{U}\left(\left({\bf u}\cdot\nabla\right){\bf u}\right)\cdot\nabla\phi\operatorname{d}x\operatorname{d}t\leqslant C\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s.$ (2.157)
For $-{\bf u}\operatorname{d}t$, we directly estimate
$\int_{0}^{t}\int_{U}\left|-{\bf u}\cdot\nabla\phi\right|\operatorname{d}x\operatorname{d}s\leqslant\frac{1}{2}\int_{0}^{t}\left\|\nabla\phi\right\|^{2}\operatorname{d}s+\frac{1}{2}\int_{0}^{t}\left\|{\bf u}\right\|^{2}\operatorname{d}s.$ (2.158)
For the stochastic term, since $\left|\mathbb{F}\right|\leqslant C\left|\rho{\bf u}\right|^{2}$ and $\triangle\phi=\sigma$, we estimate
$\mathbb{E}\left[\left|\int_{0}^{t}\int_{U}\frac{\mathbb{F}}{\bar{\rho}+\sigma}\operatorname{d}W\cdot\nabla\phi\operatorname{d}x\right|^{m}\right]\leqslant\mathbb{E}\left[\left|C\int_{0}^{t}\left|\int_{U}\frac{\mathbb{F}}{\bar{\rho}+\sigma}\cdot\nabla\phi\operatorname{d}x\right|^{2}\operatorname{d}s\right|^{\frac{m}{2}}\right]$
$\leqslant$ $\mathbb{E}\left[\left|C\int_{0}^{t}\left|\int_{U}\left|\bar{\rho}+\sigma\right|\left|{\bf u}\right|^{2}\left|\nabla\phi\right|\operatorname{d}x\right|^{2}\operatorname{d}s\right|^{\frac{m}{2}}\right]\leqslant\mathbb{E}\left[\left|C\int_{0}^{t}\left\|{\bf u}\right\|^{2}\left\|\bar{\rho}+\sigma\right\|_{\infty}^{2}\left\|\nabla\phi\right\|_{\infty}^{2}\operatorname{d}s\right|^{\frac{m}{2}}\right]$ (2.159)
$\leqslant$ $\mathbb{E}\left[\left(\frac{1}{4}\sup_{s\in[0,t]}\left\|{\bf u}\right\|^{2}\right)^{m}\right]+\mathbb{E}\left[\left(C\int_{0}^{t}\left\|{\bf u}\right\|^{2}\left\|\sigma\right\|_{1}^{2}\operatorname{d}s\right)^{m}\right]$
$+\mathbb{E}\left[\left(\frac{1}{4}\sup_{s\in[0,t]}\left\|{\bf u}\right\|^{2}\right)^{m}\right]+\mathbb{E}\left[\left(C\int_{0}^{t}\left\|{\bf u}\right\|^{2}\left\|\sigma\right\|_{\infty}^{2}\left\|\sigma\right\|_{1}^{2}\operatorname{d}s\right)^{m}\right]$
$\leqslant$ $\mathbb{E}\left[\left(\frac{1}{2}\sup_{s\in[0,t]}\left\|{\bf u}\right\|^{2}\right)^{m}\right]+\mathbb{E}\left[\left(C\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right)^{m}\right].$
Therefore, we have
$\mathbb{E}\left[\left(\int_{0}^{t}\left\|\nabla\phi\right\|^{2}\operatorname{d}s\right)^{m}\right]$
$\leqslant$ $\mathbb{E}\left[\left(\int_{0}^{t}\operatorname{d}\left(\frac{1}{2}\left\|\nabla\phi\right\|^{2}+\frac{1}{2}\left\|{\bf u}\right\|^{2}\right)\right)^{m}\right]+\mathbb{E}\left[\left(\frac{1}{2}\sup_{s\in[0,t]}\left\|{\bf u}\right\|^{2}\right)^{m}\right]$ (2.160)
$+\mathbb{E}\left[\left(C\sup_{s\in[0,t]}\left\|\nabla\phi\right\|^{2}\right)^{m}\right]+\mathbb{E}\left[\left(C\int_{0}^{t}\left\|\mathbf{w}\right\|^{2}\operatorname{d}s\right)^{m}\right]+\mathbb{E}\left[\left(C\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right)^{m}\right].$
Multiplying a small constant to it, and we plus the zero-order estimates (2.2.1) such that
$\mathbb{E}\left[\left(\sup\limits_{s\in[0,t]}\int_{0}^{s}\operatorname{d}\left(\frac{1}{2}\left\|\sigma\right\|^{2}+\left\|{\bf u}\right\|^{2}+\frac{1}{2}\left\|\nabla\phi\right\|^{2}\right)\right)^{m}\right]+\mathbb{E}\left[\left(\frac{1}{2}\sup_{s\in[0,t]}\left\|{\bf u}\right\|^{2}\right)^{m}\right]$ (2.161)
$+\mathbb{E}\left[\left(C\sup_{s\in[0,t]}\left\|\nabla\phi\right\|^{2}\right)^{m}\right]+\mathbb{E}\left[\left(C\int_{0}^{t}\left\|\mathbf{w}\right\|^{2}\operatorname{d}s\right)^{m}\right]$
can be balanced by (2.2.1). Then we obtain
$\mathbb{E}\left[\left(\sup\limits_{s\in[0,t]}\int_{0}^{s}\operatorname{d}\left(\left\|\mathbf{w}\right\|^{2}+\left\|\nabla\phi\right\|^{2}\right)+\zeta_{5}\left(\int_{0}^{t}\left\|\mathbf{w}\right\|^{2}+\left\|\nabla\phi\right\|^{2}\right)\operatorname{d}s\right)^{m}\right]$ (2.162)
$\leqslant$ $\mathbb{E}\left[\left(C\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right)^{m}\right].$
where $C$ depends on $m$.
2.2.2. First order estimates
Taking derivative to (2.39), we have
$\nabla\left(\mathcal{D}\operatorname{d}\mathbf{w}\right)+\nabla\left(\left(\tilde{\mathcal{A}}^{1}\mathbf{w}_{,1}+\tilde{\mathcal{A}}^{2}\mathbf{w}_{,2}+\tilde{\mathcal{A}}^{3}\mathbf{w}_{,3}\right)+\nabla\left(\tilde{\mathcal{B}}\mathbf{w}\right)+\nabla\left(\tilde{\mathcal{L}}_{{\bf u}}\right)\right)\operatorname{d}t$ (2.163)
=(~$=\nabla\left(\tilde{\mathcal{L}}_{\phi}\right)\operatorname{d}t+\nabla\tilde{f}.$
Recalling $\mathcal{D}=\operatorname{diag}\left[Q^{\prime}\left(\bar{\rho}+\sigma\right),\bar{\rho}+\sigma,\bar{\rho}+\sigma,\bar{\rho}+\sigma\right]$, we calculate
$\nabla\left(\mathcal{D}\operatorname{d}\mathbf{w}\right)=$ $\partial_{j}\left(\begin{array}[]{c}Q^{\prime}\left(\bar{\rho}+\sigma\right)\sigma_{t}\\ \left(\bar{\rho}+\sigma\right)\operatorname{d}u^{1}\\ \left(\bar{\rho}+\sigma\right)\operatorname{d}u^{2}\\ \left(\bar{\rho}+\sigma\right)\operatorname{d}u^{3}\end{array}\right)=\left(\begin{array}[]{c}\partial_{j}Q^{\prime}\left(\bar{\rho}+\sigma\right)\sigma_{t}+Q^{\prime}\left(\bar{\rho}+\sigma\right)\partial_{j}\sigma_{t}\\ \partial_{j}\left(\bar{\rho}+\sigma\right)\operatorname{d}u^{1}+\left(\bar{\rho}+\sigma\right)\partial_{j}\operatorname{d}u^{1}\\ \partial_{j}\left(\bar{\rho}+\sigma\right)\operatorname{d}u^{2}+\left(\bar{\rho}+\sigma\right)\partial_{j}\operatorname{d}u^{2}\\ \partial_{j}\left(\bar{\rho}+\sigma\right)\operatorname{d}u^{3}+\left(\bar{\rho}+\sigma\right)\partial_{j}\operatorname{d}u^{3}\end{array}\right)$ (2.172)
$=$ $\nabla\cdot\mathcal{D}\operatorname{d}\mathbf{w}+\mathcal{D}\nabla\operatorname{d}\mathbf{w},$
$\nabla\left(\tilde{\mathcal{A}}^{i}\mathbf{w}_{,i}\right)=\partial_{l}\left(\tilde{\mathcal{A}}^{i}_{jk}w_{k,i}\right)=\partial_{l}\tilde{\mathcal{A}}^{i}_{jk}w_{k,i}+\tilde{\mathcal{A}}^{i}_{jk}\partial_{l}w_{k,i}=\nabla\tilde{\mathcal{A}}^{i}\mathbf{w}_{,i}+\tilde{\mathcal{A}}^{i}\nabla\mathbf{w}_{,i},$ (2.173)
(~$\nabla\left(\tilde{\mathcal{B}}\mathbf{w}\right)=\partial_{l}\left(\tilde{\mathcal{B}}_{jk}w_{k}\right)=\partial_{l}\tilde{\mathcal{B}}_{jk}w_{k}+\tilde{\mathcal{B}}_{jk}\partial_{l}w_{k}=\nabla\tilde{\mathcal{B}}\mathbf{w}+\tilde{\mathcal{B}}\nabla\mathbf{w},$ (2.174)
(~$\nabla\left(\tilde{\mathcal{L}}_{{\bf u}}\right)=\left[\begin{array}[]{c}0\\ \nabla{\bf u}\end{array}\right],$ (2.177)
(~$\nabla\left(\tilde{\mathcal{L}}_{\phi}\right)=\left[\begin{array}[]{c}0\\ \nabla\left(\left(\bar{\rho}+\sigma\right)\nabla\phi\right)\end{array}\right],$ (2.180)
and
$\nabla\left(\tilde{f}\right)=\left[\begin{array}[]{c}0\\ \nabla\left(O\left(\sigma^{2}\right)-\mathbb{F}\operatorname{d}W\right)\end{array}\right]=\left[\begin{array}[]{c}0\\ O\left(\sigma\nabla\sigma\right)\operatorname{d}t-\nabla\mathbb{F}\operatorname{d}W\end{array}\right].$ (2.185)
Hence (2.163) is deduced to
$\operatorname{d}\mathbf{w}\nabla\cdot\mathcal{D}+\mathcal{D}\nabla\operatorname{d}\mathbf{w}+\left(\nabla\tilde{\mathcal{A}}^{i}\mathbf{w}_{,i}+\tilde{\mathcal{A}}^{i}\nabla\mathbf{w}_{,i}+\nabla\tilde{\mathcal{B}}\mathbf{w}+\tilde{\mathcal{B}}\nabla\mathbf{w}+\nabla{\bf u}\right)\operatorname{d}t$
$=$ $\left[\begin{array}[]{c}0\\ \nabla\left(\left(\bar{\rho}+\sigma\right)\nabla\phi\right)\end{array}\right]\operatorname{d}t+\left[\begin{array}[]{c}0\\ O\left(\sigma\nabla\sigma\right)\end{array}\right]\operatorname{d}t-\left[\begin{array}[]{c}0\\ \nabla\mathbb{F}\operatorname{d}W\end{array}\right].$ (2.192)
Multiplying (2.2.2) with $\nabla\mathbf{w}$, and integrating it on $U$, we have
$\int_{U}\frac{1}{2}\mathcal{D}\operatorname{d}\nabla\mathbf{w}:\nabla\mathbf{w}\operatorname{d}x+\int_{U}\tilde{\mathcal{A}}^{i}\partial_{i}\left(\left|\nabla\mathbf{w}\right|^{2}\right)\operatorname{d}x\operatorname{d}t+\int_{U}\tilde{\mathcal{B}}\left|\nabla\mathbf{w}\right|^{2}\operatorname{d}x\operatorname{d}t+\int_{U}\left|\nabla{\bf u}\right|^{2}\operatorname{d}x\operatorname{d}t$
$=$ $\int_{U}\left[\begin{array}[]{c}0\\ \nabla\left(\left(\bar{\rho}+\sigma\right)\nabla\phi\right)\end{array}\right]:\nabla\mathbf{w}\operatorname{d}x\operatorname{d}t-\int_{U}\nabla\tilde{\mathcal{A}}^{i}\mathbf{w}_{,i}:\nabla\mathbf{w}\operatorname{d}x\operatorname{d}t$ (2.195)
$-\int_{U}\nabla\tilde{\mathcal{B}}\mathbf{w}:\nabla\mathbf{w}\operatorname{d}x\operatorname{d}t-\int_{U}\operatorname{d}\mathbf{w}\nabla\cdot\mathcal{D}:\nabla\mathbf{w}\operatorname{d}x+\int_{U}\left[\begin{array}[]{c}0\\ O\left(\sigma\nabla\sigma\right)\end{array}\right]:\nabla\mathbf{w}\operatorname{d}x\operatorname{d}t$ (2.198)
$-\int_{U}\left[\begin{array}[]{c}0\\ \nabla\mathbb{F}\operatorname{d}W\end{array}\right]:\nabla\mathbf{w}\operatorname{d}x.$ (2.201)
Since $\sigma_{t}=-\nabla\cdot\left(\left(\bar{\rho}+\sigma\right){\bf u}\right)$, we estimate
$\int_{U}\frac{1}{2}\mathcal{D}_{t}\nabla\mathbf{w}:\nabla\mathbf{w}\operatorname{d}x\operatorname{d}t$
$=$ $\int_{U}\nabla\mathbf{w}\operatorname{diag}\left[\left(Q^{\prime}\left(\bar{\rho}+\sigma\right)\right)_{t},\left(\bar{\rho}+\sigma\right)_{t},\left(\bar{\rho}+\sigma\right)_{t},\left(\bar{\rho}+\sigma\right)_{t}\right]\nabla\mathbf{w}\operatorname{d}x\operatorname{d}t$
$=$ $\int_{U}\left(\sigma_{t}Q^{\prime\prime}\left(\bar{\rho}+\sigma\right)\sigma^{2}+\sigma_{t}\left|\nabla{\bf u}\right|^{2}\right)\operatorname{d}x\operatorname{d}t$ (2.202)
$=$ $\int_{U}\left(-Q^{\prime\prime}\left(\bar{\rho}+\sigma\right)\nabla\cdot\left(\left(\bar{\rho}+\sigma\right){\bf u}\right)\sigma^{2}-\nabla\cdot\left(\left(\bar{\rho}+\sigma\right){\bf u}\right)\left|\nabla{\bf u}\right|^{2}\right)\operatorname{d}x\operatorname{d}t$
$\leqslant$ $C\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}t.$
Due to the boundary conditions ${\bf u}\cdot\nu=0$, it follows that
$\int_{U}\tilde{\mathcal{A}}^{i}\partial_{i}\left(\left|\nabla\mathbf{w}\right|^{2}\right)\operatorname{d}x+\int_{U}\tilde{\mathcal{B}}\left|\nabla\mathbf{w}\right|^{2}\operatorname{d}x=0,$ (2.203)
and
$\int_{0}^{t}\int_{U}\left[\begin{array}[]{c}0\\ \nabla\left(\left(\bar{\rho}+\sigma\right)\nabla\phi\right)\end{array}\right]:\nabla\mathbf{w}\operatorname{d}x\operatorname{d}s-\int_{0}^{t}\left(\int_{U}\nabla\tilde{\mathcal{A}}^{i}\mathbf{w}_{,i}:\nabla\mathbf{w}\operatorname{d}x+\int_{U}\nabla\tilde{\mathcal{B}}\mathbf{w}:\nabla\mathbf{w}\operatorname{d}x\right)\operatorname{d}s$ (2.206)
$\leqslant$ $C\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s+\frac{\delta_{5}}{3}\sup\limits_{s\in[0,t]}\left\|\nabla{\bf u}\right\|^{2}+C_{\delta_{5}}\int_{0}^{t}\left\|\nabla\phi\right\|^{2}\operatorname{d}s,$ (2.207)
where $\int_{0}^{t}\left\|\nabla\phi\right\|^{2}\operatorname{d}s=\int_{0}^{t}\left\|\tilde{e}\right\|^{2}\operatorname{d}s$ can be bounded by $\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s$ from the zeroth-order energy estimates, $\delta_{5}$ being determined later. Similarly, we estimate
$\mathbb{E}\left[\left|\int_{0}^{t}\int_{U}\nabla\mathbb{F}\operatorname{d}W:\nabla\mathbf{w}\operatorname{d}x\right|^{m}\right]\leqslant\mathbb{E}\left[\left|\frac{\delta_{5}}{3}\sup\limits_{s\in[0,t]}\left\|{\bf u}\right\|^{2}\right|^{m}\right]+\mathbb{E}\left[\left|C_{\delta_{5}}\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right|^{m}\right].$ (2.208)
From (2.5), we have
$-\int_{U}\operatorname{d}\mathbf{w}\nabla\mathcal{D}:\nabla\mathbf{w}\operatorname{d}x$
$=$ $-\int_{U}\left(\mathcal{A}^{1}\mathbf{w}_{,1}+\mathcal{A}^{2}\mathbf{w}_{,2}+\mathcal{A}^{3}\mathbf{w}_{,3}+\mathcal{B}\mathbf{w}+\mathcal{L}_{{\bf u}}\right)\nabla\mathcal{D}:\nabla\mathbf{w}\operatorname{d}x\operatorname{d}t$ (2.209)
$+\int_{U}\mathcal{L}_{\mathcal{\phi}}\nabla\mathcal{D}:\nabla\mathbf{w}\operatorname{d}x\operatorname{d}t+\int_{U}\left(\mathcal{L}_{\mathcal{\phi}}+O\left(\sigma^{2}\right)\right)\nabla\mathcal{D}:\nabla\mathbf{w}\operatorname{d}x\operatorname{d}t-\int_{U}\mathbb{F}\operatorname{d}W\nabla\mathcal{D}:\nabla\mathbf{w}\operatorname{d}x$
$\leqslant$ $C\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}t+\int_{U}\mathbb{F}\operatorname{d}W\nabla\mathcal{D}:\nabla\mathbf{w}\operatorname{d}x.$
Similarly, we have
$\mathbb{E}\left[\left|\int_{0}^{t}\int_{U}\mathbb{F}\operatorname{d}W\nabla\mathcal{D}:\nabla\mathbf{w}\operatorname{d}x\right|^{m}\right]$ (2.210)
$\leqslant$ $\mathbb{E}\left[\left(\frac{\delta_{5}}{3}\sup\limits_{s\in[0,t]}\left\|{\bf u}\right\|^{2}\right)^{m}\right]+\mathbb{E}\left[\left(C_{\delta_{5}}\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right)^{m}\right].$ (2.211)
We take $\delta_{5}$ such that $\frac{\delta_{5}}{3}\left\|{\bf u}\right\|^{2}$ and $\frac{\delta_{5}}{3}\left\|\nabla{\bf u}\right\|^{2}$ can be balanced by the left side of energy estimates. Similar as the estimates for (2.2.1), we have the estimate of $\int_{0}^{t}\left\|\nabla\sigma\right\|^{2}\operatorname{d}s$. In conclusion, we have
$\mathbb{E}\left[\left|\sup\limits_{s\in[0,t]}\int_{0}^{s}\operatorname{d}\left(\frac{1}{2}\int_{U}\mathcal{D}\nabla\mathbf{w}:\nabla\mathbf{w}\operatorname{d}x\right)\right|^{m}\right]+\mathbb{E}\left[\left|c_{4}\int_{0}^{t}\int_{U}\left|\nabla\mathbf{w}\right|^{2}\operatorname{d}x\operatorname{d}s\right|^{m}\right]$ (2.212)
$\leqslant$ $\mathbb{E}\left[\left|\int_{0}^{t}C\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right|^{m}\right],$
where $C$ is independent on $t$.
2.2.3. Second order estimates
We write (2.5) in the form of components, and the $i-$th equation is
$d_{i}\operatorname{d}w_{i}+\left(\left(\tilde{\mathcal{A}}^{1}\right)_{ij}w_{j,1}+\left(\tilde{\mathcal{A}}^{2}\right)_{ij}w_{j,2}+\left(\tilde{\mathcal{A}}^{3}\right)_{ij}w_{j,3}+\left(\tilde{\mathcal{B}}\right)_{ij}w_{j}+d_{i}w_{i}\right)\operatorname{d}t$ (2.213)
$=$ $\left(d_{i}\phi_{,i}+h\left(\sigma\right)_{i}\right)\operatorname{d}t-\mathbb{F}_{i}\operatorname{d}W.$
Taking the second-order derivatives, we have
$\partial_{kl}^{2}\left(d_{i}\operatorname{d}w_{i}+\left(\left(\tilde{\mathcal{A}}^{1}\right)_{ij}w_{j,1}+\left(\tilde{\mathcal{A}}^{2}\right)_{ij}w_{j,2}+\left(\tilde{\mathcal{A}}^{3}\right)_{ij}w_{j,3}+\left(\tilde{\mathcal{B}}\right)_{ij}w_{j}+d_{i}w_{i}\right)\operatorname{d}t\right)$ (2.214)
$=\partial_{kl}^{2}\left(\left(d_{i}\phi_{,i}+h\left(\sigma\right)_{i}\right)\operatorname{d}t-\mathbb{F}_{i}\operatorname{d}W\right).$
Multiplying (2.214) with $\partial_{k}\partial_{l}w_{i}$ and integrating it over $U$, we have
$\int_{U}d_{i}\partial_{k}\partial_{l}\operatorname{d}w_{i}\partial_{k}\partial_{l}w_{i}\operatorname{d}x=\int_{U}d_{i}\operatorname{d}\left|\partial_{k}\partial_{l}w_{i}\right|^{2}\operatorname{d}x,$ (2.215)
By the insulated boundary condition ${\bf u}\cdot\nu=0$, for all $i,j$, there holds
$\int_{U}\partial_{k}\partial_{l}w_{j}\left(-\frac{1}{2}\left(\left(\tilde{\mathcal{A}}^{1}\right)_{ij,1}+\left(\tilde{\mathcal{A}}^{2}\right)_{ij,2}+\left(\tilde{\mathcal{A}}^{3}\right)_{ij,3}\right)+\left(\tilde{\mathcal{B}}\right)_{ij}\right)\partial_{k}\partial_{l}w_{i}\operatorname{d}x=0.$ (2.216)
By(2.5), the integral in the deterministic terms are bounded by $C\left\|\mathbf{w}\right\|_{3}^{3}$. The stochastic term is estimated as follows
$\mathbb{E}\left[\left|\int_{0}^{t}\int_{U}\partial_{kl}^{2}\mathbb{F}_{i}\partial_{k}\partial_{l}w_{i}\operatorname{d}x\operatorname{d}W\right|^{m}\right]$ (2.217)
$\leqslant$ $\mathbb{E}\left[\left(\delta_{6}\sup\limits_{s\in[0,t]}\left\|\nabla\mathbf{w}\right\|^{2}\right)^{m}\right]+\mathbb{E}\left[\left(C_{\delta_{6}}\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right)^{m}\right],$
where $\delta_{6}$ is taken such that $\delta_{6}\sup\limits_{s\in[0,t]}\left\|\nabla\mathbf{w}\right\|^{2}$ can be obtained by left side in first-order estimates. Similar to the estimates (2.2.1), we have the estimates for $\int_{0}^{t}\int_{U}\left|\partial^{2}\sigma\right|^{2}\operatorname{d}x\operatorname{d}s$. Taking the sum over all the index $i=1,2,3,4$, we have
$\mathbb{E}\left[\left|\sup\limits_{s\in[0,t]}\int_{0}^{s}\operatorname{d}\left(\int_{U}\frac{1}{2}\left|\partial^{2}\mathbf{w}\right|^{2}\operatorname{d}x\right)\right|^{m}\right]+\mathbb{E}\left[\left|c_{5}\int_{0}^{t}\int_{U}\left|\partial^{2}\mathbf{w}\right|^{2}\operatorname{d}x\operatorname{d}s\right|^{m}\right]$ (2.218)
$\leqslant$ $\mathbb{E}\left[\left|\int_{0}^{t}C\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right|^{m}\right],$
with the assumption that $\bar{\rho}$ have a positive lower bound, where $C$ is independent on $t$.
2.2.4. Third-order estimates
Considering the $3$-order estimates, we take an additional derivative of (2.214). Repeating the argument in subsection 2.2.3, we have
$\mathbb{E}\left[\left|\sup\limits_{s\in[0,t]}\int_{0}^{s}\operatorname{d}\left(\frac{1}{2}\int_{U}\mathcal{D}\nabla\mathbf{w}:\nabla\mathbf{w}\operatorname{d}x\right)\right|^{m}\right]+\mathbb{E}\left[\left|c_{4}\int_{0}^{t}\int_{U}\left|\nabla\mathbf{w}\right|^{2}\operatorname{d}x\operatorname{d}s\right|^{m}\right]$ (2.219)
$\leqslant$ $\mathbb{E}\left[\left|\int_{0}^{t}C\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right|^{m}\right],$
where $C$ is independent on $t$.
$\mathbb{E}\left[\left|\sup\limits_{s\in[0,t]}\int_{0}^{s}\operatorname{d}\left(\int_{U}\frac{1}{2}\left|\partial^{3}\mathbf{w}\right|^{2}\operatorname{d}x\right)\right|^{m}\right]+\mathbb{E}\left[\left|c_{6}\int_{0}^{t}\int_{U}\left|\partial^{3}\mathbf{w}\right|^{2}\operatorname{d}x\operatorname{d}s\right|^{m}\right]$ (2.220)
$\leqslant$ $\mathbb{E}\left[\left|\int_{0}^{t}C\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right|^{m}\right],$
with the assumption that $\bar{\rho}$ have a positive lower bound.
Step 5: Global existence.
2.3. Global existence
In this subsection, we show the global existence for both cases on stochastic forces under (1.20) and general forces.
2.3.1. For stochastic forces under (1.20) and small perturbation for initial data (1.21)
We combine the energy estimates up to third order. Then, the assumption that $\bar{\rho}$ have a positive lower bound, leads to the following inequality:
$\mathbb{E}\left[\left|\sup\limits_{s\in[0,t]}\left(\left\|\mathbf{w}\right\|_{3}^{2}(s)+\left\|\nabla\phi\right\|^{2}(s)\right)+\alpha\int_{0}^{t}\left(\left\|\mathbf{w}\right\|_{3}^{2}+\left\|\nabla\phi\right\|^{2}\right)(s)\operatorname{d}s\right|^{m}\right]$ (2.221)
$\leqslant$ $\mathbb{E}\left[\left|C\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right|^{m}\right]+\mathbb{E}\left[\left(C\left(\left\|\mathbf{w}_{0}\right\|_{3}^{2}+\left\|\nabla\phi_{0}\right\|^{2}\right)\right)^{m}\right],$
where $\alpha\leqslant c_{i},i=1,\cdots,6$, and $C$ depends on $\bar{\rho}$, $m$ and the domain $U$, but is independent on $t$. Since $\left\|\mathbf{w}\right\|_{3}$ is small, we have
$\mathbb{E}\left[\sup\limits_{s\in[0,t]}\left(\left|\left\|\mathbf{w}\right\|_{3}^{2}+\left\|\nabla\phi\right\|^{2}\right)+\alpha\int_{0}^{t}\left(\left\|\mathbf{w}\right\|_{3}^{2}+\left\|\nabla\phi\right\|^{2}\right)(s)\operatorname{d}s-C\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right|^{m}\right]$ (2.222)
$\leqslant$ $\mathbb{E}\left[\left(C\left(\left\|\mathbf{w}_{0}\right\|_{3}^{2}+\left\|\nabla\phi_{0}\right\|^{2}\right)\right)^{m}\right],$
and,
$\mathbb{E}\left[\left(\sup\limits_{s\in[0,t]}\left(\left\|\mathbf{w}\right\|_{3}^{2}+\left\|\nabla\phi\right\|^{2}\right)\right)^{m}\right]\leqslant\mathbb{E}\left[\left(C\left(\left\|\mathbf{w}_{0}\right\|_{3}^{2}+\left\|\nabla\phi_{0}\right\|^{2}\right)\right)^{m}\right],$ (2.223)
where $C$ is independent on $t$. With the above uniform estimates for any time $t$, and the local existence on $\left[0,T_{1}\right]$, we can extend the existence to $\left[T_{1},T_{1}+\tilde{T}\right]$, and extend to any time $T_{1}+k\tilde{T},\forall~{}k\in\mathds{N}^{+}$. More specifically, for the estimate of onto mapping, if
$\mathbb{E}\left[\left(\sup\limits_{s\in[T_{1},t]}\left\|\mathbf{w}_{n-1}(s)\right\|_{3}^{2}\right)^{m}\right]\leqslant$ $4\mathbb{E}\left[\left(\left(\left\|\mathbf{w}(T_{1})\right\|_{3}^{2}+\left\|\nabla\phi(T_{1})\right\|^{2}\right)\right)^{m}\right]$ (2.224)
$\leqslant$ $4\mathbb{E}\left[\left(C\left(\left\|\mathbf{w}_{0}\right\|_{3}^{2}+\left\|\nabla\phi_{0}\right\|^{2}\right)\right)^{m}\right],$
then
$\mathbb{E}\left[\left(\sup\limits_{s\in[T_{1},t]}\left\|\mathbf{w}_{n}(s)\right\|_{3}^{2}\right)^{m}\right]\leqslant 4\mathbb{E}\left[\left(C\left(\left\|\mathbf{w}_{0}\right\|_{3}^{2}+\left\|\nabla\phi_{0}\right\|^{2}\right)\right)^{m}\right].$ (2.225)
Similarly, the contraction holds from $T_{1}$ to $T_{1}+\tilde{T}$. Then the existence is extended to $T_{1}+k\tilde{T}$ for any $k\in\mathds{N}^{+}$. In conclusion, we obtain the global existence of $\mathbf{w}$ and $\phi$, which is equivalent to the global existence of strong solutions $(\rho,{\bf u},\Phi)$ stated by the following proposition.
Proposition 2.1.
In $(\Omega,\mathcal{F},\mathbb{P})$, there exists a unique global-in-time strong solution $(\rho,{\bf u},\Phi)$ to (1.1):
$\rho,{\bf u}\in C\left([0,T];H^{3}\left(U\right)\right),\Phi\in C\left([0,T];H^{5}\left(U\right)\right),\forall~{}T>0,$ (2.226)
up to a modification, where $m\geqslant 2$ is a constant.
2.3.2. For general stochastic forces
If the stochastic forces has linear growth in $\rho{\bf u}$, then the following energy estimates hold
$\mathbb{E}\left[\left|\sup\limits_{s\in[0,t]}\int_{0}^{s}\operatorname{d}\left(\left\|\mathbf{w}\right\|_{3}^{2}(s)+\left\|\nabla\phi\right\|^{2}(s)\right)+\alpha\int_{0}^{t}\int_{U}\left(\left\|\mathbf{w}\right\|_{3}^{2}(s)+\left\|\nabla\phi\right\|^{2}(s)\right)\operatorname{d}x\operatorname{d}s\right|^{m}\right]$
$\leqslant$ $\mathbb{E}\left[\left|\int_{0}^{t}C\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right|^{m}\right]+\mathbb{E}\left[\left|\int_{0}^{t}C\left\|\mathbf{w}\right\|_{3}^{2}\operatorname{d}s\right|^{m}\right].$ (2.227)
Without the small perturbation of initial data (1.21), we can use the generalized Grönwall’s inequality to obtain
$\mathbb{E}\left[\left(\sup\limits_{s\in[0,t]}\left(\left\|\mathbf{w}\right\|_{3}^{2}+\left\|\nabla\phi\right\|^{2}\right)\right)^{m}\right]\leqslant\mathbb{E}\left[\left(C(t)\left(\left\|\mathbf{w}_{0}\right\|_{3}^{2}+\left\|\nabla\phi_{0}\right\|^{2}\right)\right)^{m}\right],$ (2.228)
where $C(t)$ is increasing with respect to $t$. Similarly, if the stochastic forces have cubic growth in $\rho{\bf u}$, then the energy estimates become
$\mathbb{E}\left[\left|\sup\limits_{s\in[0,t]}\int_{0}^{s}\operatorname{d}\left(\left\|\mathbf{w}\right\|_{3}^{2}(s)+\left\|\nabla\phi\right\|^{2}(s)\right)+\alpha\int_{0}^{t}\int_{U}\left(\left\|\mathbf{w}\right\|_{3}^{2}(s)+\left\|\nabla\phi\right\|^{2}(s)\right)\operatorname{d}x\operatorname{d}s\right|^{m}\right]$
$\leqslant$ $\mathbb{E}\left[\left|\int_{0}^{t}C\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right|^{m}\right]+\mathbb{E}\left[\left|\int_{0}^{t}C\left\|\mathbf{w}\right\|_{3}^{4}\operatorname{d}s\right|^{m}\right].$ (2.229)
By the generalized Grönwall’s inequality, there also holds (2.228). Hence, for the smooth $Y$ in (1.4) and can be bounded by the homogeneous polynomials, the estimates of (2.228) holds as well. For the estimate of onto mapping, for any fixed $T$, $t\in[0,T]$, if
$\mathbb{E}\left[\left(\sup\limits_{s\in[T_{1},t]}\left\|\mathbf{w}_{n-1}(s)\right\|_{3}^{2}\right)^{m}\right]\leqslant$ $4\mathbb{E}\left[\left(\left(\left\|\mathbf{w}(T_{1})\right\|_{3}^{2}+\left\|\nabla\phi(T_{1})\right\|^{2}\right)\right)^{m}\right]$ (2.230)
$\leqslant$ $4\mathbb{E}\left[\left(C(T)\left(\left\|\mathbf{w}_{0}\right\|_{3}^{2}+\left\|\nabla\phi_{0}\right\|^{2}\right)\right)^{m}\right],$
then
$\mathbb{E}\left[\left(\sup\limits_{s\in[T_{1},t]}\left\|\mathbf{w}_{n}(s)\right\|_{3}^{2}\right)^{m}\right]\leqslant 4\mathbb{E}\left[\left(C(T)\left(\left\|\mathbf{w}_{0}\right\|_{3}^{2}+\left\|\nabla\phi_{0}\right\|^{2}\right)\right)^{m}\right].$ (2.231)
Thus, we extend the local existence on $\left[0,T_{1}\right]$ to $\left[0,T_{1}+\tilde{T}\right]$, and to $\left[0,T_{1}+k\tilde{T}\right],\forall~{}k\in\mathds{N}^{+}$. By Zorn’s lemma, the global existence holds.
3. Asymptotic stability of solutions
In this section, we consider the stability under the assumptions of (1.20) and (1.21). The a priori estimates (2.221) shows the stability of solutions around the steady state. However, (2.221) is insufficient for investigating the decay rate since the a priori estimates are already in the form of time integrals rather than a differential inequality. Integrating twice with respect time might not be wise as it could lead to disappearance of the favorable temporal properties. The asymptotic decay of solution is then derived from the following weighted estimates up to second-order. To manipulate the weighted energy estimates for stochastic system, we need multiply $\operatorname{d}\left(\frac{1}{2}\mathcal{D}\mathbf{w}\cdot\mathbf{w}\right)$ directly with $e^{\alpha t}$ first, where $\alpha$ is in (2.221). Then we integrate it with respect to $x$, $t$, and $\omega$, to estimate the time integral.
3.1. Weighted decay estimates
3.1.1. Zeroth-order weighted estimates
We multiply (2.2.1) with $e^{\alpha t}$, then we have
$\int_{U}e^{\alpha t}\operatorname{d}\left(\frac{1}{2}\mathcal{D}\mathbf{w}\cdot\mathbf{w}\right)\operatorname{d}x$
$=$ $e^{\alpha t}\int_{U}\frac{1}{2}\mathbf{w}\left(\operatorname{d}\mathcal{D}\right)\mathbf{w}\operatorname{d}x-e^{\alpha t}\int_{U}\left(\tilde{\mathcal{A}}^{1}\mathbf{w}_{,1}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{2}\mathbf{w}_{,2}\cdot\mathbf{w}+\tilde{\mathcal{A}}^{3}\mathbf{w}_{,3}\cdot\mathbf{w}+\tilde{\mathcal{B}}\mathbf{w}\cdot\mathbf{w}\right)\operatorname{d}x\operatorname{d}t$
$-e^{\alpha t}\int_{U}\tilde{\mathcal{L}}_{{\bf u}}\cdot\mathbf{w}\operatorname{d}x\operatorname{d}t+e^{\alpha t}\int_{U}\tilde{\mathcal{L}}_{\phi}\cdot\mathbf{w}\operatorname{d}x\operatorname{d}t+e^{\alpha t}\int_{U}\nabla h\left(\sigma\right)\cdot\mathbf{w}\operatorname{d}x\operatorname{d}t$ (3.1)
$+e^{\alpha t}\int_{U}\mathcal{D}\mathbb{F}\operatorname{d}W\cdot\mathbf{w}\operatorname{d}x+e^{\alpha t}\int_{U}\mathcal{D}\mathbb{F}\cdot\mathbb{F}\operatorname{d}x\operatorname{d}t.$
From the estimates of zeroth-order estimates in subsection 2.2, we conclude the following estimates omitting detailed calculation:
$\int_{U}e^{\alpha t}\left(-\frac{1}{2}\left(\mathbf{w}\tilde{\mathcal{A}}^{1}_{1}\mathbf{w}+\mathbf{w}\tilde{\mathcal{A}}^{2}_{2}\mathbf{w}+\mathbf{w}\tilde{\mathcal{A}}^{3}_{3}\mathbf{w}\right)+\tilde{\mathcal{B}}\left|\mathbf{w}\right|^{2}\right)\operatorname{d}x\operatorname{d}t\leqslant Ce^{\alpha t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}t;$ (3.2)
$e^{\alpha t}\int_{U}\tilde{\mathcal{L}}_{{\bf u}}\cdot\mathbf{w}\operatorname{d}x\operatorname{d}t\geqslant e^{\alpha t}\int_{U}C\bar{\rho}\left|{\bf u}\right|^{2}\operatorname{d}x\operatorname{d}t\geqslant\alpha e^{\alpha t}\int_{U}C\bar{\rho}\left|{\bf u}\right|^{2}\operatorname{d}x\operatorname{d}t;$ (3.3)
$e^{\alpha t}\int_{U}\tilde{\mathcal{L}}_{\phi}\cdot\mathbf{w}\operatorname{d}x\operatorname{d}t=-e^{\alpha t}\operatorname{d}\int_{U}\left|\nabla\phi\right|^{2}\operatorname{d}x;$ (3.4)
$e^{\alpha t}\int_{U}\mathbf{w}\left(\operatorname{d}\mathcal{D}\right)\mathbf{w}\operatorname{d}x\leqslant Ce^{\alpha t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}t;$ (3.5)
$e^{\alpha t}\int_{U}\frac{1}{2}\mathcal{D}\mathbb{F}\cdot\mathbb{F}\operatorname{d}x\operatorname{d}t\leqslant Ce^{\alpha t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}t.$ (3.6)
For the estimates of stochastic integral, it holds
$e^{\alpha t}\int_{U}\tilde{f}\cdot\mathbf{w}\operatorname{d}x\leqslant e^{\alpha t}\left\|\mathbf{w}\right\|_{3}^{3}+e^{\alpha t}\left|\int_{U}\mathbb{F}\operatorname{d}W\cdot{\bf u}\operatorname{d}x\right|.$ (3.7)
For $\left|\mathbb{F}\right|\leqslant C\left|\rho{\bf u}\right|^{2}$,
$\mathbb{E}\left[\left|\int_{0}^{t}e^{\alpha s}\int_{U}\mathbb{F}\cdot{\bf u}\operatorname{d}x\operatorname{d}W\right|^{m}\right]\leqslant\mathbb{E}\left[\left(C\int_{0}^{t}e^{2\alpha s}\left|\int_{U}\mathbb{F}\cdot{\bf u}\operatorname{d}x\right|^{2}\operatorname{d}s\right)^{\frac{m}{2}}\right]$
$\leqslant$ $\mathbb{E}\left[\left(C\int_{0}^{t}e^{2\alpha s}\left|\int_{U}\left|\bar{\rho}{\bf u}\right|^{2}\left|{\bf u}\right|\operatorname{d}x\right|^{2}\operatorname{d}s\right)^{\frac{m}{2}}\right]\leqslant\mathbb{E}\left[\left(C\sup\limits_{s\in[0,t]}\left\|{\bf u}\right\|^{2}\int_{0}^{t}e^{2\alpha s}\left\|{\bf u}\right\|_{3}^{4}\operatorname{d}s\right)^{\frac{m}{2}}\right]$
$\leqslant$ $e^{\alpha mt}\mathbb{E}\left[\left(\sup\limits_{s\in[0,t]}\left\|{\bf u}\right\|^{2}\right)^{m}\right]+e^{\alpha mt}\mathbb{E}\left[\left(C\int_{0}^{t}\left\|{\bf u}\right\|_{3}^{3}\operatorname{d}s\right)^{m}\right]$ (3.8)
$\leqslant$ $e^{\alpha mt}\mathbb{E}\left[\left(C\int_{0}^{t}\left\|{\bf u}\right\|_{3}^{3}\operatorname{d}s\right)^{m}\right],$
where the last inequality holds due to the zeroth-order estimates in subsection 2.2, $C$ is a general constant. In summary, as $\bar{\rho}$ have a positive lower bound, we have
$\mathbb{E}\left[\left(\int_{0}^{t}e^{\alpha s}\operatorname{d}\left(\int_{U}\left|\mathbf{w}\right|^{2}\operatorname{d}x+\int_{U}\left|\nabla\phi\right|^{2}\operatorname{d}x\right)\right)^{m}\right]+\mathbb{E}\left[\left(\int_{0}^{t}e^{\alpha s}\int_{U}\left|{\bf u}\right|^{2}\operatorname{d}x\operatorname{d}s\right)^{m}\right]$
$\leqslant$ $e^{\alpha mt}\mathbb{E}\left[\left(C\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right)^{m}\right].$ (3.9)
Next, we give the estimates of $\int_{0}^{t}e^{\alpha s}\int_{U}\left\|\sigma\right\|^{2}\operatorname{d}x\operatorname{d}s$. From the velocity equation (2.1), we have
$e^{\alpha t}\left(\nabla Q\left(\bar{\rho}+\sigma\right)-\nabla Q\left(\bar{\rho}\right)\right)\operatorname{d}t$ (3.10)
$=$ $-e^{\alpha t}\operatorname{d}{\bf u}-e^{\alpha t}\left(\left({\bf u}\cdot\nabla\right){\bf u}-{\bf u}\right)\operatorname{d}t+e^{\alpha t}\nabla\phi\operatorname{d}t+e^{\alpha t}\frac{\mathbb{F}}{\bar{\rho}+\sigma}\operatorname{d}W,$
with
$\nabla\left(Q\left(\bar{\rho}+\sigma\right)-Q\left(\bar{\rho}\right)\right)=Q^{\prime}\left(\bar{\rho}+\sigma\right)\nabla\sigma+Q^{\prime\prime}\left(\bar{\rho}\right)\sigma\nabla\bar{\rho}+\mathbf{h},$
where
$h_{i}=O\left(\sigma^{2}\right).$ (3.11)
We multiply the equation (3.10) with $\left(\sigma,\sigma,\sigma\right)^{T}$. The left side of (3.10) is
$e^{\alpha t}\int_{U}\left|Q^{\prime\prime}\left(\bar{\rho}\right)\nabla\bar{\rho}\right|\left|\sigma\right|^{2}\operatorname{d}x+e^{\alpha t}\int_{U}O\left(\sigma^{3}\right)\operatorname{d}x.$ (3.12)
By Itô’s formula,
$e^{\alpha t}\left(\operatorname{d}u^{i}\right)\sigma=e^{\alpha t}\operatorname{d}\left(u^{i}\sigma\right)-e^{\alpha t}u^{i}\operatorname{d}\sigma,$ (3.13)
where
$-\int_{0}^{t}e^{\alpha s}\operatorname{d}\int_{U}\left(u^{i}\sigma\right)\operatorname{d}x\leqslant\int_{0}^{t}e^{\alpha s}\operatorname{d}\left(\frac{1}{2}\left\|\sigma\right\|^{2}+\frac{1}{2}\left\|u^{i}\right\|^{2}\right).$ (3.14)
By the continuity equation, it holds
$\int_{0}^{t}e^{\alpha s}\int_{U}\left|u^{i}\operatorname{d}\sigma\right|\operatorname{d}x\leqslant C\int_{0}^{t}e^{\alpha s}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s.$ (3.15)
For $-{\bf u}\operatorname{d}t$, we directly estimate
$\int_{0}^{t}\int_{U}e^{\alpha s}\left|-u^{i}\sigma\right|\operatorname{d}x\operatorname{d}s\leqslant\frac{\delta_{4}}{2}\int_{0}^{t}e^{\alpha s}\left\|\sigma\right\|^{2}\operatorname{d}s+C_{\delta_{4}}\int_{0}^{t}\left\|u^{i}\right\|^{2}\operatorname{d}s,$ (3.16)
where $\delta_{4}$ is small such that $\delta_{4}\int_{0}^{t}e^{\alpha s}\left\|\sigma\right\|^{2}\operatorname{d}s$ can be balanced by the left side. For the term $\nabla\phi\operatorname{d}t$ in (2.141), we estimate
$\int_{0}^{t}\int_{U}e^{\alpha s}\left|-\phi_{,i}\sigma\right|\operatorname{d}x\operatorname{d}s\leqslant\frac{\delta_{4}}{2}\int_{0}^{t}e^{\alpha s}\left\|\sigma\right\|^{2}\operatorname{d}s+C_{\delta_{4}}e^{\alpha t}\sup_{s\in[0,t]}\left\|\phi_{,i}\right\|^{2}.$ (3.17)
For the stochastic term, since $\left|\mathbb{F}\right|\leqslant C\left|\rho{\bf u}\right|^{2}$, we estimate
$\mathbb{E}\left[\left|\int_{0}^{t}e^{\alpha s}\int_{U}\frac{\mathbb{F}^{i}}{\bar{\rho}+\sigma}\operatorname{d}W\sigma\operatorname{d}x\right|^{m}\right]\leqslant\mathbb{E}\left[\left|C\int_{0}^{t}e^{\alpha s}\left|\int_{U}\frac{\mathbb{F}^{i}}{\bar{\rho}+\sigma}\sigma\operatorname{d}x\right|^{2}\operatorname{d}s\right|^{\frac{m}{2}}\right]$
$\leqslant$ $\mathbb{E}\left[\left|C\int_{0}^{t}e^{\alpha s}\left|\int_{U}\left|\bar{\rho}+\sigma\right|\left|{\bf u}\right|^{2}\sigma\operatorname{d}x\right|^{2}\operatorname{d}s\right|^{\frac{m}{2}}\right]\leqslant\mathbb{E}\left[\left|C\int_{0}^{t}e^{\alpha s}\left\|{\bf u}\right\|^{2}\left\|\bar{\rho}\sigma+\sigma^{2}\right\|_{\infty}^{2}\operatorname{d}s\right|^{\frac{m}{2}}\right]$
$\leqslant$ $e^{\alpha mt}\mathbb{E}\left[\left(\frac{1}{4}\sup_{s\in[0,t]}\left\|{\bf u}\right\|^{2}\right)^{m}\right]+\mathbb{E}\left[\left(C\int_{0}^{t}e^{\alpha s}\left\|{\bf u}\right\|^{2}\left\|\sigma\right\|_{\infty}^{2}\operatorname{d}s\right)^{m}\right]$
$+e^{\alpha mt}\mathbb{E}\left[\left(\frac{1}{4}\sup_{s\in[0,t]}\left\|{\bf u}\right\|^{2}\right)^{m}\right]+\mathbb{E}\left[\left(C\int_{0}^{t}e^{\alpha s}\left\|{\bf u}\right\|^{2}\left\|\sigma\right\|_{\infty}^{4}\operatorname{d}s\right)^{m}\right]$ (3.18)
$\leqslant$ $e^{\alpha mt}\mathbb{E}\left[\left(\frac{1}{2}\sup_{s\in[0,t]}\left\|{\bf u}\right\|^{2}\right)^{m}\right]+\mathbb{E}\left[\left(C\int_{0}^{t}e^{\alpha s}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right)^{m}\right]$
$\leqslant$ $e^{\alpha mt}\mathbb{E}\left[\left(C\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right)^{m}\right]+\mathbb{E}\left[\left(C\int_{0}^{t}e^{\alpha s}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right)^{m}\right].$
Therefore, we have
$\mathbb{E}\left[\left(\int_{0}^{t}e^{\alpha s}\left\|\sigma\right\|^{2}\operatorname{d}s\right)^{m}\right]$
$\leqslant$ $\mathbb{E}\left[\left(\int_{0}^{t}e^{\alpha s}\operatorname{d}\left(\frac{1}{2}\left\|\sigma\right\|^{2}+\frac{1}{2}\left\|{\bf u}\right\|^{2}\right)\right)^{m}\right]$ (3.19)
$+e^{\alpha mt}\mathbb{E}\left[\left(C\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right)^{m}\right]+\mathbb{E}\left[\left(C\int_{0}^{t}e^{\alpha s}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right)^{m}\right]$
$\leqslant$ $\mathbb{E}\left[\left(\int_{0}^{t}e^{\alpha s}\operatorname{d}\left(\frac{1}{2}\left\|\sigma\right\|^{2}+\frac{1}{2}\left\|{\bf u}\right\|^{2}\right)\right)^{m}\right]+e^{\alpha mt}\mathbb{E}\left[\left(C\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right)^{m}\right].$
Similarly, the estimates for $\mathbb{E}\left[\left(\int_{0}^{t}e^{\alpha s}\left\|\nabla\phi\right\|^{2}\operatorname{d}s\right)^{m}\right]$ holds:
$\mathbb{E}\left[\left(\int_{0}^{t}e^{\alpha s}\left\|\nabla\phi\right\|^{2}\operatorname{d}s\right)^{m}\right]$
$\leqslant$ $\mathbb{E}\left[\left(\int_{0}^{t}e^{\alpha s}\operatorname{d}\left(\frac{1}{2}\left\|\nabla\phi\right\|^{2}+\frac{1}{2}\left\|{\bf u}\right\|^{2}\right)\right)^{m}\right]$ (3.20)
$+e^{\alpha mt}\mathbb{E}\left[\left(C\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right)^{m}\right]+\mathbb{E}\left[\left(C\int_{0}^{t}e^{\alpha s}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right)^{m}\right]$
$\leqslant$ $\mathbb{E}\left[\left(\int_{0}^{t}e^{\alpha s}\operatorname{d}\left(\frac{1}{2}\left\|\nabla\phi\right\|^{2}+\frac{1}{2}\left\|{\bf u}\right\|^{2}\right)\right)^{m}\right]+e^{\alpha mt}\mathbb{E}\left[\left(C\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right)^{m}\right].$
Multiplying a small constant to (3.1.1) and (3.1.1), we plus the zero-order estimates (2.2.1) such that
$\mathbb{E}\left[\left(\int_{0}^{t}e^{\alpha s}\operatorname{d}\left(\frac{1}{2}\left\|\sigma\right\|^{2}+\left\|{\bf u}\right\|^{2}+\frac{1}{2}\left\|\nabla\phi\right\|^{2}\right)\right)^{m}\right]$ (3.21)
can be balanced by (2.2.1). Then we obtain
$\mathbb{E}\left[\left(\int_{0}^{t}e^{\alpha s}\operatorname{d}\left(\left\|\mathbf{w}\right\|^{2}+\left\|\nabla\phi\right\|^{2}\right)+\alpha\int_{0}^{t}e^{\alpha s}\left(\left\|\mathbf{w}\right\|^{2}+\left\|\nabla\phi\right\|^{2}\right)\operatorname{d}s\right)^{m}\right]$ (3.22)
$\leqslant$ $e^{\alpha mt}\mathbb{E}\left[\left(C\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}t\right)^{m}\right].$
3.1.2. First-order weighted estimates
Multiplying (2.2.2) by $e^{\alpha t}\nabla\mathbf{w}$ and integrating it over $U$, we can repeat the argument from subsection 3.1.1 to obtain:
$\mathbb{E}\left[\left|\int_{0}^{t}e^{\alpha s}\operatorname{d}\left(\int_{U}\mathcal{D}\nabla\mathbf{w}:\nabla\mathbf{w}\operatorname{d}x\right)\right|^{m}\right]+\mathbb{E}\left[\left|\int_{0}^{t}\alpha e^{\alpha s}\int_{U}\left|\nabla\mathbf{w}\right|^{2}\operatorname{d}x\operatorname{d}s\right|^{m}\right]$ (3.23)
$\leqslant$ $\mathbb{E}\left[\left|e^{\alpha t}\int_{0}^{t}C\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right|^{m}\right].$
3.1.3. Second-order weighted estimates
Similarly, we multiply (2.214) with $e^{\alpha t}\partial^{2}\mathbf{w}$, and then integrate it on $U$. Repeating the procedure in subsection 3.1.1, we have
$\mathbb{E}\left[\left|\int_{0}^{t}e^{\alpha s}\operatorname{d}\left(\int_{U}\left|\partial^{2}\mathbf{w}\right|^{2}\operatorname{d}x\right)\right|^{m}\right]+\mathbb{E}\left[\left|\int_{0}^{t}\alpha e^{\alpha s}\int_{U}\left|\partial^{2}\mathbf{w}\right|^{2}\operatorname{d}x\operatorname{d}s\right|^{m}\right]$ (3.24)
$\leqslant$ $\mathbb{E}\left[\left|e^{\alpha t}\int_{0}^{t}C\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right|^{m}\right].$
3.1.4. Third-order weighted estimates
Considering the $3$-order weighted estimates, following the standard bootstrap of subsection 3.1.1, we have
$\mathbb{E}\left[\left|\int_{0}^{t}e^{\alpha s}\operatorname{d}\left(\int_{U}\left|\partial^{3}\mathbf{w}\right|^{2}\operatorname{d}x\right)\right|^{m}\right]+\mathbb{E}\left[\left|\int_{0}^{t}\alpha e^{\alpha s}\int_{U}\left|\partial^{3}\mathbf{w}\right|^{2}\operatorname{d}x\operatorname{d}s\right|^{m}\right]$ (3.25)
$\leqslant$ $\mathbb{E}\left[\left|e^{\alpha t}\int_{0}^{t}C\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right|^{m}\right].$
3.2. Asymptotic stability
Combining the weighted estimates in the previous subsections, we obtain
$\mathbb{E}\left[\left|\int_{0}^{t}e^{\alpha s}\operatorname{d}\left(\left\|\mathbf{w}\right\|_{3}^{2}+\left\|\nabla\phi\right\|^{2}\right)+\int_{0}^{t}\alpha e^{\alpha s}\left(\left\|\mathbf{w}\right\|_{3}^{2}+\left\|\nabla\phi\right\|^{2}\right)\operatorname{d}s\right|^{m}\right]$ (3.26)
$\leqslant$ $\mathbb{E}\left[\left|e^{\alpha t}\int_{0}^{t}C\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right|^{m}\right].$
Therefore, we have
$\mathbb{E}\left[\left|e^{\alpha t}\left(\left\|\mathbf{w}\right\|_{3}^{2}+\left\|\nabla\phi\right\|^{2}\right)\right|^{m}\right]$ (3.27)
$\leqslant$ $\mathbb{E}\left[\left|\left(\left\|\mathbf{w}_{0}\right\|_{3}^{2}+\left\|\nabla\phi_{0}\right\|^{2}\right)\right|^{m}\right]+\mathbb{E}\left[\left|e^{\alpha t}\int_{0}^{t}C\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right|^{m}\right].$
Since $\left\|\mathbf{w}_{0}\right\|_{3}^{2}+\left\|\nabla\phi_{0}\right\|^{2}$ is small, we have
$\mathbb{E}\left[\left|e^{\alpha t}\left(\left\|\mathbf{w}\right\|_{3}^{2}+\left\|\nabla\phi\right\|^{2}\right)-e^{\alpha t}\int_{0}^{t}C\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\right|^{m}\right]\leqslant\mathbb{E}\left[\left|\left(\left\|\mathbf{w}_{0}\right\|_{3}^{2}+\left\|\nabla\phi_{0}\right\|^{2}\right)\right|^{m}\right].$ (3.28)
We estimate
$e^{\alpha t}\int_{0}^{t}\left\|\mathbf{w}\right\|_{3}^{3}\operatorname{d}s\leqslant e^{\alpha t}t\sup\limits_{s\in[0,t]}\left\|\mathbf{w}\right\|_{3}^{3}\leqslant e^{\frac{3\alpha t}{2}}\sup\limits_{s\in[0,t]}\left\|\mathbf{w}\right\|_{3}^{3}.$
Therefore, we obtain the asymptotic decay estimates
$\mathbb{E}\left[\left|\sup\limits_{s\in[0,t]}\left(\left\|\mathbf{w}\right\|_{3}^{2}+\left\|\nabla\phi\right\|^{2}\right)\right|^{m}\right]\leqslant e^{-\alpha mt}\mathbb{E}\left[\left|C\left(\left(\left\|\mathbf{w}_{0}\right\|_{3}^{2}+\left\|\nabla\phi_{0}\right\|^{2}\right)\right)\right|^{m}\right],$ (3.29)
on account that $\left\|\mathbf{w}_{0}\right\|_{3}^{2}+\left\|\nabla\phi_{0}\right\|^{2}$ is sufficiently small, where $m\geqslant 2$.
4. Invariant measures
The law generated by the initial data ${\bf{z}}_{0}:=\left(\rho_{0},{\bf u}_{0},\Phi_{0}\right)$ in probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$ is denoted by $\mathcal{L}\left({\bf{z}}_{0}\right)$. We denote $\mathscr{H}:=H^{3}\left(U\right)\times H^{3}\left(U\right)\times H^{5}\left(U\right)$. With the initial data ${\bf{z}}_{0}:=\left(\rho_{0},{\bf u}_{0},\Phi_{0}\right)\in\mathscr{H}$ and the assumptions of (1.20) and (1.21), SEP system (1.1) admits a unique strong solution
${\bf{z}}(t,x,\omega):=\left(\rho,{\bf u},\Phi\right)\in\mathscr{H}.$ (4.1)
Let $\mathcal{S}_{t}$ be the transition semigroup [48]:
$\mathcal{S}_{t}\psi({\bf{z}}_{0})=\mathbb{E}[\psi\left({\bf{z}}((t,{\bf{z}}_{0}))\right)],\quad t\geqslant 0,$ (4.2)
where $\psi$ is the bounded function on $\mathscr{H}$, i.e., $\psi\in C_{b}(\mathscr{H})$. $\mathcal{S}(t,{\bf{z}}_{0},\Gamma)$ is the transition function:
$\qquad\mathcal{S}(t,{\bf{z}}_{0},\Gamma):=\mathcal{S}_{t}({\bf{z}}_{0},\Gamma)=\mathcal{S}_{t}\mathds{1}_{\Gamma}({\bf{z}}_{0})=\mathcal{L}\left({\bf{z}}(t,{\bf{z}}_{0})\right)(\Gamma),~{}{\bf{z}}_{0}\in\mathscr{H},~{}\Gamma\in\mathscr{B}(\mathscr{H}),~{}t\geqslant 0.$ (4.3)
For ${\bf{v}}_{0}:=\left(\rho_{0}-\bar{\rho},{\bf u}_{0},\Phi_{0}-\bar{\Phi}\right)$ in probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$, the perturbed system (2.1) admits a unique strong solution
${\bf{v}}(t,x,\omega):=\left(\rho-\bar{\rho},{\bf u},\Phi-\bar{\Phi}\right)\in\mathscr{H}.$ (4.4)
$\tilde{\mathcal{S}}_{t}$ is the transition semigroup:
$\tilde{\mathcal{S}}_{t}\psi({\bf{v}}_{0})=\mathbb{E}[\psi\left({\bf{v}}((t,{\bf{v}}_{0}))\right)],\quad t\geqslant 0,$ (4.5)
where $\psi$ is the bounded function on $\mathscr{H}$, i.e., $\psi\in C_{b}(\mathscr{H})$. The transition function for the perturbed system (2.1) is denoted by $\tilde{\mathcal{S}}(t,{\bf{z}}_{0},\Gamma)$.
We give the definition of stationary solution for (1.1).
Definition 4.1.
A strong solution $\left(\rho;{\bf u};\Phi\right)$ to system (1.1) under the initial boundary conditions (1.5)-(1.6) is called stationary, provided that the transition function $\left(\mathcal{S}_{\tau}\rho,\mathcal{S}_{\tau}{\bf u},\mathcal{S}_{\tau}\Phi\right)$ on $C\left([0,T];H^{3}\left(U\right)\right)\times C\left([0,T];H^{3}\left(U\right)\right)\times C\left([0,T];H^{5}\left(U\right)\right)$ is independent of $\tau\geqslant 0$.
Let $\mathscr{M}\left(\mathscr{H}\right)$ be the space of all bounded measures on $\left(\mathscr{H},\mathscr{B}\left(\mathscr{H}\right)\right)$. For any $\psi\in C_{b}\left(\mathscr{H}\right)$ and any $\mu\in\mathscr{M}\left(\mathscr{H}\right)$, we set
$\langle\psi,\mu\rangle_{\mathscr{H}}=\int_{\mathscr{H}}\psi(x)\mu(\operatorname{d}x).$ (4.6)
For $t\geqslant 0$, $\mu\in\mathscr{M}\left(\mathscr{H}\right)$, $\mathcal{S}_{t}^{*}$ acts on $\mathscr{M}\left(\mathscr{H}\right)$ by
$\mathcal{S}_{t}^{*}\mu(\Gamma)=\int_{\mathscr{H}}\mathcal{S}(t,x,\Gamma)\mu(\operatorname{d}x),\quad\Gamma\in\mathscr{B}\left(\mathscr{H}\right).$ (4.7)
Moreover, there holds
$\left\langle\psi,\mathcal{S}_{t}^{*}\mu\right\rangle_{\mathscr{H}}=\left\langle\mathcal{S}_{t}\psi,\mu\right\rangle_{\mathscr{H}},\quad\forall~{}\psi\in C_{b}(\mathscr{H}),\quad\mu\in\mathscr{M}\left(\mathscr{H}\right).$ (4.8)
Particularly, for the perturbed system (2.1) and ${\bf{v}}_{0}:=\left(\rho_{0}-\bar{\rho},{\bf u}_{0},\Phi_{0}-\bar{\Phi}\right)$ in probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$, there holds $\mathcal{S}_{t}^{*}\mathscr{L}({\bf{v}}_{0})=\mathscr{L}({\bf{v}}(t,{\bf{v}}_{0}))$. In other words,
$\left(\mathcal{S}_{t}\psi\right)\mathcal{L}\left({\bf{v}}_{0}\right)=\mathbb{E}\left[\psi\left({\bf{v}}(t)\right)\right],$ (4.9)
where $\psi\in C_{b}(\mathscr{H})$.
Definition 4.2.
A measure $\mu$ in $\mathscr{M}(\mathscr{H})$ is said to be an invariant (stationary) measure if
$P_{t}^{*}\mu=\mu,\quad\forall~{}t>0.$ (4.10)
The Dirac measure centered at the steady state $\left(\bar{\rho},0,\bar{\Phi}\right)$ is the invariant measure for the (1.10), since it keeps unchange after the action of the transition semigroup for (1.10).
For ${\bf{z}}_{0}\in\mathscr{H}$ and $T>0$, the formula
$\frac{1}{T}\int_{0}^{T}\mathcal{S}_{t}({\bf{z}}_{0},\Gamma)\operatorname{d}t=R_{T}({\bf{z}}_{0},\Gamma),\quad\Gamma\in\mathscr{B}\left(\mathscr{H}\right),$ (4.11)
defines a probability measure. For any $\nu\in\mathscr{M}(H),R_{T}^{*}\nu$ is defined as follows:
$R_{T}^{*}\nu(\Gamma)=\int_{\mathscr{H}}R_{T}(x,\Gamma)\nu(\operatorname{d}x),\quad\Gamma\in\mathscr{B}\left(\mathscr{H}\right).$ (4.12)
For any $\psi\in C_{b}(\mathscr{H})$, there holds
$\left\langle R_{T}^{*}\nu,\psi\right\rangle_{\mathscr{H}}=\frac{1}{T}\int_{0}^{T}\left\langle\mathcal{S}_{t}^{*}\nu,\psi\right\rangle_{\mathscr{H}}\operatorname{d}t.$ (4.13)
$\mathcal{S}_{t}$, is a Feller semigroup provided that, for arbitrary $\psi\in C_{b}\left(\mathscr{H}\right)$, the function
$[0,+\infty)\times\mathscr{H},\quad(t,x)\mapsto\mathcal{S}_{t}\psi(x)$ (4.14)
is continuous. Since the solution is continuous and unique, we do not need the Markov selection as in [17, 27].
The method of constructing an invariant measure described in the following theorem is due to Krylov-Bogoliubov [38].
Theorem 4.1.
If for some $\nu\in\mathscr{M}\left(\mathscr{H}\right)$ and some sequence $T_{n}\uparrow+\infty,R_{T_{n}}^{*}\nu\rightarrow\mu$ weakly as $n\rightarrow\infty$, then $\mu$ is an invariant measure for Feller semigroup $\mathcal{S}_{t},t\geqslant 0$.
The following lemma is obtained similarly to [6], and we provide a proof for the convenience of the readers. ${\bf{v}}_{t}^{{\bf{v}}_{0}}$ represents the stochastic process initiated from ${\bf{v}}_{0}$ for the sake of expediency in exposition.
Lemma 4.1.
The SEP (2.1) defines a Feller-Markov process, i.e., $\tilde{\mathcal{S}}_{t}:C_{b}(\mathscr{H})\rightarrow C_{b}(\mathscr{H})$, and
$\mathbb{E}\left[\left.\psi\left({\bf{v}}_{t+s}^{{\bf{v}}_{0}}\right)\right|\mathcal{F}_{t}\right]=\left(\tilde{\mathcal{S}}_{s}\psi\right)\left({\bf{v}}_{t}^{{\bf{v}}_{0}}\right),\quad\forall~{}{\bf{v}}_{0}\in\mathscr{H},\quad\psi\in C_{b}(\mathscr{H}),\quad\forall~{}t,s>0,$ (4.15)
Proof. From the continuity of solutions, it is easy to see the Feller property that $\mathcal{S}_{t}:C_{b}(\mathscr{H})\rightarrow C_{b}(\mathscr{H})$ is continuous. For the Markov property, it suffices to prove
$\mathbb{E}\left[\psi\left({\bf{v}}_{t+s}^{{\bf{v}}_{0}}\right)X\right]=\mathbb{E}\left[\mathcal{S}_{s}\psi\left({\bf{v}}_{t}^{{\bf{v}}_{0}}\right)X\right],$ (4.16)
where $X\in\mathcal{F}_{t}$.
Let $\mathbf{D}$ be any $\mathcal{F}_{t}$-measurable random variable. We denote $\mathbf{D}_{n}=\sum\limits_{i=1}^{n}\mathbf{D}^{i}\mathbf{1}_{\Omega^{i}}$, where $\mathbf{D}^{i}\in H$ are deterministic and $\left(\Omega^{i}\right)\subset\mathcal{F}_{t}$ is a collection of disjoint sets such that $\bigcup\limits_{i}\Omega^{i}=\Omega$. $\mathbf{D}_{n}\rightarrow\mathbf{D}$ in $\mathscr{H}$ implies $\mathcal{S}_{t}\varphi\left(\mathbf{D}_{n}\right)\rightarrow\mathcal{S}_{t}\varphi(\mathbf{D})$ in $\mathscr{H}$. For every deterministic $\mathbf{D}\in\mathcal{F}_{t}$, the random variable ${\bf{v}}_{t,t+s}^{\mathbf{D}}$ depends only on the increments of the Brownian motion $W_{t+s}-W_{t}$ and hence it is independent of $\mathcal{F}_{t}$. Therefore, it holds
$\mathbb{E}\left[\psi\left({\bf{v}}_{t,t+s}^{\mathbf{D}}\right)X\right]=\mathbb{E}\left[\psi\left({\bf{v}}_{t,t+s}^{\mathbf{D}}\right)\right]\mathbb{E}[X],\quad\forall~{}\mathbf{D}\in\mathcal{F}_{t}.$ (4.17)
Since ${\bf{v}}_{t,t+s}^{\mathbf{D}}$ has the same law as ${\bf{v}}_{s}^{\mathbf{D}}$ by uniqueness, we have
$\mathbb{E}\left[\psi\left({\bf{v}}_{t,t+s}^{\mathbf{D}}\right)X\right]=\mathbb{E}\left[\psi\left({\bf{v}}_{s}^{\mathbf{D}}\right)\right]\mathbb{E}[X]=\mathcal{S}_{s}\psi(\mathbf{D})\mathbb{E}[X]=\mathbb{E}\left[\mathcal{S}_{s}\psi(\mathbf{D})X\right].$ (4.18)
Thus, there holds
$\mathbb{E}\left[\varphi\left({\bf{v}}_{t,t+s}^{\mathbf{D}}\right)X\right]=\mathbb{E}\left[\left(\mathcal{S}_{s}\varphi\right)(\mathbf{D})X\right]$ (4.19)
for every $\mathbf{D}$. By uniqueness, we have
${\bf{v}}_{t+s}^{{\bf{v}}_{0}}={\bf{v}}_{t,t+s}^{{\bf{v}}_{t}},\quad\mathbb{P}\quad{\rm a.s.},$ (4.20)
which completes the proof. $\square$ We shall prove the tightness of the law
$\left\{\frac{1}{T}\int_{0}^{T}\mathcal{L}\left(\mathbf{w}(t)\right)\times\mathcal{L}\left(\phi(t)\right)\operatorname{d}t,\quad T>0\right\},$ (4.21)
so as to apply Krylov-Bogoliubov’s theorem.
Theorem 4.2.
There exists an invariant measure for the system (2.1).
Proof. From the energy estimates of global existence, we know that
$\mathbb{E}\left[\left(\sup\limits_{t\in[0,T]}\left\|\mathbf{w}(t)\right\|_{3}^{2}\right)^{m}\right]\leqslant\mathbb{E}\left[\left(C\left(\left\|\mathbf{w}_{0}\right\|_{3}^{2}+\left\|\nabla\phi_{0}\right\|^{2}\right)\right)^{m}\right].$ (4.22)
The sets
$B_{L}:=\left\{\left.\mathbf{w}(t)\in H^{3}(U)\right|\left\|\mathbf{w}(t)\right\|_{3}\leqslant L\right\},\quad L>0,$ (4.23)
is compact in $C^{1}(U)$. Consequently, there holds
$\frac{1}{T}\int_{0}^{T}\mathcal{L}\left(\mathbf{w}(t)\right)\left(B_{L}^{c}\right)\operatorname{d}t=$ $\frac{1}{T}\int_{0}^{T}\mathbb{P}\left[\left\{\left\|\mathbf{w}(t)\right\|_{3}>L\right\}\right]\operatorname{d}t$
$\leqslant$ $\frac{1}{L^{2m}T}\int_{0}^{T}\mathbb{E}\left[\left\|\mathbf{w}(t)\right\|_{3}^{2m}\right]\operatorname{d}t$ (4.24)
$\leqslant$ $\frac{1}{L^{2m}}\mathbb{E}\left[\left(C\left(\left\|\mathbf{w}_{0}\right\|_{3}^{2}+\left\|\nabla\phi_{0}\right\|^{2}\right)\right)^{m}\right]$
$\rightarrow 0,\text{ as }L\rightarrow+\infty.$
This gives the tightness of $\frac{1}{T}\int_{0}^{T}\mathcal{L}\left(\mathbf{w}(t)\right)\operatorname{d}t$. The tightness of $\frac{1}{T}\int_{0}^{T}\mathcal{L}\left(\phi(t)\right)\operatorname{d}t$ is obtained similarly due to the energy estimate
$\mathbb{E}\left[\left|\sup\limits_{t\in[0,T]}\left\|\nabla\phi(t)\right\|^{2}\right|^{m}\right]\leqslant\mathbb{E}\left[\left(C\left(\left\|\mathbf{w}_{0}\right\|_{3}^{2}+\left\|\nabla\phi_{0}\right\|^{2}\right)\right)^{m}\right].$ (4.25)
Hence the tightness of (4.21) holds. Therefore, there exists an invariant measure by Krylov-Bogoliubov’s theorem. $\square$
Remark 4.1.
In the above proof, we need the constant in energy estimate (4.22) is independent on $T$. That is the reason why we assume (1.20) and (1.21).
(1.1) define a Feller-Markov process as well, similarly to (2.1). Since $\left(\bar{\rho},\bar{{\bf u}},\bar{\phi}\right)$ is smooth, by the uniqueness of solutions, $\frac{1}{T}\int_{0}^{T}\mathcal{L}\left(\rho\right)\times\mathcal{L}\left({\bf u}\right)\times\mathcal{L}\left(\Phi\right)\operatorname{d}s$ is also a tight measure, which generates an invariant measure. Actually, for compact sets
$B_{\rho,L}=\left\{\left.\rho\in H^{3}(U)\right|\left\|\rho\right\|_{3}\leqslant L\right\},\quad L>0,$ (4.26)
in $C^{1}(U)$, there holds
$\frac{1}{T}\int_{0}^{T}\mathcal{L}\left(\rho\right)\left(B_{L}^{c}\right)\operatorname{d}t=$ $\frac{1}{T}\int_{0}^{T}\mathbb{P}\left[\left\{\left\|\rho\right\|_{3}>L\right\}\right]\operatorname{d}t$
$\leqslant$ $\frac{1}{L^{2m}T}\int_{0}^{T}\mathbb{E}\left[\left\|\rho\right\|_{3}^{2m}\right]\operatorname{d}t$ (4.27)
$\leqslant$ $\frac{1}{L^{2m}}C\left(\mathbb{E}\left[\left\|\rho_{0}\right\|_{3}^{2m}\right]+\mathbb{E}\left[\left\|\bar{\rho}\right\|_{3}^{2m}\right]\right)$
$\rightarrow 0,\text{ as }L\rightarrow+\infty.$
We also care about what the limit of $\frac{1}{T}\int_{0}^{T}\mathcal{L}\left(\rho\right)\times\mathcal{L}\left({\bf u}\right)\times\mathcal{L}\left(\Phi\right)\operatorname{d}t$ is.
Theorem 4.3.
The invariant measure generated by $\frac{1}{T}\int_{0}^{T}\mathcal{L}\left(\rho\right)\times\mathcal{L}\left({\bf u}\right)\times\mathcal{L}\left(\Phi\right)\operatorname{d}t$, for system (1.1), is the Dirac measure of the steady state $\left(\bar{\rho},0,\bar{\Phi}\right)$. That is, the limit
$\lim\limits_{T\rightarrow+\infty}\frac{1}{T}\int_{0}^{T}\mathcal{L}\left(\rho\right)\times\mathcal{L}\left({\bf u}\right)\times\mathcal{L}\left(\Phi\right)\operatorname{d}t=\delta_{\bar{\rho}}\times\delta_{0}\times\delta_{\bar{\Phi}}$ (4.28)
holds weakly.
Proof. For any $\psi\in C_{b}\left(H^{3}\right)$, we have
$\lim_{T\rightarrow+\infty}\frac{1}{T}\int_{0}^{T}\langle\mathcal{L}\left(\rho\right),\psi\rangle_{\mathscr{H}}\operatorname{d}t=$ $\lim_{T\rightarrow+\infty}\frac{1}{T}\int_{0}^{T}\mathbb{E}\left[\psi\left(\rho\right)\right]\operatorname{d}t$ (4.29)
$=$ $\lim_{T\rightarrow+\infty}\frac{1}{T}\int_{0}^{T}\left(\mathbb{E}\left[\psi\left(\rho\right)-\psi\left(\bar{\rho}\right)\right]+\mathbb{E}\left[\psi\left(\bar{\rho}\right)\right]\right)\operatorname{d}t.$
We claim that $\lim\limits_{T\rightarrow+\infty}\frac{1}{T}\int_{0}^{T}\mathbb{E}\left[\psi\left(\rho\right)-\psi\left(\bar{\rho}\right)\right]\operatorname{d}s=0$. Actually, we separate $\Omega$ into
$\Omega_{t}=\left\{\psi\left(\rho\right)-\psi\left(\bar{\rho}\right)\leqslant\frac{1}{\sqrt{t}}\right\},\quad t>0,$ (4.30)
and $\Omega_{t}^{c}$. Then there holds
$\mathbb{E}\left[\psi\left(\rho\right)-\psi\left(\bar{\rho}\right)\right]=$ $\int_{\Omega}\left(\psi\left(\rho\right)-\psi\left(\bar{\rho}\right)\right)\mathbb{P}\left(\operatorname{d}\omega\right)$
$=$ $\int_{\Omega\cap\Omega_{t}}\left(\psi\left(\rho\right)-\psi\left(\bar{\rho}\right)\right)\mathbb{P}\left(\operatorname{d}\omega\right)+\int_{\Omega\cap\Omega_{t}^{c}}\left(\psi\left(\rho\right)-\psi\left(\bar{\rho}\right)\right)\mathbb{P}\left(\operatorname{d}\omega\right)$ (4.31)
$=$ $I_{1}+I_{2}.$
For $I_{1}$, it holds
$\lim_{T\rightarrow+\infty}\frac{1}{T}\int_{0}^{T}\int_{\Omega\cap\Omega_{t}}\left(\psi\left(\rho\right)-\psi\left(\bar{\rho}\right)\right)\mathbb{P}\left(\operatorname{d}\omega\right)\operatorname{d}t\leqslant\lim_{T\rightarrow+\infty}\frac{1}{T}\int_{0}^{T}\frac{1}{\sqrt{t}}\operatorname{d}t=0.$ (4.32)
For $I_{2}$, by the weighted energy estimates and Chebyshev’s inequality, there holds
$\int_{\Omega\cap\Omega_{t}^{c}}\left(\psi\left(\rho\right)-\psi\left(\bar{\rho}\right)\right)\mathbb{P}\left(\operatorname{d}\omega\right)\leqslant\int_{\Omega\cap\Omega_{t}^{c}}\left(\left|\psi\left(\rho\right)\right|+\left|\psi\left(\bar{\rho}\right)\right|\right)\mathbb{P}\left(\operatorname{d}\omega\right)$
$\leqslant$ $C\int_{\Omega\cap\Omega_{t}^{c}}\left(\left\|\rho\right\|_{3}+\left\|\bar{\rho}\right\|_{3}\right)\mathbb{P}\left(\operatorname{d}\omega\right)\leqslant C\mathbb{P}\left[\left\{\psi\left(\rho\right)-\psi\left(\bar{\rho}\right)>\frac{1}{\sqrt{t}}\right\}\right]$ (4.33)
$\leqslant$ $C\frac{\mathbb{E}\left[\left|\psi\left(\rho\right)-\psi\left(\bar{\rho}\right)\right|^{2m}\right]}{\left(\frac{1}{\sqrt{t}}\right)^{2m}}\leqslant Ct^{m}e^{-\gamma mt}\mathbb{E}\left[\left|\rho_{0}-\bar{\rho}\right|^{2m}\right].$
Hence, we have
$\lim_{T\rightarrow+\infty}\frac{1}{T}\int_{0}^{T}\int_{\Omega\cap\Omega_{t}^{c}}\left(\psi\left(\rho\right)-\psi\left(\bar{\rho}\right)\right)\mathbb{P}\left(\operatorname{d}\omega\right)\operatorname{d}t\leqslant\lim_{T\rightarrow+\infty}C\frac{1}{T}\int_{0}^{T}Ct^{m}e^{-\gamma mt}\operatorname{d}t=0.$ (4.34)
Therefore, there holds
$\lim_{T\rightarrow+\infty}\frac{1}{T}\int_{0}^{T}\langle\mathcal{L}\left(\rho\right),\psi\rangle\operatorname{d}t==$ $\lim_{T\rightarrow+\infty}\frac{1}{T}\int_{0}^{T}\mathbb{E}\left[\psi\left(\bar{\rho}\right)\right]\operatorname{d}t=\mathbb{E}\left[\psi\left(\bar{\rho}\right)\right]=\langle\delta_{\bar{\rho}},\psi\rangle.$ (4.35)
A similar calculation shows that
$\lim\limits_{T\rightarrow+\infty}\int_{0}^{T}\frac{1}{T}\mathcal{L}\left({\bf u}\right)\operatorname{d}t=\delta_{0};$ (4.36)
and
$\lim\limits_{T\rightarrow+\infty}\int_{0}^{T}\frac{1}{T}\mathcal{L}\left(\Phi\right)\operatorname{d}t=\delta_{\bar{\Phi}}.$ (4.37)
This completes the proof by the tightness of a joint distributions. $\square$
5. Appendix
We provide an overview of the fundamental theory concerning stochastic analysis. Let $E$ be a separable Banach space and $\mathscr{B}(E)$ be the $\sigma$-field of its Borel subsets, respectively. Let $\left(\Omega,\mathcal{F},\mathbb{P}\right)$ be a stochastic basis. A filtration $\mathcal{F}=\left(\mathcal{F}_{t}\right)_{t\in\mathbf{T}}$ is a family of $\sigma$-algebras on $\Omega$ indexed by $\mathbf{T}$ such that $\mathcal{F}_{s}\subseteq$ $\mathcal{F}_{t}\subseteq\mathcal{F}$, $s\leq t$, $s,t\in\mathbf{T}$. $\left(\Omega,\mathcal{F},\mathbb{P}\right)$ is also called a filtered space. We first list some definitions.
  1. 1.
    $E$-valued random variables. [48] For $(\Omega,\mathscr{F})$ and $(E,\mathscr{E})$ being two measurable spaces, a mapping $X$ from $\Omega$ into $E$ such that the set $\{\omega\in\Omega:X(\omega)\in A\}=\{X\in A\}$ belongs to $\mathscr{F}$ for arbitrary $A\in\mathscr{E}$, is called a measurable mapping or a random variable from $(\Omega,\mathscr{F})$ into $(E,\mathscr{E})$ or an $E$-valued random variable.
  2. 2.
    Strongly measurable operator valued random variables. [48] Let $\mathcal{U}$ and $\mathcal{H}$ be two separable Hilbert spaces which can be infinite-dimensional, and denote by $L(\mathcal{U},\mathcal{H})$ the set of all linear bounded operators from $\mathcal{U}$ into $\mathcal{H}$. A functional operator $\Psi(\cdot)$ from $\Omega$ into $L(\mathcal{U},\mathcal{H})$ is said to be strongly measurable, if for arbitrary $X\in\mathcal{U}$ the function $\Psi(\cdot)X$ is measurable, as a mapping from $(\Omega,\mathscr{F})$ into $(\mathcal{H},\mathscr{B}(\mathcal{H}))$. Let $\mathscr{L}$ be the smallest $\sigma$-field of subsets of $L(\mathcal{U},\mathcal{H})$ containing all sets of the form
    $\{\Psi\in L(\mathcal{U},\mathcal{H}):\Psi X\in A\},\quad X\in\mathcal{U},~{}A\in\mathscr{B}(\mathcal{H}),$ (5.1)
    then $\Psi:\Omega\rightarrow L(\mathcal{U},\mathcal{H})$ is a strongly measurable mapping from $(\Omega,\mathscr{F})$ into $(L(\mathcal{U},\mathcal{H}),\mathscr{L})$.
  3. 3.
    Law of a random variable. For an $E$-valued random variable $X:(\Omega,\mathcal{F})\rightarrow(E,\mathscr{E})$, we denote by $\mathcal{L}[X]$ the law of $X$ on $E$, that is, $\mathcal{L}[X]$ is the probability measure on $(E,\mathscr{E})$ given by
    $\mathcal{L}[X](A)=\mathbb{P}[X\in A],\quad A\in\mathscr{E}.\$ (5.2)
  4. 4.
    Stochastic process. [48] A stochastic process $X$ is defined as an arbitrary family $X=\{X_{t}\}_{t\in\mathbf{T}}$ of $E$-valued random variables $X_{t}$, $t\in\mathbf{T}$. $X$ is also regarded as a mapping from $\Omega$ into a Banach space like $C([0,T];E)$ or $L^{p}=L^{p}(0,T;E),1\leq p<+\infty$, by associating $\omega\in\Omega$ with the trajectory $X(\cdot,\omega)$.
  5. 5.
    Cylindrical Wiener Process valued in Hilbert space. [48] A $\mathcal{U}$-valued stochastic process $W(t),t\geqslant 0$, is called a cylindrical Wiener process if
    • $W(0)=0$;
    • $W$ has continuous trajectories;
    • $W$ has independent increments;
    • The distribution of $(W(t)-W(s))$ is $\mathscr{N}(0,(t-s)),\quad 0\leqslant s\leqslant t$.
  6. 6.
    Adapted stochastic process. A stochastic process $X$ is $\mathcal{F}$-adapted if $X_{t}$ is $\mathcal{F}_{t}$-measurable for every $t\in\mathbf{T}$;
  7. 7.
    Martingale. The $E$-valued process $X$ is called integrable provided $\mathbb{E}\left[\|X_{t}\|\right]<+\infty$ for every $t\in\mathbf{T}$. An integrable and adapted $E$-valued process $X_{t},t\in\mathbf{T}$, is a martingale if
    • $X$ is adapted;
    • $X_{s}=\mathbb{E}\left[X_{t}\mid\mathcal{F}_{s}\right]$, for arbitrary $t,s\in\mathbf{T},~{}0\leqslant s\leqslant t$.
  8. 8.
    Stopping time. On $\left(\Omega,\mathcal{F},\mathbb{P}\right)$, a random time is a measurable mapping $\tau:\Omega\rightarrow\mathbf{T}\cup\infty$. A random time is a stopping time if $\{\tau\leq t\}\in\mathcal{F}_{t}$ for every $t\in\mathbf{T}$. For a process $X$ and a subset $V$ of the state space we define the hitting time of $X$ in $V$ as
    $\tau_{V}(\omega)=\inf\left\{\left.t\in\mathbf{T}\right|X_{t}(\omega)\in V\right\}.$ (5.3)
    If $X$ is a continuous adapted process and $V$ is closed, then $\tau_{V}$ is a stopping time.
  9. 9.
    Modification. A stochastic process $Y$ is called a modification or a version of $X$ if
    [$\mathbb{P}[\{\omega\in\Omega:X(t,\omega)\neq Y(t,\omega)\}]=0\quad\text{ for all }t\in\mathbf{T}.$ (5.4)
  10. 10.
    Progressive measurability. In $\left(\Omega,\mathcal{F},\mathbb{P}\right)$, stochastic process $X$ is progressively measurable or simply progressively measurable, if for $\omega\in\Omega$, $(\omega,s)\mapsto X(s,\omega),~{}s\leqslant t$ is $\mathcal{F}_{t}\otimes\mathscr{B}(\mathbf{T}\cap[0,t])$-measurable for every $t\in\mathbf{T}$.
  11. 11.
    Progressive measurability of continuous functions. Let $X(t),t\in[0,T]$, be a stochastically continuous and adapted process with values in a separable Banach space $E$. Then $X$ has a progressively measurable modification.
  12. 12.
    Cross quadratic variation. Fixing a number $T>0$, we denote by $\mathcal{M}_{T}^{2}(E)$ the space of all $E$-valued continuous, square integrable martingales $M$, such that $M(0)=0$. If $M\in\mathcal{M}_{T}^{2}\left(\mathbb{R}^{1}\right)$ then there exists a unique increasing predictable process $\langle M(\cdot)\rangle$, starting from 0 , such that the process
    $M^{2}(t)-\langle M(\cdot)\rangle,\quad t\in[0,T]$ (5.5)
    is a continuous martingale. The process $\langle M(\cdot)\rangle$ is called the quadratic variation of $M$. If $M_{1},M_{2}\in\mathcal{M}_{T}^{2}\left(\mathbb{R}^{1}\right)$ then the process
    $\left\langle M_{1}(t),M_{2}(t)\right\rangle=\frac{1}{4}\left[\left\langle\left(M_{1}+M_{2}\right)(t)\right\rangle-\left\langle\left(M_{1}-M_{2}\right)(t)\right\rangle\right]$ (5.6)
    is called the cross quadratic variation of $M_{1},M_{2}$. It is the unique, predictable process with trajectories of bounded variation, starting from 0 such that
    $M_{1}(t)M_{2}(t)-\left\langle M_{1}(t),M_{2}(t)\right\rangle,\quad t\in[0,T]$ (5.7)
    is a continuous martingale.
    For $M\in\mathcal{M}_{T}^{2}(\mathcal{H})$, where $\mathcal{H}$ is Hilbert space, the quadratic variation is defined by
    $\langle M(t)\rangle=\sum_{i,j=1}^{\infty}\left\langle M_{i}(t),M_{j}(t)\right\rangle e_{i}\otimes e_{j},\quad t\in[0,T],$ (5.8)
    as an integrable adapted process, where $M_{i}(t)$ and $M_{j}(t)$ are in $\mathcal{M}_{T}^{2}\left(\mathbb{R}^{1}\right)$. If $a\in\mathcal{H}_{1},b\in\mathcal{H}_{2}$, then $a\otimes b$ denotes a linear operator from $\mathcal{H}_{2}$ into $\mathcal{H}_{1}$ given by the formula
    $(a\otimes b)x=a\langle b,x\rangle_{\mathcal{H}_{2}},~{}x\in\mathcal{H}_{2}.$ (5.9)
    We define a cross quadratic variation for $M^{1}\in\mathcal{M}_{T}^{2}\left(\mathcal{H}_{1}\right)$, $M^{2}\in\mathcal{M}_{T}^{2}\left(\mathcal{H}_{2}\right)$ where $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are two Hilbert spaces. Namely we define
    $\left\langle M^{1}(t),M^{2}(t)\right\rangle=\sum_{i,j=1}^{\infty}\left\langle M_{i}^{1}(t),M_{j}^{2}(t)\right\rangle e_{i}^{1}\otimes e_{j}^{2},\quad t\in[0,T],$ (5.10)
    where $\left\{e_{i}^{1}\right\}$ and $\left\{e_{j}^{2}\right\}$ are complete orthonormal bases in $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ respectively.
  13. 13.
    Stochastic integral. Let $W$ be the Wiener process. Let $\Psi(t),t\in[0,T]$, be a measurable Hilbert–Schmidt operators in $L(\mathcal{U},\mathcal{H})$, which is set in the space $\mathcal{L}_{2}$ such that
    $\mathbb{E}\left[\int_{0}^{t}\|\Psi(s)\|_{\mathcal{L}_{2}}^{2}\operatorname{d}s\right]:=\mathbb{E}\int_{0}^{t}\langle\Psi(s),\Psi^{\star}(s)\rangle_{\mathcal{H}}\operatorname{d}s<+\infty,$ (5.11)
    where $\langle\cdot,\cdot\rangle_{\mathcal{H}}$ means the inner product in $\mathcal{H}$. For the stochastic integral $\int_{0}^{t}\Psi\operatorname{d}W$, there holds
    $\mathbb{E}\left[\left(\int_{0}^{t}\Psi\operatorname{d}W\right)^{2}\right]=\mathbb{E}\left[\int_{0}^{t}\|\Psi(s)\|_{\mathcal{L}_{2}}^{2}\operatorname{d}s\right].$ (5.12)
    Furthermore, the following properties hold
    • Linearity: $\int(a\Psi_{1}+b\Psi_{2})\operatorname{d}W=a\int\Psi_{1}\operatorname{d}W+b\int\Psi_{2}\operatorname{d}W$ for constants $a$ and $b$;
    • Stopping property: $\int 1_{\{\cdot\leq\tau\}}\Psi\operatorname{d}W=\int\Psi\operatorname{d}M^{\tau}=\int_{0}^{\cdot\wedge\tau}\Psi\operatorname{d}W$;
    • Itô-isometry: for every $t$,
      $\mathbb{E}\left[\left(\int_{0}^{t}\Psi\operatorname{d}W\right)^{2}\right]=\mathbb{E}\left[\int_{0}^{t}\|\Psi(s)\|_{\mathcal{L}_{2}}^{2}\operatorname{d}s\right].$ (5.13)
  14. 14.
    Dirac measure. Let $(E,\mathscr{B}(E))$ be a measurable space. Given $x\in E$, the Dirac measure $\delta_{x}$ at $x$ is the measure defined by
    $\delta_{x}(A):=\begin{cases}1,&x\in A\\ 0,&x\notin A\end{cases}$ (5.14)
    for each measurable set $A\subseteq E$. In this paper, there holds
    $\delta_{\bar{\rho}}=\mathcal{L}[\bar{\rho}](A)=\mathbb{P}\left[\left\{\omega\in\Omega|\bar{\rho}(x)\in A\right\}\right]=1.$
  15. 15.
    Tightness of measures. [2] Let $E$ be a Hausdorff space, and let $\mathscr{E}$ be a $\sigma$-algebra on $E$. Let $\mathscr{M}$ be a collection of measures defined on $\mathscr{E}$. The collection $\mathscr{M}$ is called tight if, for any $\varepsilon>0$, there is a compact subset $K_{\varepsilon}$ of $E$ such that, for all measures $\mu\in\mathscr{M}$,
    $|\mu|\left(E\backslash K_{\varepsilon}\right)<\varepsilon,$ (5.15)
    where $|\mu|$ is the total variation measure of $\mu$. More specially, for probability measures $\mu$, (5.15) can be written as
    $\mu\left(K_{\varepsilon}\right)>1-\varepsilon.$ (5.16)
We list some important theorems in stochastic analysis.
  1. 1.
    Itô’s formula. [34, 48] Assume that $\Psi$ is an $\mathcal{L}_{2}$-valued process stochastically integrable in $[0,T],\varphi$ being a $\mathcal{H}$-valued predictable process Bochner integrable on $[0,T],\mathbb{P}$-a.s., and $X(0)$ being a $\mathscr{F}_{0}$-measurable $\mathcal{\mathcal{H}}$-valued random variable. Then the following process
[…]
    is well defined. Assume that a function $F:[0,T]\times\mathcal{H}\rightarrow\mathbb{R}^{1}$ and its partial derivatives $F_{t},F_{x},F_{xx}$, are uniformly continuous on bounded subsets of $[0,T]\times\mathcal{H}$. Under the above conditions, $\mathbb{P}$-a.s., for all $t\in[0,T]$,
    $~{}\qquad F(t,X(t))=$ $F(0,X(0))+\int_{0}^{t}\left\langle F_{x}(s,X(s)),\Psi(s)\operatorname{d}W(s)\right\rangle_{\mathcal{H}}$ (5.18)
    $+\int_{0}^{t}\left\{F_{t}(s,X(s))+\left\langle F_{x}(s,X(s)),\varphi(s)\right\rangle_{\mathcal{H}}+\frac{1}{2}F_{xx}(s,X(s))\|\Psi(s)\|_{\mathcal{L}_{2}}^{2}\right\}\operatorname{d}s.$
    Applying the above formula for $F=\langle x,x\rangle_{\mathcal{H}}$, we have Itô’s formula for $\langle X,X\rangle_{\mathcal{H}}$. Then by
    $\langle X,Y\rangle_{\mathcal{H}}=\frac{\langle X+Y,X+Y\rangle_{\mathcal{H}}-\langle X-Y,X-Y\rangle_{\mathcal{H}}}{4}$ (5.19)
    in Hilbert space, the following Itô’s formula holds for $X$ and $Y$ in form of (5.17),
    $\langle X,Y\rangle_{\mathcal{H}}$ $=\langle X_{0},Y_{0}\rangle_{\mathcal{H}}+\int\langle X,\operatorname{d}Y\rangle_{\mathcal{H}}+\int\langle Y,\operatorname{d}X\rangle_{\mathcal{H}}+\int\operatorname{d}\left\langle~{}\langle X,Y\rangle,\langle X,Y\rangle~{}\right\rangle_{\mathcal{H}}^{\frac{1}{2}}$ (5.20)
    $=\langle X_{0},Y_{0}\rangle_{\mathcal{H}}+\int\langle X,\operatorname{d}Y\rangle_{\mathcal{H}}+\int\langle Y,\operatorname{d}X\rangle_{\mathcal{H}}+\left\langle~{}\langle X,Y\rangle,\langle X,Y\rangle~{}\right\rangle_{\mathcal{H}}^{\frac{1}{2}},$
    where $\langle X,Y\rangle$ means the cross quadratic variation of $X$ and $Y$ defined above.
  2. 2.
    Chebyshev’s inequality. Let $Y$ be a random variable in probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$, $\varepsilon>0$. For every $0<r<\infty$, Chebyshev’s inequality reads
    [$\mathbb{P}[\left\{|Y|\geq\varepsilon\right\}]\leq\frac{1}{\varepsilon^{r}}\mathbb{E}\left[|Y|^{r}\right].$ (5.21)
  3. 3.
    Burkholder-Davis-Gundy’s inequality. [7, 48] Let $M$ be a continuous local martingale in $\mathcal{H}$. Let $M^{\ast}=\max\limits_{0\leqslant s\leqslant t}|M(s)|$, for any $m\geqslant 1$. $\langle M\rangle_{T}$ denotes the quadratic variation stopped by $T$. Then there exist constants $K^{m}$ and $K_{m}$ such that
    $K_{m}\mathbb{E}\left[\left(\langle M\rangle_{T}\right)^{m}\right]\leqslant\mathbb{E}\left[\left(M^{\ast}_{T}\right)^{2m}\right]\leqslant K^{m}\mathbb{E}\left[\left(\langle M\rangle_{T}\right)^{m}\right],$ (5.22)
    for every stopping time $T$. For $m\geqslant 1$, $K^{m}=\left(\frac{2m}{2m-1}\right)^{\frac{2m(2m-2)}{2}}$, which is equivalent to $e^{m}$ as $m\rightarrow\infty$. Specifically, for every $m\geqslant 1$, and for every $t\geqslant 0$, there holds
    $\mathbb{E}\left[\sup_{s\in[0,t]}\left|\int_{0}^{t}\Psi(s)\operatorname{d}W(s)\right|^{2m}\right]\leq K^{m}\left(\mathbb{E}\left[\int_{0}^{t}\|\Psi(s)\|_{\mathcal{L}_{2}}^{2}\operatorname{d}s\right]\right)^{m}$ (5.23)
  4. 4.
    Stochastic Fubini theorem. Assume that $(E,\mathscr{E})$ is a measurable space and let
    $\Psi:(t,\omega,x)\rightarrow\Psi(t,\omega,x)$
    be a measurable mapping from $\left(\Omega_{T}\times E,\mathscr{B}(\Omega_{T})\times\mathscr{B}(E)\right)$ into $\left(\mathcal{L}^{2},\mathscr{B}\left(\mathcal{L}^{2}\right)\right)$. Assume moreover that
    $\int_{E}\left[\mathbb{E}\int_{0}^{T}\langle\Psi(s),\Psi^{\star}(s)\rangle_{\mathcal{H}}\operatorname{d}t\right]^{\frac{1}{2}}\mu(\operatorname{d}x)<+\infty,$ (5.24)
    then $\mathbb{P}$-a.s. there holds
    $\int_{E}\left[\int_{0}^{T}\Psi(t,x)\operatorname{d}W(t)\right]\mu(\operatorname{d}x)=\int_{0}^{T}\left[\int_{E}\Psi(t,x)\mu(\operatorname{d}x)\right]\operatorname{d}W(t).$ (5.25)
  5. 5.
    Kolmogorov-Centov’s continuity theorem. [35, 48] Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space and $\bar{X}$ a process on $[0,T]$ with values in a complete metric space $(E,\mathscr{E})$. Suppose that
    $\mathbb{E}\left[\left|\bar{X}_{t}-\bar{X}_{s}\right|^{a}\right]\leq C|t-s|^{1+b},$ (5.26)
    for every $s<t\leq T$ and some strictly positive constants $a,b$ and $C$. Then $\bar{X}$ admits a continuous modification $X$, [$\mathbb{P}\left[\left\{X_{t}=\bar{X}_{t}\right\}\right]=1$ for every $t$, and $X$ is locally Hölder continuous for every exponent $0<\gamma<\frac{b}{a},$ namely,
    [$\mathbb{P}\left[\left\{\omega:\sum_{0<t-s<h(\omega),t,s\leq T}\frac{\left|X_{t}(\omega)-X_{s}(\omega)\right|}{|t-s|^{\gamma}}\leq\delta\right\}\right]=1,$ (5.27)
    where $h(\omega)$ is an strictly positive random variable a.s., and the constant satisfies $\delta>0$.
Acknowledgements. The authors would like to express their thanks to Prof. Deng Zhang for the valuable discussions. This work was commenced when L. Zhang visited McGill University as a joint Ph.D training student supported by China Scholarship Council (CSC). She would like to express her gratitude to McGill University and CSC. The research of Y. Li was supported in part by National Natural Science Foundation of China under grants 12371221, 12161141004, and 11831011. Y. Li was also grateful to the supports by the Fundamental Research Funds for the Central Universities and Shanghai Frontiers Science Center of Modern Analysis. The research of M. Mei was supported by NSERC grant RGPIN 2022-03374 and NNSFC Grant W2431005.